! 1

A Complete Bibliography of Publications in HardwareX

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

04 December 2023 Version 1.00

Title word cross-reference

1000μ [BM18]. 16 [DB22]. 2 [NIKK21, ZCB^{+}20]. 2.5T [PAMH18]. 3 [ASIM21, AHUS19, BS21, BHF^{+}22, BOB^{+}19, BEES22, BVMHP22a, BM18, CVT^{+}21, CKB^{+}22, DS21, DÖ23, DBP^{+}21, DKM23, DOG21, GPO^{+}20, GCM21, HS21, HSS22, HW22, HHSP18, JKKK22, KLS22, KTK22, KTT^{+}21, MWG^{+}22, MSWC21, MKL^{+}21, NSHA22, OKH22, PTKG21, PWC23, PDR^{+}22, PHF18, RJL^{+}23, RWK22, SMS^{+}23, SK23, SSL^{+}20, STE^{+}20, STP20, SKK^{+}23, SFWW22, TSF21, TPLT22, TBBI21, TLP^{+}23, VBK^{+}22, VCS^{+}23, WS18b, WBD^{+}21, WMPP18, YXES21]. 32 [FLL23]. _{2} [BH22, BGE^{+}20, LKL^{+}21, MKM^{+}23, TLO^{+}22]. H [PRR^{+}23]. ∞ [RMOB^{+}19]. L [BM18]. μ [AP22, GHHR20].

T [BGJGC23].

-based [SG21, DB22]. -board [AP22].

-bridge [PRR^{+}23]. -channel

[DB22, FLL23]. -D

[ASIM21, BVMHP22a, GPO^{+}20, HHSP18, STP20, TPLT22, WMPP18]. -measuring [BH22]. -photon [ZCB^{+}20].

1 [GCKF22]. 19 [CABA^{+}21, GPO^{+}20, HP20, MNMN^{+}20, MDR20, NMZMT20, NHAH^{+}21, Pea21, RRS^{+}23, VA21].

2.0 [HPU21]. 2017 [Ano17d, Ano17c]. 2018 [Ano18b, Ano18c]. 2019 [Ano19a, Ano19b]. 2020 [Ano20a, Ano20b]. 2021 [Ano21a, Ano21c]. 2022 [Ano22a, Ano22b]. 2023 [Ano23b, Ano23c, Ano23a, Ano23d].

3D [JTD20].

4.0 [PBD^{+}22].

7 [KTK22].

abdominal [GCHP22]. ablated [RGSGD^{+}23]. Ablation [NAN19]. aboard [Pri19]. absolute [BVMHP22b]. absorbance [POT^{+}22]. AC [MRKP23]. AC/off [MRKP23]. AC/off-grid [MRKP23]. academic [PDMV23]. access [SGD19]. accessible [MPB21, VS22]. acidification [OFP^{+}23]. acoustic [HPS^{+}19, hTFC^{+}23, ZBOB22]. acoustical [MBML21]. acoustically [ZBOB22]. Acousto [BPP^{+}19]. Acousto-Optic [BPP^{+}19]. acquisition [BBJ21, BCB^{+}22, BVVAD21, BVVA21, KwMMM23, LBK22, MSR^{+}23, NAC^{+}22, OWHP23, PBD^{+}22, SKC^{+}20, SRC^{+}22, THL^{+}20]. acrylic [WDW21]. Actifield [KAT18]. actimeter [KAT18]. active [FPC^{+}19, McM17, MMR^{+}23, SL22, SAB^{+}18, UMEA19]. activity [SRB^{+}22]. actuation [UMEA19]. actuator [BS22b]. AD5933 [MMR22]. Adaptable [LB17, LHN^{+}23, CJ20, NLYB23]. adapted [MNMN^{+}20]. adapter [KSB22]. Adaptive [Kod18, MPASIM23]. additive [POP^{+}20, RNRP23]. adherent [PK21]. adhesion [LSD22, MZLG20, PNM^{+}23, VSTB19]. adjustable [CKMW21, RAMM20, WSI^{+}22]. administration [KLF^{+}20]. ADO [SBTPV22]. Adsorption [GAB^{+}22, RHR19]. advanced [PMCL20, VKM^{+}23, ZSLF23]. aerial [LHZ^{+}18, MHBVBVMA23]. aerially [NMZMT20]. aerodynamic [IPR22]. aerosol [KWP22]. AFE4300 [FLL23]. affordable [BSE20, CFV^{+}19, DÖ23, DKH^{+}23, GST^{+}22, PHL^{+}20, SMS^{+}23, THL^{+}20]. Africa [BSS^{+}20]. against [SWPG19]. Agenator [LRSC19]. aging [LRSC19]. agricultural [DMS^{+}22]. agriculture [BVMHP22c]. air [ARLS22, BSO^{+}23, BVBTMVP23, HP20, JBG^{+}22, PTKG21, PHLM22, RAMM20, Sur21]. air-purifying [HP20]. airborne [JBG^{+}22]. AirQo [BSO^{+}23]. albedometer [BVMHP22b]. algal [NND^{+}22]. algorithm [IBL^{+}17]. algorithms [PMCL20]. aligner [GST^{+}22]. All-in-One [SIY^{+}23, WS19, NGSU21]. all-iron [KYA21, YA19]. allowing [KHRC21]. allows [PJS22]. alone [KNAR22, PMA^{+}23]. alternative [KLF^{+}20]. altitude [LLO22, Pri19]. aluminum [TBBI21]. amplification [HYdM21]. amplifier [CKMW21]. anaerobic [CPB^{+}22, HG21]. analogue [MBP19]. analyses [CED20]. Analysis [GME^{+}22, CFRF^{+}23, FLL23, INS^{+}23, KKV20, LOSM18, LDN^{+}22, MCLSN19, MHBVBVMA23, MSR^{+}23, MTPK20, RMB18, SMS^{+}23, VOS22, WNKP22]. analytical [LWS^{+}22]. analyzer [CNBH22, HSS22, LKL^{+}21, POP^{+}20, Rom21, RJL^{+}23]. Android [BGJGC23, Poó23]. angle [BVVAGMRP22, BVORMH^{+}22, HSS22]. Animal [GMB20, CFRF^{+}23, LHN^{+}23, OHF17, PHLM22]. antenna [CPVCRLS17, DMA^{+}23, Kod18]. antennas [JKKK22]. anthropomorphic [Lup17, PK20]. Antibody [LNP^{+}23, HS21]. ants [SRB^{+}22]. AntVideoRecord [SRB^{+}22]. any [JTD20]. APELLA [SIY^{+}23]. apertures [LKN21]. Aplysia [KOBT23]. apparatus [CBTGAG^{+}18, FK19, GLG^{+}23, HP20, KAB^{+}20, Pri19, SWPG19, SK21, WCR22]. apparent [CPCQY23]. Application [SKK^{+}23, CKB^{+}22, KOBT23, Kod18]. applications [AC20, BOB^{+}19, BVMHBV21, Bie23, BKM^{+}21, BVVA21, BVMHP22c, BJL^{+}20, BG21, CFD^{+}20, DMS^{+}22, GLD18, ICBL22, KW23, LKL^{+}21, LWS^{+}22, LBK22, MBP19, MPB21, MMR^{+}23, NRK^{+}23, POT^{+}22, RJL^{+}23, STP20, SKK^{+}23, TLŠ^{+}22]. applied [GC23, MBML21]. approach [BGE^{+}20, ILL18, RGSGD^{+}23]. Approaches [HHSP18, MBH^{+}21]. appropriate [DMA^{+}23]. April [Ano17d, Ano18c, Ano19a, Ano20a, Ano21a, Ano22a]. aquaculture [SBTPV22]. aquatic [HJvdG^{+}22, MBC^{+}20, TS22]. aqueous [Bow23]. arbitrary [PK23, ZCB^{+}20]. arc [POP^{+}20]. ARDS [CABA^{+}21]. Arduino [LNP^{+}23, BMOTGHL23, CNPP21, CRLF^{+}23, DC23, DKM23, GFLMRV^{+}23, GAB^{+}22, KW23, LLO22, LBK22, MPB21, PLP23, SOB^{+}23, SGD19, dCSMVFR23]. Arduino-based [CNPP21, KW23, LBK22, MPB21, SGD19]. Arduino-operated [LNP^{+}23]. area [MD22]. areas [PPAP21, PEJ^{+}22, RBB^{+}22]. arm [BSE20, KFB^{+}23, KG22, TCG^{+}21]. array [MPB^{+}23, Poó23]. arrays [FRJ^{+}22, MD22]. Arsenic [GME^{+}22]. art [MDR20]. arthropods [McM17]. articles [Ano21b]. ASAP [SMS^{+}23]. assay [HS21, TBM20]. assembled [DHY18, KST19, Kur21]. assembly [GST^{+}22]. Assessing [CAR^{+}23]. assessment [DHY18, MBML21]. asymmetric [VOS22]. ATMO [MNMN^{+}20]. ATMO-vent [MNMN^{+}20]. atmosphere [MNMN^{+}20]. atmospheric [MMP^{+}19, Pri19, RBB^{+}22, SMS^{+}23]. atomic [GFLMRV^{+}23, LAW^{+}22]. atomically [PÁBS23]. attachment [GCK^{+}23]. AudioMoth [HPS^{+}19]. Augmentation [ZSLF23]. Australia [KMB^{+}22]. Authors [Pea21]. Auto [CSH18]. Auto-HPGe [CSH18]. automatable [SGH^{+}18]. automate [LDN^{+}22]. Automated [DOV^{+}21, GME^{+}22, HHKB20, LBK^{+}21, MECH21, MKH^{+}20, MD20, BCN^{+}21, BS22a, BGJGC23, BKM^{+}21, BM19, BPM^{+}18, CHBMCV^{+}23, DCY^{+}22, ELL21, EMMH20, FDD17, FES^{+}21, FK19, ILG23, KAT18, MCLSN19, NIKK21, PTD^{+}20, SMW^{+}20, SKG22, VOS22, VSTB19, WS19, YLW^{+}23]. automatic [AG23, OHF17, NAN19]. automating [DB22]. Automation [GFLMRV^{+}23, MN22, OWHP23, WLML23]. automized [HG21]. Autonomous [MDD20, MDD^{+}21, NLYB23, SRB^{+}22, CFV^{+}19, CAR^{+}23, LKN21, MBC^{+}20, PEJ^{+}22, RJL^{+}23, Ryu22a, SMB^{+}21]. autonomously [KKwM22]. autosampler [CSH18, CM18, CED20, Car20, Car21, DOG21]. autostainer [PDR^{+}22]. Autotitrator [dCSMVFR23]. avalanche [MKVS21]. avoidance [CFV^{+}19]. axis [AFK^{+}19, GMB20].

bacteria [CPB^{+}22]. bacterial

[PPAP21, WDW21]. bag [PTD^{+}20]. ball [MRKP23]. balloon [LLO22]. balloons [Pri19]. bandwidth [LPB18]. banks [DBRE21]. Based [AB23, BS22a, BMOTGHL23, CBT^{+}22, CNPP21, CABA^{+}21, CNBH22, DCY^{+}22, DOV^{+}21, FMD^{+}22, FKM^{+}21, FLL23, GC23, GBDS22, GAB^{+}22, HSH^{+}22, HYdM21, HW22, HKMdB22, IBL^{+}17, JSG^{+}22, JSS^{+}18, KwMMM23, KW23, LMC18, LBK22, LPB18, MBP19, MGJ^{+}22, MPB21, MKL^{+}21, MKVS21, MMR22, NFBR23, OCS21, PRR^{+}23, PTD^{+}20, PPG^{+}22, PDR^{+}22, Ryu22b, SG21, SMS^{+}22, SGD19, SJB22, SBTPV22, TCC^{+}18, TLO^{+}22, TLP^{+}23, VCS^{+}23, WCR22, YXES21, ZR20, dCSMVFR23, vAvdGP^{+}22, DB22, LNP^{+}23]. batch [THMZ22]. bath [SGD^{+}22]. battery [DVD17, KYA21, YA19]. battle [Pea21]. BCI [MP20]. bead [ZR20]. beam [FKM^{+}21]. beamline [LKN21]. bearing [DHY18]. beef [LRSC19]. behavior [UMEA19]. behavioral [DHY18]. behaviour [CFRF^{+}23]. behavioural [AdT^{+}20]. Bench [AKJT22, KDH^{+}19, SWPG19]. Bench-top [AKJT22, KDH^{+}19]. benchmark [KFB^{+}23]. benchmarking [THMZ22]. benchtop [DOG21, Poó22]. bespoke [OhMAN23]. between [KHRC21]. beyond [GAMK23]. bi [GCHP22, SKK^{+}23]. bi-directional [GCHP22, SKK^{+}23]. Biaxial [CSS^{+}22, ILG23, PCCS23, STE^{+}20]. Bilateral [TLP^{+}23]. billion [SKS23]. BIMMS [RBRK23]. bin [GC23]. bio [KHRC21, VSTB19]. bio-adhesion [VSTB19]. bio-printer [KHRC21]. biocompatible [KPY^{+}18]. biodiversity [HPS^{+}19]. bioimpedance [MMR22]. BioIn [dOdF23]. BioIn-Tacto [dOdF23]. biological [FES^{+}21, MPB21, RBRK23, SNAO20]. biology [Poó22]. biomedical [MMR^{+}23]. Biomimetic [vdBSRW22, CLCT19]. bioprinter [KTT^{+}21]. bioprinting [BOB^{+}19, TSF21]. bioreactor [LPK^{+}23, RMOB^{+}19]. bioreactors [ELL21]. biosafety [NMZMT20]. BioSamplr [ELL21]. biosensing [KSB22, POT^{+}22]. bipolar [AC20]. birdcage [KTK22]. blenders [PÁBS23]. blood [AAUP23, BDPJGS23]. BLOSM [vAvdGP^{+}22]. blot [BM19]. blow [DOV^{+}21]. board [AP22, KKwM22, ZTS22, Ano17a, Ano17b, Ano18a, NGSU21]. body [FLL23]. bolus [KMG^{+}21]. bone [BBPP23, JCBBFVC23]. Boron [vAvdGP^{+}22]. Boron-based [vAvdGP^{+}22]. BoSL [MSWC21]. bottle [FK19, MBKP23]. Bottom [NGS20]. Bottom-illuminated [NGS20]. boundaries [PWC23]. box [CC18, CHBMCV^{+}23, LDN^{+}22, VCS^{+}23, ARV^{+}22, BMD^{+}20, LLL22]. BREAD [OWHP23]. breadboard [TBBI21]. breathing [HP20, MNMN^{+}20]. bricks [TBBI21]. bridge [PRR^{+}23]. Bringing [PSL^{+}20]. broad [CNBH22]. Broadly [OWHP23]. buffer [WM19]. build [AOAPH19, DMG23, GAB^{+}22]. building [TBBI21]. buildings [PAMC17]. built [GFM^{+}19, PCCS23]. bulk [RAMM20].

Buoy [AB23]. butterfly [KCW21b].

C [BFV^{+}21]. Cable [NKM^{+}21]. cage [WS18b]. calibrating [SKK^{+}23]. calibration [BGE^{+}20, Sur21]. call [CHBMCV^{+}23]. cam [Poó23]. Camera [HOW23, JSG^{+}22, MBC^{+}20, PHL^{+}20, RKD^{+}22, SKC^{+}20, SK23, SG21, SMS^{+}22, SGH^{+}18, WRH^{+}20, ZSLF23]. cameras [BRC^{+}20, PHL^{+}20, SKC^{+}20]. campaigns [BGJGC23]. Can [MDR20]. Canada [KMB^{+}22]. cannula [KLF^{+}20]. canvas [ILG23]. capabilities [CABA^{+}21, KPKK19]. capture [SRB^{+}22]. capturing [SG21]. card [ARV^{+}22]. CAREDAQ [NAC^{+}22]. carrier [CWB^{+}21]. cart [KFB^{+}23, TCC^{+}18]. case [CC18]. cavity [BCN^{+}21]. cEEGrid [KSB22]. cell [CWB^{+}21, KWP^{+}21, KST19, MKM^{+}22, MKL^{+}21, NSZ^{+}23, RHWP22, SZZ22]. cells [FAM^{+}18, PRR^{+}23, PK21, RMOB^{+}19, SGD^{+}22]. cellular [SGD19]. cellulose [MD22]. centrifugal [GLG^{+}23, RKD^{+}22, SMW^{+}20]. centrifuge [BJL^{+}20, Poó22]. ceramic [PWC23]. cerebral [BMOTGHL23, RMOB^{+}19]. chamber [HG21, LS20a, MKM^{+}22, SDDL20, VS22]. Change [CAR^{+}23, HJvdG^{+}22, NAN19]. channel [DB22, FLL23, JTD20, LBK22, REP19, hTFC^{+}23]. chaos [KFB^{+}23]. characterization [AFK^{+}19, CSS^{+}22, DVD17, IPR22, KWP22, RBRK23, SHS^{+}23, SKS23, VSTB19]. characterizing [LS20b]. chatter [SL22]. chemical [HPU21, LOSM18, MBH^{+}21, NIKK21, POT^{+}22, VBK^{+}22]. children [AOAPH19, Lup17]. chip [BRC^{+}20, CWB^{+}21, JCBBFVC23, RGSGD^{+}23]. choice [FK19]. chromatography [MCLSN19, MLBR^{+}23, SS23]. chuck [SOB^{+}23]. cigarette [CSB21]. circuitry [KFH22]. circuits [HPLC21, Kur21]. circularity [PDMV23]. circulatory [SRC^{+}22]. circumferential [GCHP22]. cities [JC17]. Citizen [CGB^{+}19, FKM^{+}21]. citizen-science-based [FKM^{+}21]. clamp [LHN^{+}23, SMSR21]. clamping [SSL^{+}20]. classification [FMD^{+}22, MBML21]. cleanroom [NMZMT20, RGSGD^{+}23]. Climate [CAR^{+}23, BVMHP22c, HJvdG^{+}22, NAN19]. clinical [SRC^{+}22]. clockwork [PPHP^{+}23]. closed [BJL^{+}20, REP19]. closed-loop [BJL^{+}20, REP19]. clusters [THMZ22]. CNN [FMD^{+}22]. CNN-based [FMD^{+}22]. CO [BH22, BGE^{+}20, LKL^{+}21, MKM^{+}23, TLO^{+}22]. coaching [KNAR22]. coastal [CFV^{+}19, RBB^{+}22, SMB^{+}21]. coater [RWK22, SOB^{+}23]. coating [DFS^{+}22]. coefficients [RHWP22]. cold [MYG^{+}23]. collecting [CSB21]. collection [GC23, GCHP22, GCK^{+}23, MTPK20, Pri19]. collector [MLBR^{+}23, FATM22]. Collector/MALDI [FATM22]. colony [PPAP21]. color [NSZ^{+}23, ZTS22]. colorimeter [Kur21]. colorimetric [VCS^{+}23]. ColoriSens [ZTS22]. column [DCY^{+}22, SS23]. combination [JBG^{+}22]. commercial [HSS22, KPY^{+}18, KLF^{+}20]. common [CLCT19]. communication [ARV^{+}22, BVMH21, BVMHP22c, BVBTMVP23, SGD19, ZTS22]. communications [FAR22]. community [MYG^{+}23]. commutator [KK23]. Compact [AB23, CAR^{+}23, BDPJGS23, HYdM21, ICY21, LDN^{+}22, PMA^{+}23]. Comparison [MBH^{+}21]. compatible [STE^{+}20]. compensation [GLPM19]. Compensator [GLPM19]. complex [JC17]. compliance [KCW21b]. compliant [dOdF23]. components [CKB^{+}22, MCAR^{+}21]. composite [DBP^{+}21]. composition [FLL23]. compounds [VOS22]. Comprehensive [MZW^{+}21]. compression [LPK^{+}23]. computational [JC17]. computer [LRSC19]. computer-controlled [LRSC19]. condensate [CSB21]. condition [LMC18]. conditions [RBB^{+}22]. conduct [CHBMCV^{+}23]. conductive [WOM^{+}22]. conductivity [BGJGC23]. connected [KKV20]. connectivity [MZW^{+}21]. constructed [MSWC21]. constructing [PK23]. construction [AZD19, ILL18, JBG^{+}22]. consumption [GBDS22]. contact [HSS22, MCAR^{+}21, SWPG19]. contagion [TLO^{+}22]. contained [HP20, LS20a]. container [MN22]. container-grown [MN22]. contingency [BSE20]. Continuous [SFWW22, DAS22, ICBL22, JKKK22, NND^{+}22, RNRP23]. Control [LRP^{+}23, ARV^{+}22, AHUS19, AG23, BJL^{+}20, BPM^{+}18, CPLM23, CWB^{+}21, DMC22, DKH^{+}23, FZZ^{+}23, HJvdG^{+}22, KFB^{+}23, KCW21a, KW23, MPB21, MSR^{+}23, MMR^{+}23, OFP^{+}23, PHLM22, PMCL20, SL22, SDDL20, TTA^{+}21, TLP^{+}23, WS19]. controllable [PK21, PÁBS23]. controlled [AAUP23, CPB^{+}22, CPCQY23, DC23, DMA^{+}23, GCCCBF22, LRSC19, RHR19, SOB^{+}23, SK21, TLO^{+}22, VdB22, YLW^{+}23, DVD17]. controlled-adsorption [RHR19]. Controller [WS18a, AC20, FZZ^{+}23, GCM21, KCW21b, LAW^{+}22, MZW^{+}21, VKM^{+}23, ZCB^{+}20, OFP^{+}23]. controlling [LKN21]. controls [LS20a]. Conversion [HP20, BOB^{+}19]. converter [BCB^{+}22, CPLM23, MMR22]. converting [WMPP18]. conveying [DMG23]. cooled [BEES22]. cooling [BPP^{+}19, MDC^{+}19, RHWP22]. cope [MDR20]. copter [YM23]. CoralCam [GFTD20]. cores [SMS^{+}22]. coring [THL^{+}20]. corneocyte [LAW^{+}22]. Cost [AB23, HSH^{+}22, KLS22, AIB21, AdT^{+}20, ARLS22, AHUS19, ACHB19, AKJT22, BHF^{+}22, BBJ21, BPW18, BV18, BSD^{+}22, BS22a, BPP^{+}19, BMD^{+}20, BGJGC23, BKM^{+}21, BCB^{+}22, BVVAD21, BVVA21, BVMHP22b, BVMHP22c, BVVAGMRP22, BVORMH^{+}22, BVBTMVP23, BG21, BGE^{+}20, BPSO23, CJ20, CSB21, CBTGAG^{+}18, CLCT19, CKB^{+}22, CPR^{+}21, CRLF^{+}23, CHBMCV^{+}23, CPCQY23, DC23, DYL^{+}18, DAS22, DCY^{+}22, DOV^{+}21, DMS^{+}22, DFS^{+}22, ELL21, FLL23, FRJ^{+}22, FPC^{+}19, GLD18, GCKF22, GML^{+}20, GFTD20, GV20, GBDS22, GFM^{+}19, HS21, HSS22, HCS20, HG21, HPS^{+}19, HKMdB22, ICBL22, IBL^{+}17, IPR22, JZS^{+}23, JSG^{+}22, JCFK^{+}23, JBG^{+}22, JKKK22, JHA^{+}19, JTD20, KwMMM23, KWP^{+}21, KK23, KPY^{+}18, KLF^{+}20, KFH22, KDH^{+}19, KG22, KS21, KHRC21, KMG^{+}21, LZV^{+}20, LSSA20, LNP^{+}23, LKL^{+}21, LWS^{+}22, LAW^{+}22, MLG^{+}20, MKM^{+}22, MWG^{+}22, MGJ^{+}22]. cost [MBML21, MSWC21, MPFC21, MPB21, MHBVBVMA23, MCAR^{+}21, MKVS21, MBC^{+}20, MMR22, MMR^{+}23, MBH^{+}21, MTPK20, NRK^{+}23, NGSU21, NND^{+}22, OP18, OhMA22, OhMAN23, PPAP21, PPF^{+}22, PMA^{+}23, PMCL20, PPHP^{+}23, Poó22, PDDT23, RKD^{+}22, RGSGD^{+}23, RJN^{+}23, Ryu22a, SVMV20, SDDL20, SMB^{+}21, SK23, SS23, SHS^{+}23, SMS^{+}22, SWPG19, SKG22, SKS23, SRC^{+}22, SGD19, Sur21, hTFC^{+}23, TKHB21, TLŠ^{+}22, VS22, VDV23, WSI^{+}22, WSB^{+}20, WDW21, YXES21, ZTS22, ZSLF23, ZBOB22, vAvdGP^{+}22]. Cost-effective [HSH^{+}22, AKJT22, SMS^{+}22]. cost-efficient [SMB^{+}21, vAvdGP^{+}22]. coupled [KCW21b, VCS^{+}23]. courses [BSE20]. COVENT [AIB21]. COVENT-Tester [AIB21]. COVID [CABA^{+}21, GPO^{+}20, HP20, MNMN^{+}20, MDR20, NMZMT20, NHAH^{+}21, Pea21, RRS^{+}23, VA21]. COVID-19 [CABA^{+}21, GPO^{+}20, HP20, MNMN^{+}20, MDR20, NMZMT20, NHAH^{+}21, Pea21, RRS^{+}23, VA21]. COVOX [RRS^{+}23]. creativity [AOAPH19]. creativity-stimulating [AOAPH19]. crisis [CABA^{+}21, MDR20]. CRoAK [GMB20]. crop [SGH^{+}18, SMSR21]. CropManage [SGD19]. crops [MN22]. cryoanesthesia [JCFK^{+}23]. cryopreservation [KOBT23]. cryostats [MDC^{+}19]. crystal [HKMdB22, MMR^{+}23]. CubeSat [LLO22]. cultivation [HW22, NGS20]. Cultural [ASME18]. culture [HW22, MKM^{+}22, MKL^{+}21, NND^{+}22, PK21, WDW21]. CultureLED [HW22]. curing [DS21, GPO^{+}20]. current [BS22b, Ulr19]. curve [CBT^{+}22]. custom [AC20, BSO^{+}23, PCCS23]. custom-built [PCCS23]. Customizable [GMB20, CVT^{+}21, KLF^{+}20, TCG^{+}21, WRH^{+}20]. cut [KST19, Kur21, WDW21]. cycling [LPK^{+}23]. cylindrical [FAM^{+}18].

D

[ASIM21, AHUS19, BS21, BHF^{+}22, BOB^{+}19, BEES22, BVMHP22a, BM18, CVT^{+}21, CKB^{+}22, DS21, DÖ23, DBP^{+}21, DKM23, DOG21, GPO^{+}20, GCM21, HS21, HSS22, HW22, HHSP18, JKKK22, KLS22, KTK22, KTT^{+}21, MWG^{+}22, MSWC21, MKL^{+}21, NSHA22, NIKK21, OKH22, PTKG21, PWC23, PDR^{+}22, PHF18, RJL^{+}23, RWK22, SMS^{+}23, SK23, SSL^{+}20, STE^{+}20, STP20, SKK^{+}23, SFWW22, TSF21, TPLT22, TBBI21, TLP^{+}23, VBK^{+}22, VCS^{+}23, WS18b, WBD^{+}21, WMPP18, YXES21]. D-bioprinter [KTT^{+}21]. D-Printable [WBD^{+}21, BHF^{+}22, TBBI21]. D-printed [BM18, CVT^{+}21, CKB^{+}22, DKM23, GCM21, JKKK22, KLS22, MSWC21, OKH22]. D4 [Kod18]. dams [BBJ21]. DAQ [BBJ21, BCB^{+}22, GML^{+}20, LBK22]. Dark [PZF23]. Dark-field [PZF23]. Data [NAC^{+}22, WRSU23, BBJ21, BCB^{+}22, CC18, CFD^{+}20, GCHP22, GCK^{+}23, KwMMM23, LBK22, MTPK20, NGSU21, OWHP23, PBD^{+}22, Pri19, RGM20]. database [KMB^{+}22]. datalogger [BVMHBV23, McM17, SGD19]. DC [CKMW21, Ulr19]. December [Ano23a]. dedicated [CFD^{+}20]. deep [RPHS23]. defibrillator [FDD17]. degree [UB20]. degrees [FZZ^{+}23]. dehydriding [CBTGAG^{+}18]. delivery [PHLM22]. demineralized [JCBBFVC23]. dendrometer [CHG^{+}21]. density [NND^{+}22, SMB^{+}21]. deployable [SSM^{+}22, ZBOB22]. deployed [PEJ^{+}22]. deployment [NKM^{+}21]. deposition [GAB^{+}22, NIKK21]. Design [BSE20, BVMHP22a, BVMHP22b, CJ20, DFS^{+}22, HKMdB22, ILL18, JBG^{+}22, KHRC21, LRP^{+}23, NSHA22, PRR^{+}23, PPF^{+}22, RKD^{+}22, SBT20, SG21, BEES22, GPO^{+}20, IPR22, KWP22, KSB22, LLL22, MBH^{+}21, RMOB^{+}19, SOB^{+}23, vdBSRW22]. designed [BSO^{+}23]. Designing [OhMA22]. desktop [PHF18, SFWW22]. destructive [SMSR21]. Detecting [GME^{+}22]. detection [CKMW21, DMS^{+}22, IBL^{+}17, JKKK22, POT^{+}22, VCS^{+}23, ZTS22]. detector [FMD^{+}22, HYdM21, MKVS21, VDV23, vAvdGP^{+}22]. detectors [CSH18]. determine [PNM^{+}23, SK23]. determining [RHWP22]. developing [BVMHBV21]. Development [Bie23, BVMHBV23, DÖ23, FAM^{+}18, GBDS22, KW23, LPK^{+}23, MLG^{+}20, MN22, VOS22, HS21, OCS21, YXES21]. Device [LSD22, OhMA22, ARLS22, ASME18, BV18, BFV^{+}21, BSD^{+}22, BS22a, BGJGC23, BVVAD21, BVVA21, BMOTGHL23, BPSO23, CSB21, DAS22, DOV^{+}21, DHY18, DMS^{+}22, EZ22, GCHP22, GCK^{+}23, HJvdG^{+}22, HPS^{+}19, JCFK^{+}23, KDH^{+}19, KKV20, KNAR22, LKL^{+}21, LBK^{+}21, MKH^{+}20, NRK^{+}23, NAC^{+}22, OWHP23, OFP^{+}23, POT^{+}22, RGSGD^{+}23, SBT20, TPLT22, THMZ22, VCS^{+}23, WNKP22]. Devices [PSL^{+}20, AP22, BPM^{+}18, GST^{+}22, PAMC17, SAB^{+}18]. diafiltration [LBK^{+}21]. diaphragm [LKN21]. didactic [PMCL20]. dielectric [PMA^{+}23]. different [HJvdG^{+}22]. differential [BHF^{+}22, LBK22]. diffraction [SE23]. diffuser [PTKG21]. diffusion [SZZ22]. Digital [PSL^{+}20, SKC^{+}20, BDPJGS23, HYdM21, LDN^{+}22, SBTPV22]. dimensional [MMR22]. dip [DFS^{+}22]. direct [KA20, SMS^{+}23]. directional [GCHP22, JBG^{+}22, SKK^{+}23]. Disc [SIY^{+}23]. diseases [FMD^{+}22]. disinfection [BV18, BFV^{+}21]. dispenser [HS21, LNP^{+}23]. dispensing [KLS22]. disruptions [OGP22]. dissipation [MMR^{+}23]. dissolved [NND^{+}22]. distensions [PHLM22]. Distributed [LKJ^{+}23, PBD^{+}22, TLO^{+}22, BSD^{+}22, CPR^{+}21, FZZ^{+}23, KDH^{+}19]. distributing [JZS^{+}23]. distribution [MMR22]. division [MBP19]. DIY [MP20, PJS22, TKHB21]. DNA [PJS22]. do [TKHB21]. do-it-yourself [TKHB21]. docking [SSM^{+}22]. documentation [Ban23]. DoF [TCG^{+}21]. domain [PMA^{+}23]. DONALD [PAMH18]. dongle [BH22]. Doppler [JKKK22]. dorsiflexion [RHAB22]. double [BPSO23, MTPK20]. down [BCN^{+}21, NKS22]. drill [BBPP23, THL^{+}20]. DripOMeter [VAG^{+}22]. drive [KA20]. driven [GLD18, LKN21, RHR19]. Driver [KCW21b, BPP^{+}19, BS22b]. Drone [CAR^{+}23, LHZ^{+}18, PEJ^{+}22]. drone-deployed [PEJ^{+}22]. Drosophila [CLSZ^{+}17]. drug [SZZ22]. drum [RJN^{+}23]. dry [LRSC19, LS20b]. dryer [RAMM20]. drying [RAMM20]. DSAIL [KKwM22]. dual [CPVCRLS17, DC23, DBP^{+}21, FKM^{+}21, GLD18, POT^{+}22]. dual-beam [FKM^{+}21]. dual-loop [CPVCRLS17]. dual-polarity [DC23]. dummy [ACHB19]. duration [PDDT23, YM23]. during [BSE20, DMC22, CABA^{+}21, OGP22, RRS^{+}23, TSF21, ZCB^{+}20]. dynamic [KPY^{+}18].

e-RFIDuino [CNPP21]. ear [KSB22].

ear-electrode [KSB22]. Earth [BVMHBV23]. earthquake [SAB^{+}18]. easily [BHZ^{+}22, MSWC21]. Eastern [CC18]. Easy [DMG23, GAB^{+}22, JTD20, Rom21]. EC [BGJGC23]. ECAM [CSB21]. ecological [GFTD20, MMP^{+}19, WRH^{+}20, ZSLF23]. Economic [Pea20, PWC23]. ecosystems [NLYB23]. Editorial [Ano17a, Ano17b, Ano18a]. eDNA [FES^{+}21]. education [AG23, BVMHBV21, dOFdJN^{+}22, RSV^{+}23, SAB^{+}18]. educational [PDMV23]. EEG [BMD^{+}20, MP20]. effective [AKJT22, HSH^{+}22, IBL^{+}17, OhMAN23, SMS^{+}22]. effector [TCG^{+}21]. effects [RGM20]. efficiency [OKH22]. efficient [GCM21, HSH^{+}22, IBL^{+}17, PAMC17, SMB^{+}21, vAvdGP^{+}22]. egg [KOBT23]. eGreenhouse [LKL^{+}21]. elbow [KA20]. electric [JBG^{+}22]. electric-field [JBG^{+}22]. electrical [BGJGC23, CFD^{+}20, GBDS22, RBRK23, RHAB22, REP19, BHZ^{+}22]. electro [KK23, SSM^{+}22]. electro- [KK23]. electro-permanent [SSM^{+}22]. electrochemical [GFM^{+}19, HKMdB22]. electrochemistry [ICY21]. electrode [KSB22, RBRK23]. electrode-tissue [RBRK23]. electroencephalographic [CPCQY23]. electroformation [BPW18]. Electromagnetic [WSL^{+}19, KAB^{+}20]. Electromechanical [WOM^{+}22, SMSR21]. electromyography [FPC^{+}19, MPFC21]. electron [VDV23]. Electronic [BRC^{+}20, CPLM23, CSB21, KFH22, KNAR22, MCAR^{+}21, VAG^{+}22, NAN19]. electronics [FAR22, LLL22]. electrophoresis [LSSA20]. electrophysiological [AdT^{+}20]. Electrospinning [OhMA22, OhMAN23, WSI^{+}22]. electrostatic [KWP22]. electrowritting [MECH21]. elemental [CED20]. embedded [MBP19, SFWW22]. emergency [HSH^{+}22, KMB^{+}22, RPHS23, RRS^{+}23].

EmerSense [PDDT23]. emersion

[PDDT23]. EMOTIV [MP20]. emulation [SL22]. enable [SGD19]. enabling [CPB^{+}22]. enclosure [GMB20, WS18b]. end [SBTPV22, TCG^{+}21, Bow23]. end-to-end [SBTPV22]. Ender3 [BS21]. endoscope [SMS^{+}22]. Enerduino [PLP23]. Enerduino-pro [PLP23]. energy [BCB^{+}22, BVMHP22c, FAR22, GBDS22, PNM^{+}23, PAMC17, YA19]. energy-efficient [PAMC17]. engineered [LPK^{+}23]. engineering [BG21, SAB^{+}18, SKK^{+}23]. Enhancing [NKS22]. environment [HPS^{+}19, NGSU21]. environmental [BVMHP22a, Bow23, CGB^{+}19, CNPP21, LS20a, LLO22, PJS22, TS22]. environments [ASME18]. equipment [AKJT22, BPW18, LSSA20, NHAH^{+}21, SVMV20, WOM^{+}22]. ergonomic [VA21]. Erratum [Ano21b]. ESB [BMD^{+}20]. ESP32 [Kod18]. ESP32-PICO-D4 [Kod18]. estimation [BRC^{+}20, MMR22]. evaluate [SMSR21]. evaluation [FAM^{+}18, KMB^{+}22, OKH22, Sur21]. events [PDDT23]. everyone [VdB22]. Example [PEJ^{+}22]. exchangeable [DÖ23]. excitation [MBML21]. exoskeleton [KA20, TLP^{+}23]. Expandable [OWHP23]. experimental [DMG23, KFB^{+}23, UMEA19]. experimentation [CGB^{+}19]. experiments [AdT^{+}20, ACHB19, BHF^{+}22, BCN^{+}21, CHBMCV^{+}23, GML^{+}20, HG21, LLL22, MZW^{+}21, OHF17, OFP^{+}23, TTA^{+}21, VdB22, VS22]. expressive [FSM^{+}21]. extended [PPAP21, SGG23]. extension [KST19]. external [FDD17]. extraction [PJS22]. Extruder [BOB^{+}19, DS21, dOFdJN^{+}22, KPKK19, PWC23, PHF18, SFWW22, TSF21, WMPP18]. extrusion [KHRC21, TSF21]. extrusions [TBBI21]. eye [CLSZ^{+}17]. fabric [SWPG19]. fabricate [JTD20]. Fabricating [OhMAN23]. Fabrication [MPB^{+}23, MD22, OKH22, OhMA22, DFS^{+}22, FRJ^{+}22, JCBBFVC23, NRK^{+}23, RNRP23, RKD^{+}22, SBT20]. face [FSM^{+}21, IBL^{+}17, NHAH^{+}21]. Facially [FSM^{+}21]. facility [LMB^{+}23]. factors [MMP^{+}19]. FAIMS [VOS22]. FAL [MSWC21]. fan [LS20a]. fast [BPP^{+}19, CAV^{+}22, LZV^{+}20, MPASIM23, RHR19, JTD20]. fast-scan [RHR19]. feedback [KA20]. feeder [JZS^{+}23, OHF17]. feeding [RNRP23]. fever [ASIM21]. FFF [LSD22]. fiber [AAUP23, BVORMH^{+}22, KCW21b, MKH^{+}20, SFWW22]. fiber-coupled [KCW21b]. fibers [DOV^{+}21, SBT20]. fibre [FAM^{+}18]. field [BVMHBV23, BGE^{+}20, DBRE21, HJvdG^{+}22, JBG^{+}22, LSSA20, LDN^{+}22, NSHA22, PJS22, PZF23, Sur21, SGH^{+}18, TCC^{+}18, VOS22, WSL^{+}19]. field-based [TCC^{+}18]. fighter [HP20]. filament [dOFdJN^{+}22, PWC23, WMPP18]. filaments [PWC23]. filled [MWG^{+}22]. film [DVD17, GAB^{+}22]. filtering [ILL18]. filtration [OKH22]. fire [HP20, RPHS23, SKK^{+}23]. First [GCKF22, vAvdGP^{+}22]. FISH [DB22, HCS20, UB20, vdBSRW22]. fish-robot [UB20]. FishCam [MBC^{+}20]. five [UB20]. fixation [WYT23]. flexible [CPLM23, CFRF^{+}23, FRJ^{+}22, GFTD20, KNAR22, OCS21, RC18, Sur21]. flight [BRC^{+}20, Pri19, YM23]. flood [SSM^{+}22]. Flow [GME^{+}22, BPSO23, GCHP22, GCK^{+}23, HS21, HPC21, ICBL22, KPY^{+}18, LBK^{+}21, MHBVBVMA23, MD20]. flow-through [MD20]. fluid [BKM^{+}21, DMG23, GHHR20, SRC^{+}22]. fluidics [BKM^{+}21, DB22]. fluorescence [Ban23, MPASIM23, NRK^{+}23, POT^{+}22, STE^{+}20, ZTS22]. fluorescent [DB22, HPU21]. fluorometer [BHF^{+}22]. flux [BSD^{+}22]. Fly [BJL^{+}20]. Fly-by-Pi [BJL^{+}20]. fNIRS [TKHB21]. foam [WCR22]. FoamPi [WCR22]. focus [PZF23]. food [DÖ23, JZS^{+}23, JSS^{+}18]. ForageFeeder [JZS^{+}23]. force [CNBH22, GFLMRV^{+}23, HWR^{+}19, LAW^{+}22, hTFC^{+}23]. forest [RPHS23]. forming [SKG22, WNKP22]. Fostering [PDMV23]. found [OGP22]. four [FLL23]. FPGA [DKH^{+}23, IBL^{+}17, WM19]. fraction [FATM22, MLBR^{+}23]. frame [WM19]. framework [KMB^{+}22, MSR^{+}23, OWHP23, OCS21]. Free [BDPJGS23, Pea20]. freedom [FZZ^{+}23, UB20]. freely [KK23]. freeze [BPSO23]. freeze-quench [BPSO23]. French [TPLT22]. frequency [BPP^{+}19, SKS23]. FRESH [BOB^{+}19, TSF21]. FRESH-method [BOB^{+}19]. freshwater [NLYB23]. friction [CHG^{+}21]. front [Bow23]. front-End [Bow23]. frontline [VA21]. fronts [PEJ^{+}22]. fruity [BSS^{+}20]. FTP [SGD19]. full [NHAH^{+}21]. full-face [NHAH^{+}21]. Fully [VSTB19, AG23]. functional [TKHB21]. functions [PDR^{+}22]. fused [RNRP23].

gain [CKMW21]. galvanostat [DVD17, ICY21]. gamma [CSH18]. gamma-ray [CSH18]. gantry [TTA^{+}21]. gas [BCN^{+}21, KFH22, MKM^{+}23, Sur21]. Gaseous [Bow23]. gastropod [CLCT19]. Gateway [PSL^{+}20]. GatorByte [AB23]. gauge [NKS22]. gel [KHRC21, LSSA20]. General [GLPM19, ICY21]. General-compensation-purpose [GLPM19]. generating [CSB21]. generation [RMOB^{+}19]. generator [ZR20]. gentle [OGP22]. geolocation [BG21]. geological [SK21]. geopositioning [BVMHP22b, BVMHBV23]. geotechnical [BJL^{+}20]. germanium [CSH18]. germicidal [ARLS22]. Giant [BPW18]. glacial [RBB^{+}22]. glacier [PEJ^{+}22]. glaciers [NAN19]. GNOME [Bow23].

Gordian [KOBT23]. GPS [BGJGC23, CC18]. gradients [SMB^{+}21]. grafts [JCBBFVC23]. granulator [RJN^{+}23]. Grape [GCKF22]. graphite [MD22]. gravity [DCY^{+}22, SOB^{+}23]. greenhouse [LKL^{+}21]. Greenland [PEJ^{+}22]. grid [KKwM22, MRKP23, RBB^{+}22]. Griggs [SK21]. ground [JSG^{+}22, NKM^{+}21, NKS22]. ground-based [JSG^{+}22]. grown [MN22]. growth [HG21, MZW^{+}21]. Guide [AOAPH19, PDMV23]. guided [SE23].

hand [CPCQY23, PK20, TLP^{+}23, YXES21, ZBOB22]. handheld [BS22a, WLML23]. handling [GHHR20, HG21, WLML23]. Hangprinter [RNRP23]. haptic [CPCQY23, KA20]. Hardware [MDR20, SL22, AdT^{+}20, ASME18, BVMHBV21, BHZ^{+}22, CC18, CKMW21, GCM21, HF23, HHKB20, KWP22, KCW21a, KCW21b, KOBT23, KW23, MYG^{+}23, OGP22, OCS21, PHLM22, REP19, THMZ22]. Hardware-in-the-loop [SL22]. hardware-software [THMZ22]. HardwareX [Pea21]. hare [KOBT23]. harmful [JBG^{+}22]. hazardous [PEJ^{+}22]. head [ACHB19, GCCCBF22, OKH22, WYT23]. head-fixation [WYT23]. headband [TKHB21]. headset [MP20]. health [MDR20, RRS^{+}23]. heart [SJB22]. heat [LS20b, SWPG19, STP20]. heat-sterilizable [STP20]. heatwave [HJvdG^{+}22]. heavy [CSH18]. hectometric [DMA^{+}23]. helium [MDC^{+}19]. helmet [LPB18]. helmet-mounted [LPB18]. help [MDR20]. Heritage [ASME18]. High [BEES22, CHG^{+}21, LLO22, RAMM20, BCB^{+}22, BS22b, CSH18, DC23, DHY18, GFM^{+}19, JC17, KCW21b, LMB^{+}23, LWS^{+}22, LAW^{+}22, LPB18, MPB21, MZLG20, MKM^{+}23, MTPK20, NKS22, Pri19, SIR18, STP20, SK21, TSF21, TLŠ^{+}22, VS22, vdBSRW22]. high-bandwidth [LPB18]. high-current [BS22b]. High-performance [BEES22]. high-performing [MPB21]. high-precision [GFM^{+}19, LWS^{+}22, NKS22]. high-purity [CSH18, MKM^{+}23]. high-quality [MTPK20]. high-speed [LAW^{+}22]. high-temperature [STP20]. high-throughput [LMB^{+}23, TLŠ^{+}22]. Highly [CVT^{+}21, OKH22]. Highly-customizable [CVT^{+}21]. HistoEnder [PDR^{+}22]. histological [PDR^{+}22]. hobby [KA20]. holders [SS23]. hollow [AAUP23]. hollow-fiber [AAUP23]. Holographic [BDPJGS23]. honeycomb [TBBI21]. horn [JKKK22]. horticultural [Bie23]. hot [GLG^{+}23]. house [DHY18, WSI^{+}22]. housing [GML^{+}20]. HPGe [CSH18]. HRI [PK20]. human [JKKK22, OKH22, RMOB^{+}19]. humanoid [FSM^{+}21]. Humidity [VdB22, LS20a, RAMM20]. Humidity-controlled [VdB22]. HumidOSH [LS20a]. hybrid [KTT^{+}21, MP20]. hybridization [DB22]. hydraulic [FZZ^{+}23]. hydriding [CBTGAG^{+}18]. hydriding/dehydriding [CBTGAG^{+}18]. hydrogel [SFWW22, ZR20]. hydrogels [DBP^{+}21, KTT^{+}21]. HyperRail [AHUS19]. Hyperspectral [SVMV20, BVVA21]. Hypnos [NGSU21].

I-V [CBT^{+}22]. I/O [PBD^{+}22]. i3

[BOB^{+}19]. ice

[MYG^{+}23, MZLG20, THL^{+}20]. iceberg [THL^{+}20]. IETeasy [MBML21]. iFO [HCS20]. illuminated [NGS20]. illumination [HW22, KDH^{+}19, TLŠ^{+}22]. image [FMD^{+}22, MSR^{+}23, SE23]. image-guided [SE23]. Imaging [SVMV20, AHUS19, BDPJGS23, CWB^{+}21, CFD^{+}20, DB22, KTK22, KAB^{+}20, KMG^{+}21, LAW^{+}22, LB17, LHZ^{+}18, LDN^{+}22, MKL^{+}21, MPB^{+}23, NSZ^{+}23, PPG^{+}22, RKD^{+}22, VCS^{+}23, ZCB^{+}20]. immersion [SGD^{+}22]. immunoassays [WLML23]. impact [ACHB19]. impacts [ARV^{+}22, NAN19, WSB^{+}20]. Impedance [BHZ^{+}22, HKMdB22, MMR22]. implant [WYT23]. implant-protection [WYT23]. implantable [KLF^{+}20]. implants [LPK^{+}23]. implement [Rom21]. Implementation [RC18, BVMHP22a, CJ20, IBL^{+}17, KHRC21, NSHA22, WM19]. implemented [DMS^{+}22]. impulse [MBML21]. IMU [BVMHBV23, LMC18, SG21, TLP^{+}23]. IMU-based [LMC18]. in-house [DHY18, WSI^{+}22]. in-lab [THMZ22]. In-situ [FAR22, FAM^{+}18, KAB^{+}20]. in-vitro [LPK^{+}23]. in-vivo [KMG^{+}21]. incidence [BVORMH^{+}22]. incident [BVVAGMRP22]. incline [WSB^{+}20]. including [CWB^{+}21]. incubation [BCN^{+}21, PPAP21]. incubator [CRLF^{+}23, DBRE21, MKL^{+}21, WDW21]. incubot [MKL^{+}21]. indoor [ASME18, BVVA21, GV20]. indoor/outdoor [BVVA21]. induced [KS21, NKS22]. industrial [TBBI21]. Industry [PBD^{+}22]. Inertial [BVMHBV23, LPB18, RGSGD^{+}23, SG21]. inexpensive [DBRE21, DHY18, NLYB23, TTA^{+}21]. infections [NMZMT20]. Infrared [ASIM21, MHBVBVMA23, HCS20, TKHB21]. infusion [KMG^{+}21, SRC^{+}22, VAG^{+}22]. InGaAs [MKVS21]. initiatives [MDR20]. inline [GML^{+}20, RCM21]. Inquiry [LOSM18]. insertion [BBPP23]. Instrument [GME^{+}22, MBML21, MRZMT21]. instrumentation [DMA^{+}23, FAM^{+}18, MBH^{+}21, RHR19]. Instrumented [dOFdJN^{+}22, BBPP23, GMB20]. insulin [PPF^{+}22]. integer [MBP19]. Integrated [OhMA22, KOBT23, NLYB23, TPLT22]. Integrating [PSL^{+}20]. integration [MCLSN19, WS18a, ZTS22]. intelligent [FAR22]. intensity [KDH^{+}19]. intensive [BSD^{+}22]. interaction [UB20]. interface [INS^{+}23, KW23, RBRK23]. internet [MZW^{+}21, AB23, TS22]. intertidal [PDDT23]. intra [BMOTGHL23]. intra-cerebral [BMOTGHL23]. intracerebral [KLF^{+}20]. intravenous [VAG^{+}22]. invasive [DAS22]. inverse [ILL18]. Inverter [LRP^{+}23, PRR^{+}23]. investigate [MRZMT21]. investigating [UMEA19, UB20]. ion [FAM^{+}18, RC18, RHWP22, SGD^{+}22, VOS22]. Ionic [GAB^{+}22]. IOT [SJB22, BSE20, BVVAD21, BVVA21, BVVAGMRP22, GC23, GBDS22, KKV20, LMB^{+}23, PAMC17, PSL^{+}20, WRSU23]. IoT-based [GC23]. iris [LKN21]. iron [KYA21, YA19]. irradiation [ARLS22]. irrigation [MN22, SGD19]. isocratic [MLBR^{+}23]. isolated [Ulr19]. isolation [VS22]. isothermal [SGD^{+}22]. isotopic [CED20]. iThermowall [ASIM21]. IV [VAG^{+}22]. IVRT [SZZ22].

Jones [IBL^{+}17]. June [Ano23b].

K2 [NSZ^{+}23]. keyboard [MBP19]. kg [TCG^{+}21]. killing [JBG^{+}22]. kinematic [BG21]. Kinesis [GMB20]. Kinetic [PNM^{+}23]. Kit [CGB^{+}19, BS21, BSO^{+}23, CVT^{+}21, PDMV23]. kitchen [PÁBS23]. Knot [KOBT23].

Lab [PSL^{+}20, SIY^{+}23, CKMW21, JTD20, KCW21a, KCW21b, LLL22, RGSGD^{+}23, THMZ22, AG23]. lab-on-a-chip [RGSGD^{+}23]. Lab-on-a-Disc [SIY^{+}23]. Label [BDPJGS23]. Label-Free [BDPJGS23]. Laboratory [PSL^{+}20, AG23, BCN^{+}21, CFD^{+}20, DP17, GML^{+}20, HJvdG^{+}22, LDN^{+}22, PÁBS23, RAMM20, TBM20, vAvdGP^{+}22]. Laborem [LLL22]. labs [Poó22]. Lake [LHZ^{+}18]. LAMSkyCam [JSG^{+}22]. landslide [CPR^{+}21]. lapse [LDN^{+}22, WRH^{+}20]. Large [PHF18, SGD^{+}22, DMG23, LHN^{+}23, MD22, PPAP21, RNRP23, WS18a, vAvdGP^{+}22]. large-scale [WS18a, vAvdGP^{+}22]. laser [BPP^{+}19, KST19, Kur21, MLG^{+}20, RGSGD^{+}23, SKS23, WDW21]. laser-ablated [RGSGD^{+}23]. laser-cut [KST19, Kur21, WDW21]. lasers [KCW21b]. lateral [HS21]. Layer [GAB^{+}22, BPM^{+}18, RAMM20]. layouts [PK23]. leaf [FMD^{+}22, MMP^{+}19]. leafcutter [SRB^{+}22]. learning [KFB^{+}23, RPHS23, SAB^{+}18]. learning-powered [RPHS23]. led [PLP23, HW22, MP20, MPB^{+}23, RHR19, SHS^{+}23, ZCB^{+}20]. LED-driven [RHR19]. Legacy [PSL^{+}20]. Lensless [BDPJGS23]. less [ZBOB22]. level [KwMMM23, SKS23]. levels [MKM^{+}22]. Li [FAM^{+}18, RHWP22]. Li-ion [FAM^{+}18, RHWP22]. lickometer [RMB18]. LiDAR [BS22a]. LiDAR-based [BS22a]. life [GLD18]. light [ARLS22, BVBTMVP23, CKMW21, GAMK23, JBG^{+}22, PZF23, RGM20, SHS^{+}23]. Lightweight [PEJ^{+}22, BCB^{+}22]. limb [RHAB22]. limited [RHAB22, VCS^{+}23]. limnic [RBB^{+}22]. linear [AHUS19]. liquid [BEES22, GCM21, KLS22, MCLSN19, MLBR^{+}23, MDC^{+}19, SMS^{+}23, WLML23, ZR20]. lithium [DBRE21, SGD^{+}22]. lithium-ion [SGD^{+}22]. live [CPB^{+}22, MKL^{+}21, NSZ^{+}23, STE^{+}20]. live-cell [NSZ^{+}23]. local [GC23]. lock [HYdM21]. lock-in [HYdM21]. locomotion [WSB^{+}20, vdBSRW22]. locomotor [SRB^{+}22]. lodging [SMSR21, SGH^{+}18]. Logger [MMP^{+}19, CC18, CLCT19, CFD^{+}20, PDDT23, RGM20]. logic [Rom21]. long [CWB^{+}21, HG21, MKL^{+}21, Rom21, SFWW22, YM23]. long-term [CWB^{+}21, HG21, MKL^{+}21, Rom21]. loop [BJL^{+}20, CPVCRLS17, DMA^{+}23, IPR22, KW23, REP19, SL22]. LoRa [KwMMM23, TLO^{+}22]. LoRaWAN [GBDS22]. Low [AB23, AdT^{+}20, BPW18, BBPP23, BPP^{+}19, BGJGC23, BKM^{+}21, BVMHP22c, BVORMH^{+}22, BVBTMVP23, BGE^{+}20, BPSO23, CBTGAG^{+}18, CKMW21, CRLF^{+}23, CHBMCV^{+}23, DC23, DYL^{+}18, DAS22, FRJ^{+}22, FPC^{+}19, GLD18, GML^{+}20, HS21, HSS22, KwMMM23, KFH22, KLS22, KMG^{+}21, LNP^{+}23, LWS^{+}22, MKM^{+}22, MWG^{+}22, MGJ^{+}22, MPFC21, MHBVBVMA23, MTPK20, NRK^{+}23, OhMA22, PPAP21, PPHP^{+}23, RGSGD^{+}23, RJN^{+}23, SS23, SHS^{+}23, SWPG19, SRC^{+}22, hTFC^{+}23, TLŠ^{+}22, VDV23, WSB^{+}20, AIB21, ARLS22, AHUS19, ACHB19, BSO^{+}23, BHF^{+}22, BBJ21, BV18, BSD^{+}22, BS22a, BMD^{+}20, BCB^{+}22, BVVAD21, BVVA21, BVMHP22b, BVVAGMRP22, BG21, CJ20, CSB21, CLCT19, CKB^{+}22, CPR^{+}21, CPCQY23, DCY^{+}22, DOV^{+}21, DMS^{+}22, DFS^{+}22, ELL21, FLL23, GCKF22, GFTD20, GV20, GBDS22, GFM^{+}19, HCS20, HG21, HPS^{+}19, HKMdB22, ICBL22, IPR22, JZS^{+}23, JSG^{+}22, JCFK^{+}23]. low [JBG^{+}22, JKKK22, JC17, JHA^{+}19, JTD20, KWP^{+}21, KK23, KPY^{+}18, KLF^{+}20, KDH^{+}19, KG22, KS21, KHRC21, LZV^{+}20, LSSA20, LKL^{+}21, LAW^{+}22, Lup17, MLG^{+}20, MBML21, MSWC21, MPB21, MCAR^{+}21, MKVS21, MBC^{+}20, MMR22, MMR^{+}23, MBH^{+}21, NGSU21, NND^{+}22, OP18, PPF^{+}22, PMA^{+}23, PMCL20, Poó22, PDDT23, RKD^{+}22, Ryu22a, SVMV20, SDDL20, SMB^{+}21, SK23, SKG22, SKS23, SGD19, Sur21, TKHB21, VS22, WSI^{+}22, WDW21, YXES21, ZTS22, ZSLF23, ZBOB22]. low-anthropomorphic [Lup17]. Low-Cost [AB23, KLS22, AdT^{+}20, BPW18, BGJGC23, BKM^{+}21, BVORMH^{+}22, BVBTMVP23, BPSO23, CBTGAG^{+}18, CRLF^{+}23, CHBMCV^{+}23, DC23, DYL^{+}18, FRJ^{+}22, FPC^{+}19, GLD18, GML^{+}20, HS21, HSS22, KFH22, KMG^{+}21, LNP^{+}23, LWS^{+}22, MKM^{+}22, MGJ^{+}22, MPFC21, MHBVBVMA23, MTPK20, NRK^{+}23, OhMA22, PPAP21, PPHP^{+}23, RGSGD^{+}23, RJN^{+}23, SS23, SHS^{+}23, hTFC^{+}23, TLŠ^{+}22, VDV23, WSB^{+}20, AIB21, ARLS22, AHUS19, ACHB19, BHF^{+}22, BBJ21, BV18, BS22a, BMD^{+}20, BCB^{+}22, BVVAD21, BVVA21, BVMHP22b, BVVAGMRP22, BG21, CSB21, CLCT19, CKB^{+}22, CPR^{+}21, CPCQY23, DCY^{+}22, DOV^{+}21, DMS^{+}22, DFS^{+}22, FLL23, GCKF22, GFTD20, GV20, GBDS22, GFM^{+}19, HCS20, HG21, HPS^{+}19, HKMdB22, ICBL22, IPR22, JZS^{+}23, JSG^{+}22, JCFK^{+}23, JKKK22, JHA^{+}19, JTD20, KWP^{+}21, KK23, KPY^{+}18, KLF^{+}20, KDH^{+}19, KG22, LZV^{+}20, LKL^{+}21, LAW^{+}22, MLG^{+}20, MBML21]. low-cost [MSWC21, MPB21, MCAR^{+}21, MKVS21, MBC^{+}20, MMR22, MBH^{+}21, NGSU21, NND^{+}22, OP18, PMA^{+}23, PMCL20, Poó22, PDDT23, RKD^{+}22, Ryu22a, SVMV20, SDDL20, SK23, SKG22, SKS23, SGD19, Sur21, TKHB21, VS22, WDW21, YXES21, ZTS22, ZSLF23, ZBOB22]. low-light [CKMW21]. low-power [PPHP^{+}23]. low-resource [BSO^{+}23]. Low-speed [BBPP23]. low-weight [SMB^{+}21]. lower [FZZ^{+}23, RHAB22]. LUCIA [BFV^{+}21]. Luminance [DMC22]. luminescence [MKVS21]. LunAero [HHKB20]. Lvital [DAS22]. LWIR [SMS^{+}22]. lysimeter [MN22].

m [AHUS19]. Maasi [RWK22]. MAC [KK23]. machine [AAUP23, DFS^{+}22, MLG^{+}20, OhMAN23, WSI^{+}22]. made [KA20, Poó23]. MADV [LBK22]. MADV-DAQ [LBK22]. magnet [PAMH18, SSM^{+}22]. magnetic [BVMHBV23, BS22b, CHG^{+}21, DOG21, KTK22, NSHA22, SGG23]. makers [PDMV23]. Making [VS22]. MALDI [FATM22]. management [Bie23, CHBMCV^{+}23, KKwM22, NGSU21]. manipulating [PTKG21]. manipulation [GCM21]. mannequin [OKH22]. manual [CABA^{+}21]. manufacturing [LKJ^{+}23, POP^{+}20, RNRP23]. mapper [NSHA22]. mapping [BVMHBV23, MD22, WSL^{+}19]. March [Ano23c]. marine [GV20, NLYB23, PHL^{+}20]. Masi [CABA^{+}21]. mask [HP20, NHAH^{+}21, OKH22, PTD^{+}20]. mask-based [PTD^{+}20]. masks [BFV^{+}21]. mass [SMS^{+}23]. masses [KOBT23]. Mate [PK23]. material [EZ22, PCCS23, SK21]. materials [Bow23, MBML21, NIKK21, PÁBS23, SK23, SHS^{+}23, WOM^{+}22]. matrix [JCBBFVC23]. matter [BSO^{+}23]. measure [KS21, NKS22]. Measurement [BVMHBV23, BVMH21, BVORMH^{+}22, IPR22, KDH^{+}19, KKV20, LBK22, LPB18, MGJ^{+}22, MZLG20, SK23, SG21, SRC^{+}22, VSTB19]. measurements [EZ22, ICY21, LZV^{+}20, MKM^{+}23, PMA^{+}23, PEJ^{+}22, Rom21, REP19, RBB^{+}22, SMB^{+}21, WSL^{+}19]. Measuring [Bow23, BH22, BVVAGMRP22, BVBTMVP23, GMB20, LSD22, NAN19, SWPG19]. Mechanical [MDR20, CABA^{+}21, CSS^{+}22, ILG23, MBML21, NAC^{+}22]. Mechanically [MKH^{+}20]. mechanism [DÖ23, SSM^{+}22]. mechanisms [MBH^{+}21]. mechanized [RJN^{+}23]. mechanochemical [PÁBS23]. mechanoluminescence [EZ22]. mechanotransduction [SBT20]. mediated [PHLM22]. medium [JCBBFVC23]. medium-scale [JCBBFVC23]. MEDUSA [PPG^{+}22]. melt [GLG^{+}23, MECH21]. melting [BHF^{+}22]. membrane [PTKG21]. meniscal [LPK^{+}23]. meshes [OhMAN23]. mesocosms [HJvdG^{+}22]. metal [KAB^{+}20, WNKP22]. metals [MWG^{+}22]. meter [HWR^{+}19, PAMC17, PLP23, TCG^{+}21]. method [BOB^{+}19, DOV^{+}21, GAB^{+}22]. Methods [MDC^{+}19]. MHz [Ulr19]. Miau [Car21]. mice [DHY18, KLF^{+}20, WYT23]. micro [CKB^{+}22, HJvdG^{+}22]. micro- [HJvdG^{+}22]. micro-milling [CKB^{+}22]. microalgae [NGS20, NND^{+}22]. microarrays [PK23]. Microbalance [MMR^{+}23, Car21, HKMdB22]. microbiological [DBRE21]. microbiology [DBRE21]. micrObs [WRH^{+}20]. microcontroller [BGE^{+}20, CRLF^{+}23, MGJ^{+}22]. microcontroller-based [MGJ^{+}22]. microextraction [MCLSN19]. MicroEye [CKB^{+}22]. microfibres [GLG^{+}23]. Microfluidic [VCS^{+}23, BPM^{+}18, CWB^{+}21, GST^{+}22, KCW21a, NRK^{+}23, RGSGD^{+}23, RJL^{+}23, hTFC^{+}23, TLŠ^{+}22, WLML23, WS18a, ZTS22]. microfluidics [GHHR20, RKD^{+}22, WS19, YLW^{+}23]. MicroFPGA [DKH^{+}23]. microgrid [ARV^{+}22]. microgrids [NKM^{+}21]. Microinjection [SNAO20]. microinjector [BMOTGHL23]. micromanipulators [HHSP18]. micrometric [DOV^{+}21]. MicroMI [DBRE21]. microorganisms [JBG^{+}22, LS20b]. microparticle [RGSGD^{+}23]. micropump [BM18]. microscope [AKJT22, BDPJGS23, DKH^{+}23, GFLMRV^{+}23, GFM^{+}19, MPASIM23, MKL^{+}21, MPB^{+}23, NRK^{+}23, NSZ^{+}23, PZF23, VDV23]. Microscopy [VKM^{+}23, GAMK23, HF23, LAW^{+}22, MSR^{+}23, SE23, STE^{+}20]. microsecond [TLŠ^{+}22]. microstructure [RMB18]. microsyringe [CM18]. microUSV [GV20]. Microwave [MHBVBVMA23, WSL^{+}19]. miEye [AKJT22]. migration [HHKB20]. mill [MRKP23]. Millifluidic [LOSM18]. milling [CKB^{+}22]. mini [MN22, Poó22, BCB^{+}22]. Mini-DAQ [BCB^{+}22]. mini-lysimeter [MN22]. miniature [JSG^{+}22, THL^{+}20]. miniaturized [NMZMT20, RMOB^{+}19, SIY^{+}23]. mirror [RHAB22]. mixer [DP17]. mixers [PÁBS23]. mixing [BPSO23]. mobile [BVMHBV21, BVMHP22b, BG21, CNPP21, DBRE21, HSS22, NKM^{+}21, NFBR23, OCS21, RSV^{+}23, RBB^{+}22, THMZ22]. mobility [RC18, VOS22]. modal [dOdF23]. mode [KMG^{+}21, TLP^{+}23]. model [CPB^{+}22]. models [PHLM22]. modem [SGD19]. Modification [PÁBS23, KHRC21]. Modified [NHAH^{+}21, DHY18, SZZ22]. ModLight [GAMK23]. Modular [AHUS19, BVMH21, GAMK23, LRP^{+}23, CJ20, CAR^{+}23, CKB^{+}22, DMG23, EMMH20, PRR^{+}23, PK23, Poó23, SS23, Sur21, TBBI21, WBD^{+}21]. Modulators [BPP^{+}19]. Module [PSL^{+}20, PMCL20, SK23, dOdF23]. moduli [SK23]. moisture [TTA^{+}21, vAvdGP^{+}22]. molecular [Poó22]. molecule [NSZ^{+}23, YLW^{+}23]. monitor [CNPP21, CLCT19, PDDT23, RCM21]. Monitoring [AB23, ASME18, BBJ21, BSD^{+}22, BGJGC23, CGB^{+}19, CFV^{+}19, CKB^{+}22, CPR^{+}21, DAS22, DKM23, FKM^{+}21, FES^{+}21, GFTD20, GBDS22, GCHP22, HPS^{+}19, HPU21, JHA^{+}19, KPKK19, LMC18, LHZ^{+}18, MECH21, MMR^{+}23, NAC^{+}22, NND^{+}22, OP18, POP^{+}20, Ryu22a, Ryu22b, SSM^{+}22, SGD19, Sur21, TLO^{+}22, VAG^{+}22, WCR22, ZBOB22]. mooring [ZBOB22]. Mostly [VBK^{+}22]. motion [AHUS19, UB20]. Motor [LKN21, KA20]. Motor-driven [LKN21]. motorized [KK23, TCC^{+}18]. Motrol [THMZ22]. mount [CPVCRLS17]. mounted [LPB18]. Mouthguard [LPB18]. Mouthguard-based [LPB18]. Movement [TBM20, GCCCBF22, RHAB22, SG21, SGH^{+}18]. movements [GMB20]. moving [KK23]. Multi [AFK^{+}19, LBK22, POP^{+}20, AZD19, BVVAD21, BPM^{+}18, GMB20, HW22, JBG^{+}22, JTD20, KFB^{+}23, OCS21, REP19, SKC^{+}20, YM23, dOdF23]. multi-arm [KFB^{+}23]. Multi-axis [AFK^{+}19, GMB20]. multi-camera [SKC^{+}20]. Multi-channel [LBK22, JTD20, REP19]. multi-copter [YM23]. multi-directional [JBG^{+}22]. multi-layer [BPM^{+}18]. multi-modal [dOdF23]. multi-robot [AZD19]. Multi-sensor [POP^{+}20, OCS21]. multi-spectral [BVVAD21]. multi-well [HW22]. Multichannel [BS22b]. multifunction [KPKK19]. multilayered [SWPG19]. multilevel [PRR^{+}23]. Multimodal [MPASIM23, MPB^{+}23]. multiparameter [PDDT23]. multiple [KDH^{+}19]. multiport [MD20, MDD^{+}21]. Multipurpose [RPHS23]. multispectral [BVMHP22b, PPG^{+}22]. MVs [MDR20]. MYSTAT [ICY21].

N95 [BFV^{+}21]. nanofiber [WSI^{+}22]. Nanofibrous [OhMA22, OhMAN23]. nanopositioner [LWS^{+}22]. nanotextures [LAW^{+}22]. nasopharyngeal [GPO^{+}20]. natural [DMC22]. Naval [CAR^{+}23]. Near [MHBVBVMA23, RBB^{+}22, TKHB21]. Near-Infrared [MHBVBVMA23, TKHB21]. neonatal [CRLF^{+}23, JCFK^{+}23, WSB^{+}20]. Nerita [CLCT19]. nerve [SBT20]. nerve-stretching [SBT20]. network [CPR^{+}21, KS21, TLO^{+}22]. networks [GBDS22]. neuro [REP19]. neuro-prosthesis [REP19]. neurophysiology [CJ20]. neuroscience [BMOTGHL23, TBM20]. neutron [vAvdGP^{+}22]. NIR [BVMH21]. Nitrogen [Bow23]. nocturnal [HHKB20]. noise [BVBTMVP23, CKMW21]. non [DAS22, MCAR^{+}21, SMSR21]. non-contact [MCAR^{+}21]. non-destructive [SMSR21]. non-invasive [DAS22]. NORDACC [CAR^{+}23]. NOSE [BOB^{+}19]. Novel [SZZ22, KAB^{+}20, KNAR22, MBP19, SOB^{+}23]. nozzle [DBP^{+}21]. nuclear [DOG21]. nucleation [MYG^{+}23]. Nydus [BOB^{+}19].

O [PBD^{+}22]. Observation

[HCS20, vAvdGP^{+}22]. observations [ZSLF23]. observatory [RBB^{+}22]. obstacle [CFV^{+}19]. occlusion [GCK^{+}23]. occurrence [PDDT23]. ocean [CFV^{+}19, OFP^{+}23, PEJ^{+}22]. oceanic [RBB^{+}22]. October [Ano17c, Ano18b, Ano19b, Ano20b, Ano21c, Ano22b]. OCTOPUS [HOW23]. odyssey [BSS^{+}20]. off [KKwM22, RBB^{+}22]. off-grid [MRKP23, RBB^{+}22]. offset [CKMW21]. OhmPi [CFD^{+}20]. OMIS [LOSM18]. omnidirectional [ILL18, RSV^{+}23]. on-field [Sur21]. on-site [PMA^{+}23]. One [BOB^{+}19, SIY^{+}23, NGSU21, TCG^{+}21, WS19]. one-meter [TCG^{+}21]. online [BSE20, CKB^{+}22]. Open [BVMHBV21, BEES22, BM18, BM19, BJL^{+}20, CED20, CNBH22, CPB^{+}22, CKMW21, DP17, FZZ^{+}23, FDD17, FATM22, GC23, GPO^{+}20, GCM21, GME^{+}22, HF23, HOW23, JCBBFVC23, KWP22, KMB^{+}22, KLS22, KCW21a, KCW21b, KPKK19, KTT^{+}21, KYA21, KST19, Kur21, LOSM18, LAW^{+}22, MDR20, MBKP23, MMR^{+}23, NAN19, NSZ^{+}23, OP18, OWHP23, PHLM22, PAMC17, PSL^{+}20, POP^{+}20, PHL^{+}20, RSV^{+}23, REP19, SMW^{+}20, SIY^{+}23, STP20, TS22, Ulr19, UMEA19, UB20, WMPP18, WRSU23, YA19, dCSMVFR23, AIB21, ASIM21, AdT^{+}20, AZD19, AP22, AG23, ASME18, BS21, Ban23, BSE20, BFV^{+}21, BCN^{+}21, BS22a, Bie23, BHZ^{+}22, BPM^{+}18, CC18, CJ20, CGB^{+}19, CFRF^{+}23, CM18, Car20, CAV^{+}22, CFD^{+}20, CRLF^{+}23, DB22, DKM23, DOG21, ELL21, EMMH20, FKM^{+}21, dOFdJN^{+}22, FK19, GHHR20, GFM^{+}19, HS21, HSS22, HCS20, HWR^{+}19]. open [HHSP18, HKMdB22, HP20, ILG23, IPR22, JZS^{+}23, KK23, KOBT23, KAT18, KW23, LZV^{+}20, LSSA20, LRSC19, LS20b, LLL22, LKL^{+}21, LWS^{+}22, LKJ^{+}23, Lup17, MYG^{+}23, MKM^{+}22, MBML21, MLBR^{+}23, MCAR^{+}21, MBP^{+}22, MSR^{+}23, MRKP23, MBC^{+}20, MBH^{+}21, MTPK20, NFBR23, OGP22, OHF17, OCS21, OFP^{+}23, PPAP21, PDMV23, PK20, PPF^{+}22, PK23, Pea20, PTD^{+}20, PPG^{+}22, Poó22, Poó23, RMB18, RCM21, RC18, Ryu22a, SDDL20, SK23, SIR18, SKG22, SBTPV22, TCG^{+}21, TSF21, TLO^{+}22, VAG^{+}22, VA21, WYT23, WSB^{+}20, WBD^{+}21, WCR22, ZTS22, ZCB^{+}20, KK23, OFP^{+}23]. open-control [KW23]. Open-hardware [GCM21, KWP22, KOBT23, MYG^{+}23]. open-loop [IPR22]. Open-MAC [KK23]. Open-Source [KLS22, SIY^{+}23, BEES22, CED20, CNBH22, CKMW21, FZZ^{+}23, FDD17, HOW23, JCBBFVC23, KCW21a, KCW21b, KTT^{+}21, KST19, Kur21, LAW^{+}22, NAN19, NSZ^{+}23, PHL^{+}20, RSV^{+}23, SMW^{+}20, TS22, Ulr19, UMEA19, UB20, AZD19, AP22, AG23, Ban23, BSE20, BCN^{+}21, BS22a, Bie23, BHZ^{+}22, BPM^{+}18, CC18, CJ20, CM18, Car20, CRLF^{+}23, DKM23, DOG21, EMMH20, FKM^{+}21, dOFdJN^{+}22, GHHR20, HS21, HSS22, HKMdB22, ILG23, KK23, LZV^{+}20, LKL^{+}21, LWS^{+}22, LKJ^{+}23, Lup17, MKM^{+}22, MLBR^{+}23, MCAR^{+}21, MBP^{+}22, MSR^{+}23, MRKP23, MBH^{+}21, MTPK20, NFBR23, OCS21, OFP^{+}23, PPAP21, PK20, PK23, PPG^{+}22, Poó22, Poó23, RMB18, RC18, Ryu22a, SK23, SIR18, SKG22, TSF21, VAG^{+}22, VA21, WYT23, WSB^{+}20, WBD^{+}21, WCR22, ZTS22]. Open-Source-Powered [PSL^{+}20]. Open Irr [Bie23]. OpenASAP [SMS^{+}23]. OpenBCI [KSB22]. OpenEM [WSL^{+}19]. OpenFish [vdBSRW22]. OpenHumidistat [VdB22]. openPFGE [LSSA20]. OpenTCC [SDDL20]. OpenWorkstation [EMMH20]. operated [CRLF^{+}23, LNP^{+}23]. Operating [CAR^{+}23, SG21]. operation [BPM^{+}18]. operations [JC17]. operative [CLCT19]. optic [FAM^{+}18, BPP^{+}19]. optical [BVORMH^{+}22, CNBH22, CLSZ^{+}17, HYdM21, TBBI21]. optimization [ILL18]. Optimized [NMZMT20, TSF21]. opto [KK23, RJL^{+}23, VAG^{+}22]. opto-electronic [VAG^{+}22]. opto-microfluidic [RJL^{+}23]. opto-physiological [KK23]. optofluidic [NRK^{+}23]. optogenetic [WDW21]. optogenetics [ZCB^{+}20]. optomechanical [WS18b]. OptoPi [CFRF^{+}23]. orbital [NGS20]. organic [VOS22]. Organism [HOW23]. organoids [RMOB^{+}19]. orientated [OhMAN23]. orientation [BVMHP22b, GCCCBF22]. OSH [MDR20]. OSH-MVs [MDR20]. Osmar [CM18]. other [GPO^{+}20, SMB^{+}21, STP20]. Ötz [DKM23]. Ötz-T [DKM23]. outdoor [BVVA21]. Overcoming [OGP22]. Oxide [Bow23]. oximeter [SJB22]. oxygen [MKM^{+}22, MKVS21, NND^{+}22, RRS^{+}23]. oxygenator [AAUP23].

P300 [MP20]. package [LB17]. packaging [KOBT23]. PACKMAN [MRZMT21]. Pages [Ano17d, Ano17c, Ano18c]. pain [DHY18]. paintings [ILG23]. pandemic [BSE20]. pandemics [OGP22]. paper [MD22, VCS^{+}23]. paper-based [VCS^{+}23]. PARA [TCG^{+}21]. parallel [DCY^{+}22, KWP^{+}21]. Parametric [GPO^{+}20]. PARS [KG22]. part [LSD22, SKS23]. Partially [PTD^{+}20]. participation [CGB^{+}19]. particle [PNM^{+}23, RNRP23]. particulate [BSO^{+}23, HP20]. parts [HSS22, SK23]. passive [ZBOB22]. PassStat [CAV^{+}22]. patches [SZZ22]. patients [CABA^{+}21, MNMN^{+}20]. PAWS [NLYB23]. payload [TCG^{+}21]. pedestrian [MHBVBVMA23]. peen [SKG22]. pellets [MWG^{+}22]. pendulum [KFB^{+}23, MTPK20]. people [RHAB22]. performance [ARV^{+}22, BEES22, JC17, TSF21]. performing [MPB21]. perfusion [CWB^{+}21, JCBBFVC23, PK21]. peripheral [BCN^{+}21, SBT20]. peristaltic [CVT^{+}21, JTD20, MSWC21]. permanent [PAMH18, SSM^{+}22]. permittivity [PMA^{+}23]. personal [GCKF22, NHAH^{+}21, VA21]. personalized [Pri19]. Peta [SIR18]. Peta-pico-Voltron [SIR18]. pH [MPB21, NND^{+}22, OFP^{+}23]. phase [MGJ^{+}22]. phases [KAB^{+}20]. phenotyping [HWR^{+}19, LMB^{+}23, SGH^{+}18, TCC^{+}18]. Phone [Poó23, HSS22]. phones [Poó23]. photoacoustic [KFH22]. photocatalytic [MKM^{+}23, SHS^{+}23]. photochemistry [DYL^{+}18]. photodiode [MKVS21]. photogrammetry [Poó23, YXES21]. photogrammetry-based [YXES21]. photometric [RJL^{+}23]. photon [ZCB^{+}20]. photonically [SHS^{+}23]. photopolymers [DS21]. photoreactor [MKM^{+}23]. photovoltaic [BVMHP22c, CBT^{+}22, MRKP23]. physical [ZSLF23]. physical-ecological [ZSLF23]. Physiologic [HPC21, HPLC21, MKM^{+}22]. Physiologic-range [HPC21, HPLC21]. physiological [KK23, RGM20]. Pi [KPY^{+}18, SKC^{+}20, AG23, BJL^{+}20, CBT^{+}22, KKwM22, LDN^{+}22, SK23, WCR22]. pickup [CNBH22]. PICO [Kod18, SIR18]. piezoelectric [KPY^{+}18]. PiFlow [KPY^{+}18]. pin [LHN^{+}23]. pinch [SMSR21]. pipe [DMG23]. PiRamid [LDN^{+}22]. pitfall [McM17]. pixelated [SE23]. planar [PCCS23, SHS^{+}23]. plant [FMD^{+}22, LMB^{+}23, MZW^{+}21, TTA^{+}21]. PlasPI [PHL^{+}20, ZSLF23]. plastic [GC23, KHRC21, RNRP23, WMPP18]. plate [PK21]. plate-scale [PK21]. PlateFlo [PK21]. plates [HW22, KST19, Kur21]. Platform [LRP^{+}23, AFK^{+}19, BGJGC23, CFRF^{+}23, DKH^{+}23, GFTD20, GV20, JC17, JHA^{+}19, KCW21a, KSB22, LLL22, MCLSN19, MECH21, MP20, RC18, RHAB22, Ryu22b, SIY^{+}23, SBTPV22, UB20, WLML23, ZSLF23]. platforms [BSS^{+}20, OCS21]. play [DYL^{+}18, SVMV20, WLML23]. playback [CHBMCV^{+}23]. playhouse [WDW21]. plug [DYL^{+}18, Kur21, SVMV20, WLML23]. plug-and-play [SVMV20, WLML23]. plug-in [Kur21]. plug’n’play [CPVCRLS17]. pluripotent [RMOB^{+}19]. PLUSPULS [SGG23]. PMMA [RGSGD^{+}23]. pneumatic [BPM^{+}18]. point [WNKP22]. Poisson [SK23]. polar [SMB^{+}21]. polarimetry [MPASIM23]. polarity [DC23]. pollution [BVBTMVP23]. polymer [BEES22, DOV^{+}21, MLG^{+}20, MWG^{+}22, PWC23]. polymer-ceramic [PWC23]. polymeric [PWC23]. polyurethane [WCR22]. pore [BBJ21]. Portable [BVVA21, Car20, LZV^{+}20, POT^{+}22, CFV^{+}19, DBRE21, FLL23, GST^{+}22, HSH^{+}22, KG22, MRZMT21, MMR22, RBRK23, SJB22, Sur21, WNKP22, ZTS22]. PortaLyzer [PJS22]. position [GCCCBF22]. positioned [LKL^{+}21]. positioners [HHSP18]. positioning [CJ20]. potential [Bie23]. potentiostat [CAV^{+}22, DVD17, INS^{+}23, ICY21]. potentiostat/galvanostat [DVD17, ICY21]. Power [YM23, BSD^{+}22, CPLM23, DC23, DBRE21, HSH^{+}22, JC17, KKwM22, MGJ^{+}22, NGSU21, OP18, PPHP^{+}23, SIR18]. power-efficient [HSH^{+}22]. PowEred [NMZMT20, PSL^{+}20, HP20, MRKP23, RPHS23, RBB^{+}22, SIY^{+}23, VA21]. Powering [KKwM22]. PPE [STP20]. pre [ILG23]. pre-tensioning [ILG23]. precipitator [KWP22]. precise [CAV^{+}22, GML^{+}20, Rom21]. precision [CHG^{+}21, GHHR20, GFM^{+}19, LWS^{+}22, NKS22, RAMM20, VS22]. pregnant [DAS22]. preparation [YLW^{+}23]. preparative [DCY^{+}22, MLBR^{+}23]. Presence [HOW23]. presentation [DMC22]. press [BBPP23]. Pressure [YLW^{+}23, BBJ21, GHHR20, GML^{+}20, GCHP22, GCK^{+}23, HPC21, PHLM22, PWC23, SRC^{+}22, SK21, NMZMT20]. Pressure-controlled [YLW^{+}23]. pressurized [PTKG21]. prevention [RPHS23, TLO^{+}22]. previously [Ano21b]. print [ASIM21]. printable [BHF^{+}22, HHSP18, TBBI21, WS18b, WMPP18, WBD^{+}21]. printed [AHUS19, BVMHP22a, BM18, CVT^{+}21, CKB^{+}22, DKM23, DOG21, GCM21, HS21, HSS22, JKKK22, JTD20, KLS22, MSWC21, OKH22, PTKG21, RJL^{+}23, RWK22, SMS^{+}23, SK23, SSL^{+}20, STE^{+}20, SKK^{+}23, TPLT22, VBK^{+}22, VCS^{+}23]. printer [BS21, BOB^{+}19, BEES22, DÖ23, HW22, KHRC21, MKL^{+}21, PDR^{+}22]. printer-based [HW22, MKL^{+}21, PDR^{+}22]. printers [PWC23, PHF18, SFWW22]. printing [BEES22, DS21, DBP^{+}21, GPO^{+}20, KTK22, KTT^{+}21, MWG^{+}22, PWC23, STP20, SFWW22, TLP^{+}23, WMPP18]. pro [PLP23]. Probe [VKM^{+}23, HHSP18, PLP23, SMS^{+}23, SE23]. probes [SKK^{+}23]. process [LSD22, MECH21]. processes [WNKP22]. processor [BM19]. produce [DOV^{+}21, WSI^{+}22]. producing [PWC23]. production [GLG^{+}23, HS21, RJN^{+}23]. Professor [TCC^{+}18]. profiler [PEJ^{+}22, SMB^{+}21]. profiling [THMZ22]. program [RMB18]. programable [TLŠ^{+}22]. Programmable [BV18, EZ22, AHUS19, BS21, BRC^{+}20, BPM^{+}18, CLSZ^{+}17, DFS^{+}22, GLD18, KPY^{+}18, KMG^{+}21, NLYB23]. progress [Ano23a]. Prokaryote [WDW21]. PROPER [NMZMT20]. properties [CSS^{+}22, MBML21]. prosthesis [CPCQY23, REP19]. protection [WYT23]. protective [NHAH^{+}21, OKH22, VA21]. protein [BHF^{+}22, DCY^{+}22]. proteomics [FATM22]. protocols [SGG23]. prototype [GCKF22, GCCCBF22, GLPM19, GAB^{+}22, LLO22, MGJ^{+}22, MMR^{+}23]. prototypes [AAUP23]. Prototyping [ARLS22, Ryu22a, CPLM23]. Providing [RRS^{+}23]. Prusa [BOB^{+}19]. psychological [RGM20]. published [Ano21b]. pull [ICBL22]. pulse [BS22b, SJB22, SGG23]. pulsed [LSSA20]. pulsed-field [LSSA20]. pulses [TLŠ^{+}22]. pump [BS21, CVT^{+}21, DS21, DÖ23, GHHR20, GLD18, ICBL22, JTD20, KMG^{+}21, MSWC21, PPF^{+}22, PPHP^{+}23, PHF18, SMW^{+}20, GHHR20, JTD20, dCSMVFR23]. pumping [KPY^{+}18, MD20, MDD20]. pumps [GCM21, KPY^{+}18]. purifier [DCY^{+}22, JBG^{+}22]. purifying [HP20]. purity [CSH18, MKM^{+}23]. purpose [GLPM19]. push [ICBL22]. Pushing [PWC23]. PVA [WSI^{+}22]. Python [AG23]. PytuTester [MBP^{+}22].

quadruped [FZZ^{+}23]. quality

[ARLS22, BSO^{+}23, BGJGC23, FKM^{+}21, JHA^{+}19, MTPK20, Ryu22a, Ryu22b, Sur21]. Quantitative [NHAH^{+}21, MPB^{+}23]. Quartz [MMR^{+}23, HKMdB22]. quench [BPSO23].

rack [SS23]. radar [JKKK22]. radially [KDH^{+}19]. radiation [BFV^{+}21, BVMH21, BVMHP22a]. radiative [SWPG19]. radio [CC18, Kod18]. radio-telemetry [CC18]. randomly [JZS^{+}23]. range [CNBH22, HPC21, HPLC21, Ulr19]. rapid [CPLM23, GCK^{+}23, HWR^{+}19, THL^{+}20]. rapidly [SSM^{+}22]. rapidly-deployable [SSM^{+}22]. Raspberry [AG23, CBT^{+}22, KPY^{+}18, KKwM22, LDN^{+}22, SKC^{+}20, SK23, WCR22]. RaspberryPi [MBP^{+}22]. RaspyControl [AG23]. rate [SRC^{+}22, SJB22]. ratio [SK23]. rationing [JSS^{+}18]. rats [FK19]. ray [CSH18, KAB^{+}20]. reach [TCG^{+}21]. Reaction [GAB^{+}22]. reactions [CBTGAG^{+}18, HPU21, WCR22]. reactor [DYL^{+}18]. Real [AB23, AG23, BVVAGMRP22, BG21, CPB^{+}22, MECH21, RBB^{+}22, Ryu22a, Ryu22b]. Real-Time [AB23, AG23, BVVAGMRP22, BG21, CPB^{+}22, MECH21, Ryu22a, Ryu22b]. reality [RHAB22]. realized [TBBI21]. receiver [GCKF22]. receptacle [PTKG21]. Reconfigurable [OWHP23]. reconstruction [MSR^{+}23]. Recorder [GMB20, SKC^{+}20]. recording [HHKB20]. recordings [AdT^{+}20, KK23]. Recyclebot [WMPP18]. recycled [RNRP23]. recycling [GC23]. red [SKS23]. reduction [MKM^{+}23]. reference [OCS21]. ReFiBot [PDMV23]. regarding [Ano21b]. regime [BBJ21]. regions [SMB^{+}21]. regulating [MKM^{+}22]. regulatory [KMB^{+}22]. rehabilitation [KG22, RHAB22, TLP^{+}23]. related [GBDS22]. relative [LS20a, RAMM20]. release [CPB^{+}22, SZZ22]. Reliable [DMA^{+}23, LZV^{+}20, MKM^{+}23]. Remote [HPU21, AG23, ASME18, LLL22, LB17, LBK22, RBB^{+}22, TS22]. remotely [DMA^{+}23]. renewable [YA19]. repeated [PEJ^{+}22]. replicable [BHZ^{+}22, VSTB19]. RepRap [STP20]. RepRapable [PTD^{+}20, WMPP18]. reproducible [MKM^{+}23]. requirements [KMB^{+}22]. research [BVMHBV21, BMOTGHL23, BG21, CPLM23, CAR^{+}23, CHBMCV^{+}23, dOFdJN^{+}22, FES^{+}21, GV20, HKMdB22, KDH^{+}19, KOBT23, LHN^{+}23, MYG^{+}23, MBC^{+}20, RSV^{+}23, REP19, SRC^{+}22]. residential [ARLS22]. resistance [LS20b, SWPG19, SMSR21]. resistant [MPFC21]. resistivity [CFD^{+}20]. resolution [AKJT22, BCB^{+}22]. resonance [DOG21, KTK22]. resonator [KTK22]. Resource [AB23, BSO^{+}23, DAS22, VCS^{+}23]. resource-limited [VCS^{+}23]. RespiCo [KNAR22]. Respirator [NMZMT20, HP20, VA21]. Respiratory [GCHP22, GCK^{+}23, GPO^{+}20, HPC21, HPLC21, KNAR22]. Response [SAB^{+}18, HP20, KS21, NKS22, RPHS23, VA21]. results [NHAH^{+}21, vAvdGP^{+}22]. resuscitator [CABA^{+}21]. retains [PDR^{+}22]. retention [VSTB19]. retraction [TSF21]. review [Pea20]. REVOLVER [DCY^{+}22]. RF [WSL^{+}19]. RFID [CNPP21, JSS^{+}18]. RFIDuino [CNPP21]. rig [ACHB19, SHS^{+}23, SMW^{+}20]. ring [BCN^{+}21]. ring-down [BCN^{+}21]. rings [FLL23]. river [KwMMM23]. RNA [BHF^{+}22]. road [LMC18]. Robonerite [CLCT19]. robot [AZD19, AOAPH19, BSE20, BVMHBV21, FZZ^{+}23, KLS22, Lup17, NFBR23, PK20, RSV^{+}23, UB20, VBK^{+}22, WSL^{+}19]. robotic [CPCQY23, FSM^{+}21, PDMV23, TCG^{+}21, TTA^{+}21, dOdF23, vdBSRW22]. Robotically [LKL^{+}21]. robotics [BSE20, GV20]. ROBOTONT [RSV^{+}23]. Robust [BDPJGS23, SOB^{+}23, NLYB23]. rodents [JCFK^{+}23, KK23, KAT18]. roller [MBKP23]. ROMR} [NFBR23]. room [BV18, SG21]. rope [ZBOB22]. rope-less [ZBOB22]. ROS [NFBR23, RSV^{+}23]. ROS-based [NFBR23]. ROS-supported [RSV^{+}23]. Rotating [ZR20]. Rotating-liquid-based [ZR20]. rotator [DP17]. RotoMate [DOG21]. RT [ARV^{+}22]. rupture [CSS^{+}22].

SALAD [LNP^{+}23]. Sample [MCLSN19, DP17, HG21, PAMH18, YLW^{+}23]. sampler [MDD^{+}21]. Samplers [NLYB23, FES^{+}21]. samples [CSS^{+}22, THL^{+}20]. sampling [BCN^{+}21, BKM^{+}21, ELL21, GPO^{+}20, MSWC21, McM17, MD20, MBH^{+}21, NLYB23, Ryu22b, VOS22, VKM^{+}23]. Sandwich [LS20b]. sap [BSD^{+}22]. SASe [FES^{+}21]. savings [Pea20]. Scaffold [OhMA22]. scalable [BEES22, BKM^{+}21, LLL22]. scale [CFD^{+}20, DMG23, HS21, JCBBFVC23, LOSM18, LDN^{+}22, PK21, RNRP23, SGD^{+}22, SWPG19, WS18a, vAvdGP^{+}22]. scan [RHR19]. Scan4CFU [PPAP21]. scanner [YXES21]. Scanning [LMB^{+}23, VKM^{+}23, BHF^{+}22, GFM^{+}19, SE23, VDV23]. scattering [PZF23]. scenarios [HJvdG^{+}22]. scenes [DMC22]. scheduling [NGSU21]. scheme [FZZ^{+}23]. school [PAMC17]. science [FKM^{+}21, GLD18, PDMV23]. scientific [CGB^{+}19, MBKP23, Pea20]. screen [TTA^{+}21]. screening [ASIM21]. screw [BBPP23]. sea [KOBT23]. seascapes [CLCT19]. seawater [MPB21]. sediment [DKM23]. seedball [RJN^{+}23]. selective [MLG^{+}20]. self [HP20, LS20a]. self-contained [HP20, LS20a]. semi [VOS22]. semi-automated [VOS22]. sense [CPCQY23]. sensing [BSO^{+}23, BGJGC23, BGE^{+}20, CNBH22, JC17, KFH22, KDH^{+}19, KPKK19, POT^{+}22, TS22, dOdF23]. sensitive [HYdM21]. Sensor [SVMV20, AFK^{+}19, AHUS19, BSO^{+}23, BSS^{+}20, CPR^{+}21, DKM23, DMS^{+}22, FRJ^{+}22, GML^{+}20, HPC21, KCW21a, KS21, LKL^{+}21, MD22, NGSU21, OCS21, PMA^{+}23, POP^{+}20, SG21, SE23, SSM^{+}22, TLP^{+}23, Ulr19, ZTS22]. Sensors [MD22, BVMHP22a, FAM^{+}18, LLO22, MHBVBVMA23, Sur21]. Sensors-on-paper [MD22]. separation [KAB^{+}20, RGSGD^{+}23]. September [Ano23d]. sequential [DB22]. series [PHL^{+}20]. service [BG21, PBD^{+}22]. set [BS21]. settings [BSO^{+}23, DAS22, VCS^{+}23]. Setup [SKK^{+}23, BPM^{+}18, DMG23, NIKK21, SKS23, UMEA19, VSTB19]. shaker [DP17, NGS20]. shallow [Car20]. share [Pea21]. shear [SK23]. sheet [WNKP22]. shield [DKM23]. shields [BVMHP22a, CSH18]. shift [MGJ^{+}22]. shot [SKG22]. Shybo [Lup17]. Sidekick [KLS22]. Sieverts [CBTGAG^{+}18]. Sieverts-type [CBTGAG^{+}18]. sign [JKKK22]. signals [CPCQY23]. signs [DAS22]. SiLA2 [PSL^{+}20]. silent [ZBOB22]. simple [BGE^{+}20, CAV^{+}22, SOB^{+}23, SKS23]. simplified [KSB22, RHR19]. simplify [MDC^{+}19]. simulations [HJvdG^{+}22]. simulator [SL22]. simultaneous [KTT^{+}21, MSR^{+}23, NSZ^{+}23]. simultaneously [AdT^{+}20]. single [BPM^{+}18, NSZ^{+}23, WNKP22, YLW^{+}23]. single- [BPM^{+}18]. single-molecule [NSZ^{+}23, YLW^{+}23]. Singlet [MKVS21]. sintering [MLG^{+}20]. site [PMA^{+}23]. SitkaNet [CPR^{+}21]. situ [FAR22, DB22, FAM^{+}18, KAB^{+}20]. six [KWP^{+}21]. sized [SHS^{+}23]. sizing [KOBT23]. skin [LAW^{+}22]. skull [LHN^{+}23]. sky [JSG^{+}22]. SLA [GPO^{+}20, JTD20]. SLA-3D-printed [JTD20]. slide [PDR^{+}22]. small [CFRF^{+}23, CFD^{+}20, CSS^{+}22, HS21, JHA^{+}19, LOSM18, LDN^{+}22, MSWC21, SNAO20, THL^{+}20]. small-scale [HS21, LDN^{+}22]. Smart [PLP23, BVMHP22c, BSS^{+}20, HHKB20, JC17, MLBR^{+}23, PAMC17, CGB^{+}19]. smartphone [INS^{+}23, WBD^{+}21]. smARTS Museum V1 [ASME18]. smaTo [BSS^{+}20]. smAvo [BSS^{+}20]. smear [BDPJGS23]. SnapFib [GAB^{+}22]. snorkel [NHAH^{+}21]. soft [DBP^{+}21, PCCS23, SK23, SSL^{+}20, vdBSRW22]. software [PHLM22, PK21, THMZ22]. software-controllable [PK21]. soil [Bie23, PMA^{+}23, TTA^{+}21, vAvdGP^{+}22]. solid [MKM^{+}23, SMS^{+}23]. solidification [KAB^{+}20]. solids [CED20, SMS^{+}23]. solution [NGSU21, PMA^{+}23, PBD^{+}22]. Sonification [WRSU23]. sorting [McM17]. sound [ILL18]. Source [GME^{+}22, KLS22, LRP^{+}23, MDR20, PSL^{+}20, SIY^{+}23, dCSMVFR23, AIB21, ASIM21, AZD19, AP22, AG23, Ban23, BSE20, BFV^{+}21, BCN^{+}21, BS22a, BVMHBV21, Bie23, BEES22, BM18, BM19, BHZ^{+}22, BJL^{+}20, BPM^{+}18, CC18, CJ20, CFRF^{+}23, CM18, CED20, Car20, CAV^{+}22, CNBH22, CPB^{+}22, CFD^{+}20, CKMW21, CRLF^{+}23, DB22, DP17, DKM23, DOG21, ELL21, EMMH20, FZZ^{+}23, FKM^{+}21, FDD17, FATM22, dOFdJN^{+}22, FK19, GC23, GPO^{+}20, GHHR20, GFM^{+}19, HS21, HSS22, HCS20, HWR^{+}19, HHSP18, HKMdB22, HP20, HOW23, ILL18, ILG23, JZS^{+}23, JCBBFVC23, KK23, KCW21a, KCW21b, KPKK19, KTT^{+}21, KYA21, KAT18, KST19, Kur21, LZV^{+}20, LSSA20, LRSC19, LS20b, LLL22, LKL^{+}21, LWS^{+}22, LAW^{+}22, LKJ^{+}23, Lup17, MKM^{+}22, MBML21, MLBR^{+}23, MCAR^{+}21, MBP^{+}22, MSR^{+}23, MRKP23, MBKP23, MBC^{+}20, MMR^{+}23, MBH^{+}21, MTPK20, NAN19]. source [NSZ^{+}23, NFBR23, OP18, OGP22, OWHP23, OHF17, OCS21, OFP^{+}23, PRR^{+}23, PPAP21, PK20, PHLM22, PPF^{+}22, PK23, Pea20, PTD^{+}20, PPG^{+}22, PAMC17, Poó22, Poó23, POP^{+}20, PHL^{+}20, RSV^{+}23, RMB18, RCM21, RC18, Ryu22a, SDDL20, SK23, SIR18, SMW^{+}20, SKG22, STP20, TCG^{+}21, TSF21, TLO^{+}22, TS22, Ulr19, UMEA19, UB20, VAG^{+}22, VA21, WYT23, WSB^{+}20, WBD^{+}21, WMPP18, WCR22, YA19, ZTS22, ZCB^{+}20]. sources [GAMK23, SHS^{+}23]. Southern [BSS^{+}20]. space [GCKF22, MRZMT21, PAMH18]. spectral [BVVAD21, DMS^{+}22]. spectrometers [DOG21]. spectrometry [RC18, SMS^{+}23, VOS22]. spectrophotometer [FKM^{+}21, LZV^{+}20, WBD^{+}21]. spectroscopy [BCN^{+}21, CSH18, TKHB21]. spectrum [SAB^{+}18, ZCB^{+}20]. speed [BBPP23, GCCCBF22, LS20a, LAW^{+}22, WSB^{+}20, vdBSRW22]. spin [RWK22, SOB^{+}23, RMOB^{+}19]. Spinning [MKH^{+}20, GLG^{+}23, RMOB^{+}19]. split [GCK^{+}23]. SpO2 [SJB22]. SPOT [LMB^{+}23]. spotter [FATM22]. spp [CLCT19]. spun [DOV^{+}21]. SSVEP [MP20]. SSVEP-P300 [MP20]. stability [ARV^{+}22]. stabilization [PZF23, SKS23]. stage [CKMW21, MYG^{+}23, SSM^{+}22]. stalk [HWR^{+}19, SMSR21]. Stalker [HWR^{+}19]. stand [BS22a, KNAR22, PMA^{+}23]. stand-alone [KNAR22, PMA^{+}23]. standalone [TLO^{+}22]. Standardized [SSL^{+}20, KOBT23]. standards [KMB^{+}22]. state [MDR20]. Static [GLPM19]. Station [WRSU23, BVMHP22c, BVBTMVP23, CNPP21, GCKF22, CGB^{+}19, NAN19]. stations [HHSP18]. stem [RMOB^{+}19]. step [BRC^{+}20]. sterilizable [STP20]. stimulating [AOAPH19]. stimulation [MPFC21, RHAB22, ZCB^{+}20]. stimulations [REP19]. stimulator [CLSZ^{+}17, SGG23]. stimuli [MP20]. stopped [BPSO23]. stopped-flow [BPSO23]. storage [FAR22, NGSU21, YA19]. strain [LBK22, NKS22, WNKP22]. strand [KOBT23]. straws [TPLT22]. strength [HWR^{+}19]. stress [AFK^{+}19, SK21]. stress-controlled [SK21]. stretchable [WOM^{+}22]. stretcher [KWP^{+}21, STE^{+}20]. stretching [SBT20]. strips [HS21]. stroboscopic [TLŠ^{+}22]. strong [JBG^{+}22, SMB^{+}21]. studies [CC18, CSB21, CPB^{+}22, ILG23, PHL^{+}20, WRH^{+}20]. study [CC18, CBTGAG^{+}18, CPR^{+}21, KAB^{+}20, REP19]. studying [ARV^{+}22, RGM20, WSB^{+}20]. sub [DOV^{+}21, TLŠ^{+}22]. sub-micrometric [DOV^{+}21]. sub-microsecond [TLŠ^{+}22]. subjects [SNAO20]. submersible [MDD^{+}21, NLYB23]. subsonic [IPR22]. Subsurface [FES^{+}21]. Successive [GAB^{+}22]. such [NMZMT20]. suitable [LDN^{+}22]. suites [AHUS19]. sun [RBB^{+}22]. sunlight [BVVAGMRP22, BVORMH^{+}22]. super [AKJT22, DBP^{+}21]. super-resolution [AKJT22]. supply [DC23, OGP22, SIR18]. support [SNAO20, VKM^{+}23]. supported [RSV^{+}23]. surface [CFV^{+}19, CAR^{+}23, CPVCRLS17, McM17, PNM^{+}23, REP19, Ryu22a]. surface-active [McM17]. surface-mount [CPVCRLS17]. Surveillance [HOW23]. surveying [BS22a]. suspended [DKM23]. swab [GPO^{+}20]. swarm [GV20]. switch [RHR19]. switching [BPP^{+}19, KHRC21]. Synchronization [BMD^{+}20, SE23]. synchronous [SKC^{+}20]. synchrotron [KAB^{+}20]. synthesis [LOSM18, PÁBS23, VBK^{+}22]. synthesized [KMB^{+}22]. Syringe [BOB^{+}19, DS21, BS21, DÖ23, GLD18, ICBL22, PPHP^{+}23, PHF18, TSF21, LNP^{+}23]. Syringe-based [LNP^{+}23]. syringe-pump [DÖ23]. System [LOSM18, WRSU23, AdT^{+}20, AHUS19, AZD19, Ban23, BBJ21, Bie23, BRC^{+}20, BCB^{+}22, BVMH21, BVVAGMRP22, BVORMH^{+}22, CPLM23, CJ20, CGB^{+}19, CWB^{+}21, CKB^{+}22, CSS^{+}22, CHBMCV^{+}23, DS21, DB22, DBP^{+}21, ELL21, FLL23, GPO^{+}20, GBDS22, HW22, HCS20, JCBBFVC23, JKKK22, JSS^{+}18, KwMMM23, KFB^{+}23, KPY^{+}18, KG22, KOBT23, Kod18, KST19, LRSC19, LS20b, LKN21, LMC18, LKJ^{+}23, LBK22, LPB18, MGJ^{+}22, MN22, MPB21, MHBVBVMA23, MZLG20, MD20, MDD20, MMR22, NKM^{+}21, NMZMT20, NLYB23, NND^{+}22, OP18, PK20, PCCS23, PK21, PMA^{+}23, PPG^{+}22, Poó23, RKD^{+}22, RBRK23, REP19, RHWP22, SRB^{+}22, SMS^{+}23, SK23, SSL^{+}20, SKG22, SRC^{+}22, SNAO20, SGH^{+}18, TLO^{+}22, VOS22, VAG^{+}22, WS19, WYT23, WS18b]. system-on-chip [BRC^{+}20]. systems [AG23, HPC21, JTD20, PHL^{+}20].

T [DKM23, KTK22]. tabletop [GAB^{+}22]. tactile [FRJ^{+}22, dOdF23]. Tacto [dOdF23]. tags [CNPP21]. tailings [BBJ21]. Tailored [KTK22]. tangential [LBK^{+}21]. tangential-flow [LBK^{+}21]. Tank [OFP^{+}23, SBTPV22]. task [NGSU21]. tasks [dOdF23]. Taxonomy [MDR20]. TD [BBJ21]. TD-DAQ [BBJ21]. TDM [ZSLF23]. TDT [LS20b]. teaching [BSE20, CPLM23, MYG^{+}23]. tech [Pea21]. technical [KMB^{+}22]. technique [MBML21]. techniques [MCLSN19]. technologies [DMA^{+}23]. technology [BRC^{+}20, EMMH20, Pea20, TLP^{+}23]. Telechemistry [HPU21]. telemedicine [CABA^{+}21]. telemetry [CC18]. Temperature [MHBVBVMA23, AC20, BGJGC23, CWB^{+}21, CLCT19, CPB^{+}22, HJvdG^{+}22, KCW21b, MZW^{+}21, McM17, MD22, MMR^{+}23, OFP^{+}23, RAMM20, SDDL20, STP20, SK21]. temperature-control [SDDL20]. temperature-controlled [CPB^{+}22]. temperatures [CLCT19]. tensile [ILG23, MZLG20, SSL^{+}20, WOM^{+}22]. Tension [AAUP23]. tensioning [ILG23]. TENTACLE [HJvdG^{+}22]. term [CWB^{+}21, HG21, MKL^{+}21, Rom21]. terrestrial [RBB^{+}22]. test [ACHB19, FK19, PNM^{+}23, SHS^{+}23, SMW^{+}20, SMSR21, WOM^{+}22]. tester [AIB21, ILG23, MBP^{+}22, RCM21, AIB21]. testing [AFK^{+}19, BCB^{+}22, BJL^{+}20, CSS^{+}22, EZ22, LLO22, LKJ^{+}23, PCCS23, PMCL20, SGD^{+}22, SSL^{+}20, SZZ22, SK21]. tether [YM23]. textile [SWPG19]. textiles [UMEA19]. their [Pea21]. therapy [RHAB22]. Thermal [SMS^{+}22, LS20b, MD22, SWPG19]. thermometer [ASIM21, MCAR^{+}21]. thermoplastic [GLG^{+}23]. thermoplastics [KTT^{+}21]. thin [DVD17, GAB^{+}22, PÁBS23, RAMM20]. thin-film [DVD17]. Things [AB23, TS22]. Things-Based [AB23]. thoracic [GCHP22]. three [HPLC21, LHN^{+}23, TCG^{+}21]. three-DoF [TCG^{+}21]. three-pin [LHN^{+}23]. three/two [HPLC21]. three/two-way [HPLC21]. throughput [DHY18, LMB^{+}23, MZLG20, TLŠ^{+}22]. timber [BS22a]. Time [AB23, PMA^{+}23, AG23, BRC^{+}20, BVVAGMRP22, BG21, CPB^{+}22, LDN^{+}22, McM17, MECH21, NLYB23, PHL^{+}20, RBB^{+}22, Ryu22a, Ryu22b, WRH^{+}20]. Time-domain [PMA^{+}23]. time-integrated [NLYB23]. time-lapse [LDN^{+}22, WRH^{+}20]. time-sorting [McM17]. timed [PHLM22]. times [PPAP21]. timing [DMC22]. tinyML [BVBTMVP23]. TIRF [NSZ^{+}23]. tissue [LPK^{+}23, MKL^{+}21, PCCS23, PK23, RBRK23, SSL^{+}20]. TMA [PK23]. TMA-Mate [PK23]. Tomography [BHZ^{+}22]. tool [CKB^{+}22, PJS22, Sur21]. toolkit [PK23]. tools [GFM^{+}19]. top [AKJT22, KDH^{+}19]. topical [SZZ22]. TOSSIT [ZBOB22]. touchscreen [GLD18, RWK22]. tourniquet [LKJ^{+}23]. tracer [CBT^{+}22]. Track [BGJGC23, SGH^{+}18]. tracking [MTPK20, PPAP21]. traffic [LMC18]. transcranial [SGG23]. transfer [MGJ^{+}22, SWPG19]. transformed [BS21]. transimpedance [CKMW21]. transmission [PMA^{+}23]. transmitted [NMZMT20]. transpiration [BSD^{+}22]. transportable [HJvdG^{+}22]. Trap [HOW23, McM17, hTFC^{+}23]. treadmill [WSB^{+}20]. treatment [MCLSN19]. tree [BSD^{+}22, KS21]. TreeMMoSys [KS21, NKS22]. trees [NKS22]. triad [DMS^{+}22]. Triple [WM19, NSZ^{+}23]. triple-color [NSZ^{+}23]. tripods [HSS22]. tropical [CLCT19]. truncation [MBP19]. tunnel [IPR22, SKK^{+}23]. turbidity [DKM23, KDH^{+}19]. turbine [HSH^{+}22]. turbine-based [HSH^{+}22]. turtles [CC18]. twelve [FZZ^{+}23]. two [CKMW21, FK19, MMR22, TCG^{+}21]. two-bottle [FK19]. two-dimensional [MMR22]. two-kg [TCG^{+}21]. two-stage [CKMW21]. two-way [HPLC21]. type [CBTGAG^{+}18].

UAS [Ryu22b]. UAS-based [Ryu22b].

UASWQP [Ryu22b]. UAV

[RPHS23, SSM^{+}22]. UFBGA [AP22]. UGV [NKM^{+}21]. UK [KMB^{+}22]. ultra [JBG^{+}22, OhMAN23, PPF^{+}22, WSI^{+}22]. ultra-cost-effective [OhMAN23]. ultra-low-cost [WSI^{+}22]. ultraviolet [ARLS22, BV18]. unbiased [DHY18]. undergraduate [TBM20]. undersampled [SSM^{+}22]. understory [BS22a]. Underwater [HOW23, HCS20, MDD20, ZSLF23, ZBOB22]. uniaxial [KWP^{+}21]. Uniform [MKH^{+}20]. Unilamellar [BPW18]. Unit [BVMHBV23, CNBH22, SG21, Sur21]. units [OCS21]. unmanned [NKM^{+}21, Ryu22a]. unsaturated [BBJ21]. Untangling [KOBT23]. updated [RMOB^{+}19]. upgrade [SK21]. upper [KG22]. urban [CLCT19]. USB [DVD17]. USB-controlled [DVD17]. use [BGE^{+}20, DOG21, GML^{+}20, KMB^{+}22, LDN^{+}22, MN22, SGD19]. used [BPP^{+}19, SKK^{+}23]. uses [DBRE21]. using [ARLS22, AG23, BFV^{+}21, BCN^{+}21, BRC^{+}20, BVVAGMRP22, BVORMH^{+}22, CC18, CSH18, DMA^{+}23, FAM^{+}18, GCCCBF22, HSS22, JBG^{+}22, KTK22, MWG^{+}22, MHBVBVMA23, MCAR^{+}21, MP20, OKH22, PLP23, RNRP23, SVMV20, SK23, SHS^{+}23, SSL^{+}20, SOB^{+}23, TLP^{+}23]. USMA [MKH^{+}20]. USV [JHA^{+}19]. utilizing [BOB^{+}19, KPY^{+}18, OGP22]. UV [BFV^{+}21, BVMH21, GPO^{+}20, SHS^{+}23]. UV-A [SHS^{+}23]. UV-C [BFV^{+}21]. UVGI [ARLS22].

V [CBT^{+}22]. Validation

[LRP^{+}23, HKMdB22, MLG^{+}20, MPB^{+}23, PRR^{+}23, SG21]. valve [HPLC21, PTD^{+}20]. vaping [CSB21]. vapor [NIKK21]. var [GLPM19]. variable [WSB^{+}20]. variables [GBDS22]. vehicle [CFV^{+}19, CAR^{+}23, NKM^{+}21, Ryu22a]. vehicles [LB17]. velocity [RAMM20, SKK^{+}23]. vent [MNMN^{+}20]. ventilation [NAC^{+}22, OGP22, TLO^{+}22]. Ventilator [MDR20, AIB21, CABA^{+}21, HSH^{+}22, KMB^{+}22, MBP^{+}22, PTD^{+}20, RCM21]. ventilators [KMB^{+}22]. VentMon [RCM21]. Versatile [AC20, Ban23, CAV^{+}22, DMA^{+}23, GLG^{+}23, KCW21a, NSHA22, RBRK23]. Version [GCKF22]. vertical [SZZ22]. Vesicles [BPW18]. via [GFLMRV^{+}23]. vial [SS23]. vibration [KKV20, VS22]. video [HCS20, HHKB20, MTPK20, SKC^{+}20]. view [MHBVBVMA23]. Viola [IBL^{+}17]. violet [JBG^{+}22]. viral [NMZMT20, OKH22]. virtual [RHAB22]. viruses [GPO^{+}20]. vis [BVMH21]. visible [WBD^{+}21]. visual [DMC22, ZCB^{+}20]. visualization [Ryu22a, Ryu22b]. vital [DAS22, JKKK22]. VitRad [JKKK22]. vitrification [TPLT22]. vitro [LPK^{+}23, AC20, EMMH20, CSB21, CPB^{+}22, SZZ22]. vivo [KMG^{+}21, SBT20]. VOCs [VOS22]. volatile [VOS22]. Voltage [LRP^{+}23, DC23, GBDS22, KCW21b, LBK22, PRR^{+}23, SIR18]. voltammetry [INS^{+}23, RHR19]. VoltamoStat [INS^{+}23]. Voltron [SIR18]. volume [PHF18]. volumes [NSHA22].

wall [ASIM21]. Warden [DMS^{+}22]. Water [AB23, GME^{+}22, NLYB23, Bie23, BGJGC23, CAR^{+}23, FKM^{+}21, JHA^{+}19, KwMMM23, MD20, MDD^{+}21, PTKG21, Ryu22a, Ryu22b, SGD19]. waters [Car20, MSWC21, SMB^{+}21]. watersheds [SSM^{+}22]. wave [BCB^{+}22, JKKK22]. waveform [NAC^{+}22]. wavelengths [KDH^{+}19]. way [HPLC21]. We-VoltamoStat [INS^{+}23]. wear [CKB^{+}22]. Wearable [RGM20, AP22, INS^{+}23, NMZMT20, TKHB21]. Weather [NAN19, WRSU23, GCKF22, MRZMT21, Pri19]. WeatherChimes [WRSU23]. web [Kod18]. webcam [PPG^{+}22]. weed [DMS^{+}22, DMS^{+}22]. weight [DHY18, SMB^{+}21]. weight-bearing [DHY18]. well [HW22]. wells [KWP^{+}21]. western [BM19]. wet [MKH^{+}20]. wet-Spinning [MKH^{+}20]. Wheelchair [GCCCBF22]. wide [PAMH18]. wideband [Ulr19]. width [BRC^{+}20]. WiFi [CPVCRLS17]. wildlife [CC18, CHBMCV^{+}23]. wind [IPR22, KS21, NKS22, RBB^{+}22, SKK^{+}23]. wind-induced [KS21, NKS22]. winding [AAUP23]. wire [POP^{+}20]. wireless [BVMH21, BVMHP22c, BVBTMVP23, GCM21, LZV^{+}20, LPB18, MGJ^{+}22, RKD^{+}22, TLP^{+}23, ZTS22]. within [MKL^{+}21, PDDT23]. without [RGSGD^{+}23]. women [DAS22]. workflows [DCY^{+}22, EMMH20]. world [Pea21].

X [KAB^{+}20]. X-ray [KAB^{+}20]. XYZ} [LWS^{+}22, CJ20].

YOLO [AOAPH19]. Young [SK23].

yourself [TKHB21]. Ystruder [KPKK19].

Zappelin’ [ZCB^{+}20]. zero [CHG^{+}21, KPY^{+}18]. zero-friction [CHG^{+}21]. zones [PDDT23].

References

AliAbbasi:2023:TCH

[AAUP23] Easa AliAbbasi, Anil Akseki, Azmat Ullah, and Kerem Pekkan. Tension controlled hollowfiber winding machine for blood oxygenator prototypes. HardwareX, 14: ??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722300041X.

Agade:2023:GIT

[AB23] Piyush Agade and Eban Bean. GatorByte — an Internet of Things-based low-cost, compact, and real-time water resource monitoring buoy. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000342.

Asteriti:2020:VBT

[AC20] Sabrina Asteriti and Lorenzo Cangiano. Versatile bipolar temperature controller for custom in vitro applications. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722030064X.

Alhaddad:2019:LCT

[ACHB19] Ahmad Yaser Alhaddad, John-John Cabibihan, Ahmad Hayek, and Andrea Bonarini. A low-cost test rig for impact experiments on a dummy head. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300380.

Aguiar:2020:LCO

[AdT^{+}20] Leandro A. A. Aguiar, Nivaldo A. P. de Vasconcelos, Gabriela Chiuffa Tunes, Antonio J. Fontenele, Romildo de Albuquerque Nogueira, Marcelo Bussotti Reyes, and Pedro V. Carelli. Low-cost open hardware system for behavioural experiments simultaneously with electrophysiological recordings. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300419.

Agcayazi:2019:MAS

[AFK^{+}19] Talha Agcayazi, Marc Foster, Hannah Kausche, Max Gordon, and Alper Bozkurt. Multi-axis stress sensor characterization and testing platform. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300488.

Ariza:2023:RLF

[AG23] Jonathan Álvarez Ariza and Christian Nomesqui Galvis. RaspyControl Lab: a fully open-source and real-time remote laboratory for education in automatic control systems using Raspberry Pi and Python. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000032.

Alcala:2019:HMP

[AHUS19] José M. Lopez Alcala, Marja Haagsma, Chester J. Udell, and John S. Selker. HyperRail: Modular, 3D printed, 1–100 m, programmable, and low-cost linear motion control system for imaging and sensor suites. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300483.

Abuzairi:2021:CTL

[AIB21] Tomy Abuzairi, Ahli Irfan, and Basari. COVENT-Tester: a low-cost, open source ventilator tester. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000250.

Alsamsam:2022:MBT

[AKJT22] Mohammad Nour Alsamsam, Aurimas Kopūstas, Meda Jurevičiūtė, and Marijonas Tutkus. The miEye: Bench-top superresolution microscope with cost-effective equipment. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001134.

Anonymous:2017:EBa

[Ano17a]  Anonymous.  Editorial

Board. HardwareX, 1: i, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300196.

Anonymous:2017:EBb

[Ano17b]  Anonymous.  Editorial

Board. HardwareX, 2: i, October 2017. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300743.

Anonymous:2017:PO

[Ano17c] Anonymous. Pages 1–70 (October 2017). HardwareX, 2:1–70, October 2017. CODEN ???? ISSN 2468-0672.

Anonymous:2017:PA

[Ano17d] Anonymous. Pages 1–76 (April 2017). HardwareX, 1:1–76, April 2017. CO-DEN ???? ISSN 2468-0672.

Anonymous:2018:EB

[Ano18a]  Anonymous.  Editorial

Board. HardwareX, 3: i, April 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300373.

Anonymous:2018:O

[Ano18b]  Anonymous.  October

2018. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672.

Anonymous:2018:PA

[Ano18c] Anonymous. Pages 1–160 (April 2018). HardwareX, 3:1–160, April 2018. CO-DEN ???? ISSN 2468-0672.

Anonymous:2019:A

[Ano19a] Anonymous. April 2019.

HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672.

Anonymous:2019:O

[Ano19b]  Anonymous.  October

2019. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672.

Anonymous:2020:A

[Ano20a] Anonymous. April 2020.

HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672.

Anonymous:2020:O

[Ano20b]  Anonymous.  October

2020. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672.

Anonymous:2021:A

[Ano21a] Anonymous. April 2021. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. Anonymous:2021:ERP

[Ano21b] Anonymous. Erratum regarding previously published articles. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000651.

Anonymous:2021:O

[Ano21c]  Anonymous.  October

2021. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672.

Anonymous:2022:A

[Ano22a] Anonymous. April 2022. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672.

Anonymous:2022:O

[Ano22b]  Anonymous.  October

2022. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672.

Anonymous:2023:PD

[Ano23a] Anonymous. In progress (December 2023). HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672.

Anonymous:2023:J

[Ano23b] Anonymous. June 2023. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672.

Anonymous:2023:M

[Ano23c] Anonymous. March 2023.

HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672.

Anonymous:2023:S

[Ano23d] Anonymous. September 2023. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672.

Alves-Oliveira:2019:GBY

[AOAPH19] Patrícia Alves-Oliveira, Patrícia Arriaga, Ana Paiva, and Guy Hoffman. Guide to build YOLO, a creativity-stimulating robot for children. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300890.

Antunes:2022:OSU

[AP22] Rui Azevedo Antunes and Luís Brito Palma. An open-source UFBGA μboard for wearable devices. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000268.

Al-Rawi:2022:PLC

[ARLS22] Mohammad Al-Rawi, Annette Lazonby, and Callan Smith. Prototyping a lowcost residential air quality device using ultraviolet germicidal irradiation (UVGI) light. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722100081X.

Aguilar:2022:RBC

[ARV^{+}22] Maximiliano Aguilar, Sebastián Riffo, Antonio Veliz, Catalina González-Castaño, and Carlos Restrepo. RT Box card for studying the control communication impacts on microgrid performance and stability. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000670.

Abuzairi:2021:ITW

[ASIM21] Tomy Abuzairi, Nur Imaniati Sumantri, Ahli Irfan, and Ridho Maulana Mohamad. Infrared thermometer on the wall (iThermowall): an open source and 3-D print infrared thermometer for fever screening. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300778.

Asinelli:2018:SOH

[ASME18] Mainardo Gaudenzi Asinelli, Moisès Serra Serra, Judit Molera Marimòn, and Jordi Serra Espaulella. The smARTS Museum V1: an open hardware device for remote monitoring of Cultural Heritage indoor environments. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300834.

Allwright:2019:OSM

[AZD19] Michael Allwright, Weixu Zhu, and Marco Dorigo. An open-source multirobot construction system. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300786.

Bandilla:2023:VOS

[Ban23] Leonhard Bandilla. Versatile open-source fluorescence documentation system. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000573.

Basson:2021:TDL

[BBJ21] Jack Adriaan Basson, André Broekman, and Schalk Willem Jacobsz. TD-DAQ: a low-cost data acquisition system monitoring the unsaturated pore pressure regime in tailings dams. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722100050X.

Betts:2023:LSI

[BBPP23] J. Logan Betts, Frank M. Brinkley, Lauren B. Priddy, and Matthew W. Priddy. Low-speed instrumented drill press for bone screw insertion. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000810.

Bosma:2022:MDL

[BCB^{+}22] Bret Bosma, Ryan Coe, Giorgio Bacelli, Ted Brekken, and Budi Gunawan. Mini-DAQ: a lightweight, lowcost, high resolution, data acquisition system for wave energy converter testing. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000773.

Berry:2021:OSA

[BCN^{+}21] Timothy D. Berry, Chance Creelman, Nick Nickerson, Akio Enders, and Thea Whitman. An open-source, automated, gas sampling peripheral for laboratory incubation experiments using cavity ring-down spectroscopy. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000377.

Buitrago-Duque:2023:RCD

[BDPJGS23] Carlos Buitrago-Duque, Brayan Patiño-Jurado, and Jorge Garcia-Sucerquia. Robust and compact digital lensless holographic microscope for label-free blood smear imaging. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000159.

Birkelid:2022:HPP

[BEES22] Andreas Hagerup Birkelid, Sindre W. Eikevåg, Christer W. Elverum, and Martin Steinert. Highperformance polymer 3D printing — open-source liquid cooled scalable printer design. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000104.

Bentancor:2021:LOS

[BFV^{+}21] Marcel Bentancor, Sebastián Fernández, Federico Viera, Sarita Etcheverry, Carolina Poradosú, Pablo D’Angelo, Hernán Montemuiño, Santiago Mirazo, Álvaro Irigoyen, Analía Sanabria, and Horacio Failache. LUCIA: an open source device for disinfection of N95 masks using UV-C radiation. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000109.

Broekman:2021:LCM

[BG21] André Broekman and Petrus Johannes Gräbe. A low-cost, mobile realtime kinematic geolocation service for engineering and research applications. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000328.

Brown:2020:LCC

[BGE^{+}20] S. L. Brown, C. S. Goulsbra, M. G. Evans, T. Heath, and E. Shuttleworth. Low cost CO_{2} sensing: a simple microcontroller approach with calibration and field use. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300456.

Blanco-Gomez:2023:LCA

[BGJGC23] Pablo Blanco-Gómez, José Luis Jiménez-García, and Jose M.

Cecilia. Low-cost automated GPS, electrical conductivity and temperature sensing device (EC + T track) and Android platform for water quality monitoring campaigns. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001262.

Boillat:2022:CMD

[BH22] Marc-Aurèle Boillat and Peter C. Hauser. CO_{2}- measuring dongle. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000839.

Barthels:2022:LCP

[BHF^{+}22] Fabian Barthels, Stefan J.

Hammerschmidt, Tim R. Fischer, Collin Zimmer, Elisabeth Kallert, Mark Helm, Christian Kersten, and Tanja Schirmeister. A low-cost 3D-printable differential scanning fluorometer for protein and RNA melting experiments. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000013.

Brazey:2022:OSE

[BHZ^{+}22] B. Brazey, Y. Haddab, N. Zemiti, F. Mailly, and P. Nouet. An opensource and easily replicable hardware for Electrical Impedance Tomography. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000232.

Bierer:2023:DOS

[Bie23] Andrew M. Bierer. Development of an opensource soil water potential management system for horticultural applications, “Open Irr”. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000652.

Broekman:2020:FPO

[BJL^{+}20] André Broekman, Schalk Willem Jacobsz, Hendrik Louw, Elsabé Kearsley, Tiago Gaspar, and Talia Simone Da Silva Burke. Fly-by-Pi: Open source closedloop control for geotechnical centrifuge testing applications. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300602.

Booeshaghi:2021:LCS

[BKM^{+}21] A. Sina Booeshaghi, Yeokyoung (Anne) Kil, Kyung Hoi (Joseph) Min, Jase Gehring, and Lior Pachter. Low-cost, scalable, and automated fluid sampling for fluidics applications. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000304.

Bravo-Martinez:2018:OSP

[BM18] Jorge Bravo-Martinez. Open source 3D-printed 1000μ L micropump. HardwareX, 3:110–116, April 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300378.

Bravo-Martinez:2019:OSA

[BM19] Jorge Bravo-Martinez. Open source automated western blot processor. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300525.

Bilucaglia:2020:ELC

[BMD^{+}20] Marco Bilucaglia, Riccardo Masi, Giovanni Di Stanislao, Rita Laureanti, Alessandro Fici, Riccardo Circi, Margherita Zito, and Vincenzo Russo. ESB: a low-cost EEG Synchronization Box. HardwareX, 8:??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300341.

Bravo-Martinez:2023:ABI

[BMOTGHL23] Jorge Bravo-Martínez, Sonia Ortega-Tinoco, Julieta Garduño, and Salvador Hernández-López. Arduino based intra-cerebral microinjector device for neuroscience research. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000536.

Bessler:2019:YYX

[BOB^{+}19] Nils Bessler, Dennis Ogiermann, Maj-Britt Buchholz, Alexander Santel, Jan Heidenreich, Rawas Ahmmed, Holm Zaehres, and Beate Brand-Saberi. Nydus One Syringe Extruder (NOSE): a Prusa i3 3D printer conversion for bioprinting applications utilizing the FRESHmethod. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300877.

Bowman:2023:GGN

[Bow23] Samuel Bowman. GNOME: a Gaseous Nitrogen Oxide Measuring front-End for aqueous environmental materials. HardwareX, 15:??, September 2023.

CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000664.

Brower:2018:OSP

[BPM^{+}18] Kara Brower, Robert R.

Puccinelli, Craig J. Markin, Tyler C. Shimko, Scott A. Longwell, Bianca Cruz, Rafael Gomez-Sjoberg, and Polly M. Fordyce. An open-source, programmable pneumatic setup for operation and automated control of singleand multi-layer microfluidic devices. HardwareX, 3: 117–134, April 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300433.

Bhandare:2019:LCF

[BPP^{+}19] Aniket Bhandare, Abhishek Patnaik, David Pommerenke, Sachin Sharma, and Daniel Fischer. Low cost fast frequency switching driver for Acousto-Optic Modulators used in laser cooling. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721830066X.

Bujnowicz:2023:LCS

[BPSO23] Łukasz Bujnowicz, Rafał Pietras, Marcin Sarewicz, and Artur Osyczka. Lowcost stopped-flow and freeze-quench device for double mixing. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000160.

Bellon:2018:LCE

[BPW18] Javier Alejandro Bellon, Marcelo Javier Pino, and Natalia Wilke. Low-cost equipment for electroformation of giant unilamellar vesicles. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300440.

Bolanos:2020:ESS

[BRC^{+}20] Yamir H. Bolaños, Carlos F. Rengifo, Pablo E. Caicedo, Luis E. Rodriguez, and Wilson A. Sierra. Electronic system for step width estimation using programmable system-on-chip technology and time of flight cameras. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300353.

Baas:2021:EPK

[BS21] Sander Baas and Vittorio Saggiomo. Ender3 3D printer kit transformed into open, programmable syringe pump set. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000481.

Bessell:2022:LCO

[BS22a] Brandt Bessell and Aditya Singh. A low-cost opensource handheld LiDARbased automated understory timber stand surveying device. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000840.

Bujnowicz:2022:MPH

[BS22b] Łukasz Bujnowicz and Marcin Sarewicz. Multichannel pulse high-current driver of magnetic actuator. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000311.

Beslity:2022:LCL

[BSD^{+}22] Justin Beslity, Stephen B.

Shaw, John E. Drake, Jason Fridley, John C. Stella, Jordan Stark, and Kanishka Singh. A low cost, low power sap flux device for distributed and intensive monitoring of tree transpiration. HardwareX, 12:??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000967.

Benitez:2020:DAI

[BSE20] Victor H. Benitez, Rodrigo Symonds, and David E. Elguezabal. Design of an affordable IoT opensource robot arm for online teaching of robotics courses during the pandemic contingency. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300675.

Bainomugisha:2023:ASK

[BSO^{+}23] Engineer Bainomugisha, Joel Ssematimba, Deogratius Okedi, Anold Nsubuga, Marvin Banda, George William Settala, and Gideon Lubisia. AirQo sensor kit: a particulate matter air quality sensing kit custom designed for low-resource settings. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000895.

Broekman:2020:SSF

[BSS^{+}20] André Broekman, Wynand JvdM Steyn, Johannes LP Steyn, Malick Bill, and Lise Korsten. smAvo and smaTo: a fruity odyssey of smart sensor platforms in Southern Africa. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300651.

Bentancor:2018:PLC

[BV18] Marcel Bentancor and Sabina Vidal. Programmable and low-cost ultraviolet room disinfection device. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300452.

Botero-Valencia:2023:LCA

[BVBTMVP23] J. S. Botero-Valencia, C. Barrantes-Toro, D. Marquez-Viloria, and Joshua M. Pearce. Low-cost air, noise, and light pollution measuring station with wireless communication and tinyML. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000846.

Botero-Valencia:2021:MSU

[BVMH21] J. S. Botero-Valencia and M. Mejia-Herrera. Modular system for UV–vis– NIR radiation measurement with wireless communication. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000663.

Betancur-Vasquez:2021:OSO

[BVMHBV21] D. Betancur-Vásquez, M. Mejia-Herrera, and J. S. Botero-Valencia. Open source and open hardware mobile robot for developing applications in education and research. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000468.

Botero-Valencia:2023:DIM

[BVMHBV23] J. S. Botero-Valencia, M. Mejía-Herrera, and D. Betancur-Vásquez. Development of an inertial measurement unit (IMU) with datalogger and geopositioning for mapping the Earth’s magnetic field. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000925.

Botero-Valencia:2022:DID

[BVMHP22a] J. S. Botero-Valencia, M. Mejia-Herrera, and Joshua M. Pearce. Design and implementation of 3-D printed radiation shields for environmental sensors.

HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000128.

Botero-Valencia:2022:DLC

[BVMHP22b] J. S. Botero-Valencia, M. Mejia-Herrera, and Joshua M. Pearce. Design of a low-cost mobile multispectral albedometer with geopositioning and absolute orientation. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000694.

Botero-Valencia:2022:LCC

[BVMHP22c] J. S. Botero-Valencia, M. Mejia-Herrera, and Joshua M. Pearce. Low cost climate station for smart agriculture applications with photovoltaic energy and wireless communication. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000414.

Botero-Valencia:2022:LCSb

[BVORMH^{+}22] J. S. Botero-Valencia, E. Ospina-Rojas, M. Mejia-Herrera, D. Gonzalez-Montoya, M. Durango-Flórez, and C. A. Ramos-Paja. Low-cost system for sunlight incidence angle measurement using optical fiber. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000475.

Botero-Valencia:2021:PLC

[BVVA21] J. S. Botero-Valencia and J. Valencia-Aguirre. Portable low-cost IoT hyperspectral acquisition device for indoor/outdoor applications. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000456.

Botero-Valencia:2021:LCI

[BVVAD21] J. S. Botero-Valencia, J. Valencia-Aguirre, and D. Durmus. A low-cost IoT multi-spectral acquisition device. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722100002X.

Botero-Valencia:2022:LCSa

[BVVAGMRP22] J. S. Botero-Valencia, J. Valencia-Aguirre, D. Gonzalez-Montoya, and C. A. Ramos-Paja. A low-cost system for real-time measuring of the sunlight incident angle using IoT. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000177.

Chang:2021:MMV

[CABA^{+}21] Javier Chang, Augusto Acosta, Jorge Benavides-Aspiazu, Jaime Reategui, Christiam Rojas, Jordi Cook, Richard Nole, Luigi Giampietri, Sandra Pérez-Buitrago, Fanny L. Casado, and Benjamin Castaneda. Masi: a mechanical ventilator based on a manual resuscitator with telemedicine capabilities for patients with ARDS during the COVID-19 crisis. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722100016X.

Carvalho:2020:POS

[Car20] Matheus C. Carvalho.

Portable open-source autosampler for shallow waters. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300511.

Carvalho:2021:MMA

[Car21] Matheus C. Carvalho.

Miau, a microbalance autosampler. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000444.

Carlson:2023:CMA

[CAR^{+}23] Daniel F. Carlson, Serkan Akbulut, Jeppe Fogh Rasmussen, Christian Søndergård Hestbech, Marius Hjorth Andersen, and Claus Melvad. Compact and modular autonomous surface vehicle for water research: the Naval Operating Research Drone Assessing Climate Change (NOR-DACC). HardwareX, 15: ??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000603.

Caux:2022:PSF

[CAV^{+}22] Mélicia Caux, Anis Achit, Kethsovann Var, Gabriel Boitel-Aullen, Daniel Rose, Agnès Aubouy, Sylvain Argentieri, Raymond Campagnolo, and Emmanuel Maisonhaute. PassStat, a simple but fast, precise and versatile open source potentiostat. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000359.

Casado:2022:RPB

[CBT^{+}22] P. Casado, J. M. Blanes, C. Torres, C. Orts, D. Marroquí, and A. Garrigós.

Raspberry Pi based photovoltaic I-V curve tracer. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000074.

Carrillo-Bucio:2018:LCS

[CBTGAG^{+}18] Juan Luis Carrillo-Bucio, Juan Rogelio Tena-Garcia, Eduardo Paul Armenta-Garcia, Osiry Hernandez-Silva, José Gerardo Cabañas-Moreno, and Karina Suárez-Alcántara. Low-cost Sieverts-type apparatus for the study of hydriding/dehydriding reactions. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300439.

Cain:2018:OSH

[CC18] Patrick W. Cain and Matthew D. Cross. An open-source hardware GPS data logger for wildlife radio-telemetry studies: a case study using Eastern box turtles. HardwareX, 3:82–90, April 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300779.

Carvalho:2020:OSA

[CED20] Matheus C. Carvalho, William Eickhoff, and Michael Drexl. Opensource autosampler for elemental and isotopic analyses of solids. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300328.

Clement:2020:OOS

[CFD^{+}20] Rémi Clement, Yannick Fargier, Vivien Dubois, Julien Gance, Emile Gros, and Nicolas Forquet. OhmPi: an open source data logger for dedicated applications of electrical resistivity imaging at the small and laboratory scale. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300316.

Cano-Ferrer:2023:OOS

[CFRF^{+}23] Xavier Cano-Ferrer, Ruairí J. V. Roberts, Alice S. French, Joost de Folter, Hui Gong, Luke Nightingale, Amy Strange, Albane Imbert, and Lucia L. Prieto-Godino. OptoPi: an open source flexible platform for the analysis of small animal behaviour. HardwareX, 15: ??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000500.

Carlson:2019:APA

[CFV^{+}19] Daniel F. Carlson, Alexander Fürsterling, Lasse Vesterled, Mathias Skovby, Simon Sejer Pedersen, Claus Melvad, and Søren Rysgaard. An affordable and portable autonomous surface vehicle with obstacle avoidance for coastal ocean monitoring. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300161.

Camprodon:2019:SCK

[CGB^{+}19] Guillem Camprodon, Óscar González, Víctor Barberán, Máximo Pérez, Viktor Smári, Miguel Ángel de Heras, and Alejandro Bizzotto. Smart Citizen Kit and Station: an open environmental monitoring system for citizen participation and scientific experimentation. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300203.

Cupiche-Herrera:2023:LCA

[CHBMCV^{+}23] Vianney J. Cupiche-Herrera, Geovanni I. Balan-Medina, José D. Cú-Vizcarra, Alexis A. Mora-Roche, Jeisson Rodriguez-Valenzuela, and Brian McLaren. Low-cost automated call box system to conduct playback experiments for wildlife research and management. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000251.

Clonch:2021:HPZ

[CHG^{+}21] Cameron Clonch, Mark Huynh, Bryson Goto, Alexander Levin, John Selker, and Chet Udell. High precision zero-friction magnetic dendrometer. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722100078X.

Campbell:2020:DIL

[CJ20] Thomas Campbell and James F. X. Jones. Design and implementation of a low cost, modular, adaptable and open-source XYZ positioning system for neurophysiology. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300079.

Christiand:2022:MLC

[CKB^{+}22] Christiand, Gandjar Kiswanto, Ario Sunar Baskoro, Fachryal Hiltansyah, Muhammad Ramadhani Fitriawan, Ramandika Garindra Putra, Shabrina Kartika Putri, and Tae Jo Ko. MicroEye: a low-cost online tool wear monitoring system with modular 3D-printed components for micro-milling application. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000141.

Cretu:2021:OSL

[CKMW21] Vlad F. Cretu, Florian Kehl, Brandon C. Metz, and Peter A. Willis. Opensource lab hardware: Low noise adjustable two-stage gain transimpedance amplifier with DC offset for low-light detection. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000626.

Chan:2019:RLC

[CLCT19] Shelley H. M. Chan, Lynette H. L. Loke, Sam Crickenberger, and Peter A. Todd. Robonerite: a low-cost biomimetic temperature logger to monitor operative temperatures of a common gastropod (Nerita spp.) in tropical urban seascapes. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300252.

Chen:2017:POS

[CLSZ^{+}17] Xinping Chen, Walter D. Leon-Salas, Taylor Zigon, Donald F. Ready, and Vikki M. Weake. A programmable optical stimulator for the Drosophila eye. HardwareX, 2:13– 33, October 2017. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806721730024X.

Carvalho:2018:OOS

[CM18] Matheus C. Carvalho and Rachel H. Murray. Osmar, the open-source microsyringe autosampler. HardwareX, 3:10–38, April 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300287.

Chang:2022:OSF

[CNBH22] Tien-Jen Chang, Line Hagner Nielsen, Anja Boisen, and En-Te Hwu. Opensource force analyzer with broad sensing range based on an optical pickup unit. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000530.

Cassel:2021:RAB

[CNPP21] Mathieu Cassel, Oldrich Navratil, Franck Perret, and Hervé Piégay. The e-RFIDuino: an Arduinobased RFID environmental station to monitor mobile tags. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000390.

Christfort:2022:OSA

[CPB^{+}22] Juliane Fjelrad Christfort, Chrysillis Judy Magaard Polhaus, Pi Westi Bondegaard, Tien-Jen Chang, En Te Hwu, Line Hagner Nielsen, Kinga Zór, and Anja Boisen. Open source anaerobic and temperaturecontrolled in vitro model enabling real-time release studies with live bacteria. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000207.

Cutipa-Puma:2023:LCR

[CPCQY23] Diego Ronaldo Cutipa-Puma, Cristian Giovanni Coaguila-Quispe, and Pablo Raul Yanyachi. A low-cost robotic hand prosthesis with apparent haptic sense controlled by electroencephalographic signals. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000469.

Caldognetto:2023:FPE

[CPLM23] Tommaso Caldognetto, Andrea Petucco, Andrea Lauri, and Paolo Mattavelli. A flexible power electronic converter system with rapid control prototyping for research and teaching. HardwareX, 14:??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000184.

Chu:2021:SLC

[CPR^{+}21] Max Chu, Annette Patton, Josh Roering, Cora Siebert, John Selker, Cara Walter, and Chet Udell. SitkaNet: a low-cost, distributed sensor network for landslide monitoring and study. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000201.

Chamorro-Posada:2017:PWS

[CPVCRLS17] Pedro Chamorro-Posada, José Vázquez-Cabo, José L. Rodríguez, and José M. López-Santos. A plug’n’play WiFi surface-mount dualloop antenna. HardwareX, 1:46–53, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067216300104.

Cuervo:2023:LCO

[CRLF^{+}23] Rubén Cuervo, Miguel A.

Rodríguez-Lázaro, Ramon Farré, David Gozal, Gorka Solana, and Jorge Otero. Low-cost and open-source neonatal incubator operated by an Arduino microcontroller. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000640.

Campbell:2021:ELC

[CSB21] R. T. Campbell, V. Suresh, and K. S. Burrowes. ECAM: a low-cost vaping device for generating and collecting electronic cigarette condensate for in vitro studies. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000547.

Carvalho:2018:AHA

[CSH18] Matheus C. Carvalho, Christian J. Sanders, and Ceylena Holloway. Auto-HPGe, an autosampler for gamma-ray spectroscopy using high-purity germanium (HPGe) detectors and heavy shields. HardwareX, 4:??, October 2018.

CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806721830049X.

Corti:2022:BTS

[CSS^{+}22] Andrea Corti, Tariq Shameen, Shivang Sharma, Annalisa De Paolis, and Luis Cardoso. Biaxial testing system for characterization of mechanical and rupture properties of small samples. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000785.

Ching:2021:HCP

[CVT^{+}21] Terry Ching, Jyothsna Vasudevan, Hsih Yin Tan, Chwee Teck Lim, Javier Fernandez, Yi-Chin Toh, and Michinao Hashimoto. Highly-customizable 3Dprinted peristaltic pump kit. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000316.

Cantoni:2021:MCC

[CWB^{+}21] Federico Cantoni, Gabriel Werr, Laurent Barbe, Ana Maria Porras, and Maria Tenje. A microfluidic chip carrier including temperature control and perfusion system for long-term cell imaging. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000754.

Dese:2022:LCN

[DAS22] Kokeb Dese, Gelan Ayana, and Gizeaddis Lamesgin Simegn. Low cost, noninvasive, and continuous vital signs monitoring device for pregnant women in low resource settings (Lvital device). HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000219.

Deng:2022:OSC

[DB22] Zhaojie Deng and Brian J.

Beliveau. An open source 16-channel fluidics system for automating sequential fluorescent in situ hybridization (FISH)-based imaging. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000888.

Dine:2021:DNP

[DBP^{+}21] Andi Dine, Edward Bentley, Loic A. PoulmarcK, Daniele Dini, Antonio E. Forte, and Zhengchu Tan. A dual nozzle 3D printing system for super soft composite hydrogels. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000055.

Diep:2021:MPM

[DBRE21] Tai The Diep, Samuel Bizley, Partha Pratim Ray, and Alexander Daniel Edwards. MicroMI: a portable microbiological mobile incubator that uses inexpensive lithium power banks for field microbiology. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000729.

Davis:2023:LCA

[DC23] Eric J. Davis and Brian H. Clowers. Low-cost Arduino controlled dual-polarity high voltage power supply. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001274.

delCastillo-Santaella:2023:ABA

[dCSMVFR23] Teresa del Castillo-Santaella, Julia Maldonado-Valderrama, and Miguel Angel Fernandez-Rodriguez. Autotitrator based on an Arduino Open Source Pump. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000718.

Diep:2022:RLC

[DCY^{+}22] Patrick Diep, Jose L.

Cadavid, Alexander F. Yakunin, Alison P. McGuigan, and Radhakrishnan Mahadevan. REVOLVER: a low-cost automated protein purifier based on parallel preparative gravity column workflows. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000360.

Dunlap:2022:DFL

[DFS^{+}22] Christy Dunlap, Skylar Featherstone, Matthew Smith, Man Vu, Amanda Williams, Jason Bailey, and Han Hu. Design and fabrication of a lowcost and programmable dip coating machine. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001092.

Drackley:2018:IMW

[DHY18] Brendan Drackley, Matt Holtz, and Jay Yang. An inexpensive modified weight-bearing device assembled in-house for high throughput unbiased behavioral pain assessment in mice. HardwareX, 4: ??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300063.

Deschamps:2023:MAF

[DKH^{+}23] Joran Deschamps, Christian Kieser, Philipp Hoess, Takahiro Deguchi, and Jonas Ries. MicroFPGA: an affordable FPGA platform for microscope control. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000147.

Droujko:2023:OPO

[DKM23] Jessica Droujko, Felix Kunz, Jr., and Peter Molnar. Ötz-T: 3D-printed open-source turbidity sensor with Arduino shield for suspended sediment monitoring. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000020.

Dezord:2023:RVR

[DMA^{+}23] Clément Dezord, Gilles Micolau, Chahine Abbas, Arnaud Mesgouez, and Elisabeth Pozzo Di Borgo. Reliable, versatile and remotely controlled instrumentation of an hectometric loop antenna using appropriate technologies. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000706.

Cesarei:2022:LTC

[DMC22] Andrea De Cesarei, Michele Marzocchi, and Maurizio Codispoti. Luminance and timing control during visual presentation of natural scenes. HardwareX, 12:??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001213.

Demenois:2023:EBM

[DMG23] Morgan Demenois, Hong Yan Miao, and Frédérick P. Gosselin. Easy to build, modular and large scale pipe conveying fluid experimental setup. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000676.

Duncan:2022:WWL

[DMS^{+}22] Liam Duncan, Brendan Miller, Colton Shaw, Ryan Graebner, Marcelo L. Moretti, Cara Walter, John Selker, and Chet Udell. Weed Warden: a low-cost weed detection device implemented with spectral triad sensor for agricultural applications. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000487.

Demircan:2023:DAF

[DÖ23] Evren Demircan and Beraat Özçelik. Development of affordable 3D food printer with an exchangeable syringe-pump mechanism. HardwareX, 14:??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000378.

deOliveira:2023:BTC

[dOdF23] Thiago Eustaquio Alves de Oliveira and Vinicius Prado da Fonseca. BioIn-Tacto: a compliant multi-modal tactile sensing module for robotic tasks. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000858.

Filho:2022:IOS

[dOFdJN^{+}22] Maurício de Oliveira Filho, Matheus Cerqueira de Jesus, Anderson Zenken Nakazato, Marcel Yuzo Kondo, and Luis Rogerio de Oliveira Hein. Instrumented open-source filament extruder for research and education. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001079.

Dyga:2021:ROS

[DOG21] Marco Dyga, Christoph Oppel, and Lukas J. Gooßen. RotoMate: an open-source, 3D printed autosampler for use with benchtop nuclear magnetic resonance spectrometers. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000407.

Dominguez:2021:ALC

[DOV^{+}21] José E. Domínguez, E Olivos, Carlos Vázquez, J. M. Rivera, Rigoberto Hernández-Cortes, and Javier González-Benito. Automated lowcost device to produce sub-micrometric polymer fibers based on blow spun method. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722100047X.

Dhankani:2017:OSL

[DP17] Karankumar C. Dhankani and Joshua M. Pearce. Open source laboratory sample rotator mixer and shaker. HardwareX, 1: 1–12, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067216300049.

Darling:2021:SPE

[DS21] Cory Darling and Damon A. Smith. Syringe pump extruder and curing system for 3D printing of photopolymers. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000043.

Dobbelaere:2017:UCP

[DVD17] Thomas Dobbelaere, Philippe M. Vereecken, and Christophe Detavernier. A USBcontrolled potentiostat/ galvanostat for thin-film battery characterization. HardwareX, 2:34–49, October 2017. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300317.

Dellal:2018:LCP

[DYL^{+}18] David Dellal, Emma Yee, Shefali Lathwal, Hadley Sikes, and Jose Gomez-Marquez. Low-cost plug and play photochemistry reactor. HardwareX, 3: 1–9, April 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300524.

Efromson:2021:BOS

[ELL21] John P. Efromson, Shuai Li, and Michael D. Lynch. BioSamplr: an open source, low cost automated sampling system for bioreactors. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000067.

Eggert:2020:OMO

[EMMH20] Sebastian Eggert, Pawel Mieszczanek, Christoph Meinert, and Dietmar W. Hutmacher. OpenWorkstation: a modular opensource technology for automated in vitro workflows. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300614.

Einbergs:2022:PMT

[EZ22] Ernests Einbergs and Aleksejs Zolotarjovs. Programmable material testing device for mechanoluminescence measurements. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000943.

Fleming:2018:DES

[FAM^{+}18] Joe Fleming, Tazdin Amietszajew, Euan McTurk, David P. Towers, Dave Greenwood, and Rohit Bhagat. Development and evaluation of in-situ instrumentation for cylindrical Li-ion cells using fibre optic sensors. HardwareX, 3:100–109, April 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300809.

Fleming:2022:SEC

[FAR22] Joe Fleming, Tazdin Amietszajew, and Alexander Roberts. In-situ electronics and communications for intelligent energy storage. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000396.

Ficarro:2022:OSF

[FATM22] Scott B. Ficarro, William Max Alexander, Isidoro Tavares, and Jarrod A. Marto. Open source fraction Collector/MALDI spotter for proteomics. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000505.

Ferretti:2017:OSA

[FDD17] Jacopo Ferretti, Licia Di Pietro, and Carmelo De Maria. Open-source automated external defibrillator. HardwareX, 2:61– 70, October 2017. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300354.

Formel:2021:SAS

[FES^{+}21] Nathan Formel, Ian C.

Enochs, Chris Sinigalliano, Sean R. Anderson, and Luke R. Thompson. Subsurface automated samplers for eDNA (SASe) for biological monitoring and research. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000699.

Frie:2019:OSA

[FK19] Jude A. Frie and Jibran Y. Khokhar. An open source automated two-bottle choice test apparatus for rats. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300045.

Feng:2021:OSD

[FKM^{+}21] Jiansheng Feng, Banafsheh Khakipoor, Jacob May, Melissa Mulford, Joshua Davis, Kelly Siman, Gabrielle Russell, Adam W. Smith, and Hunter King. An open-source dual-beam spectrophotometer for citizenscience-based water quality monitoring. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000717.

Fernandez:2023:LCP

[FLL23] Joel Escobar Fernández, Cristian Martínez López, and Víctor Mosquera Leyton. A low-cost, portable 32-channel EIT system with four rings based on AFE4300 for body composition analysis. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223001013.

Falaschetti:2022:CBI

[FMD^{+}22] Laura Falaschetti, Lorenzo Manoni, Denis Di Leo, Danilo Pau, Valeria Tomaselli, and Claudio Turchetti. A CNN-based image detector for plant leaf diseases classification. HardwareX, 12:??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001080.

Fortune:2019:LCA

[FPC^{+}19] Benjamin C. Fortune, Christopher G. Pretty, Logan T. Chatfield, Lachlan R. McKenzie, and Michael P. Hayes. Low-cost active electromyography. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300501.

Fiedler:2022:LCF

[FRJ^{+}22] Niklas Fiedler, Philipp Ruppel, Yannick Jonetzko, Norman Hendrich, and Jianwei Zhang. Lowcost fabrication of flexible tactile sensor arrays. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001171.

Faraj:2021:FEH

[FSM^{+}21] Zanwar Faraj, Mert Selamet, Carlos Morales, Patricio Torres, Maimuna Hossain, Boyuan Chen, and Hod Lipson. Facially expressive humanoid robotic face. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300262.

Fang:2023:OSL

[FZZ^{+}23] Lizhou Fang, Junhui Zhang, Huaizhi Zong, Ximeng Wang, Kun Zhang, Jun Shen, and Zhenyu Lu. Open-source lower controller for twelve degrees of freedom hydraulic quadruped robot with distributed control scheme. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001389.

Gupta:2022:SEB

[GAB^{+}22] Ashoke Kumar Sen Gupta, Abu Adnan, Shantanu Bhattacharjee, Nipu Kumar Das, M. A. Matin, and Muhammad Quamruzzaman. SnapFib: an easy build Arduino based tabletop prototype for thin film deposition by Successive Ionic Layer Adsorption and Reaction method. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200092X.

Gibson:2023:MLS

[GAMK23] Graham M. Gibson, Robert Archibald, Mark Main, and Akhil Kallepalli. Modular light sources for microscopy and beyond (ModLight). HardwareX, 13:??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001304.

Guevara:2022:DLC

[GBDS22] Nelson E. Guevara, Yamir H.

Bolaños, Juan P. Diago, and Juan M. Segura. Development of a low-cost IoT system based on Lo-RaWAN for monitoring variables related to electrical energy consumption in low voltage networks. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200075X.

Gabriel:2023:OSI

[GC23] Alex Gabriel and Fabio Cruz. Open source IoTbased collection bin applied to local plastic recycling. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001341.

Gonzalez-Cely:2022:WPC

[GCCCBF22] Aura Ximena González-Cely, Mauro Callejas-Cuervo, and Teodiano Bastos-Filho. Wheelchair prototype controlled by position, speed and orientation using head movement. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000517.

Guy:2022:RBD

[GCHP22] Ella F. S. Guy, J. Geoffrey Chase, and Lui R. Holder-Pearson. Respiratory bi-directional pressure and flow data collection device with thoracic and abdominal circumferential monitoring. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000992.

Guy:2023:RPS

[GCK^{+}23] Ella F. S. Guy, Jaimey A. Clifton, Jennifer L. Knopp, Lui R. Holder-Pearson, and J. Geoffrey Chase. Respiratory pressure and split flow data collection device with rapid occlusion attachment. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000962.

Gibbons:2022:GVF

[GCKF22] John Gibbons, Kristina Collins, David Kazdan, and Nathaniel Frissell. Grape Version 1: First prototype of the low-cost personal space weather station receiver. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000347.

Gervasi:2021:OHW

[GCM21] Alain Gervasi, Pierre Cardol, and Patrick E. Meyer. Open-hardware wireless controller and 3D-printed pumps for efficient liquid manipulation. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000286.

Guerrero-Felix:2023:AAF

[GFLMRV^{+}23] Jesus Gerardo Guerrero-Felix, Javier Lopez-Miras, Miguel Angel Rodriguez-Valverde, Carmen Lucia Moraila-Martinez, and Miguel Angel Fernandez-Rodriguez. Automation of an atomic force microscope via Arduino. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000548.

Guver:2019:LCH

[GFM^{+}19] Alperen Guver, Nafetalai Fifita, Peker Milas, Michael Straker, Michael Guy, Kara Green, Taha Yildirim, Ilyas Unlu, Mehmet V. Yigit, and Birol Ozturk. A low-cost and high-precision scanning electrochemical microscope built with open source tools. HardwareX, 6:??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300495.

Greene:2020:CFL

[GFTD20] Austin Greene, Zac Forsman, Robert J. Toonen, and Megan J. Donahue. CoralCam: a flexible, low-cost ecological monitoring platform. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300537.

Gao:2020:UOS

[GHHR20] Run Ze Gao, Marie Hébert, Jan Huissoon, and Carolyn L. Ren. μPump: an open-source pressure pump for precision fluid handling in microfluidics. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300043.

Garcia:2018:LCT

[GLD18] Valentina E. Garcia, Jamin Liu, and Joseph L. De-Risi. Low-cost touchscreen driven programmable dual syringe pump for life science applications. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300269.

Gunther:2023:VHM

[GLG^{+}23] Jason Gunther, Jacques Lengaigne, Mélanie Girard, Valérie Toupin-Guay, James T. Teasdale, Martine Dubé, and Ilyass Tabiai. A versatile hot melt centrifugal spinning apparatus for thermoplastic microfibres production. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000615.

Guillardi:2019:GCP

[GLPM19] Hildo Guillardi, Eduardo Verri Liberado, José Antenor Pomilio, and Fernando Pinhabel Marafão. Generalcompensation-purpose static var compensator prototype. HardwareX, 5: ??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300518.

Gupta:2020:CRA

[GMB20] Saumya Gupta, Peter M.

Marchetto, and Mark A. Bee. Customizable recorder of animal kinesis (CRoAK): a multi-axis instrumented enclosure for measuring animal movements. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300250.

Gutierrez:2022:OSA

[GME^{+}22] Julián Gutiérrez, Juan Pablo Mochen, Gabriel Eggly, Marcelo Pistonesi, and Rodrigo Santos. Open source automated flow analysis instrument for detecting arsenic in water. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000293.

Goertzen:2020:LCP

[GML^{+}20] Luke Goertzen, Nicole Mehr, Manuel Lopez, Chet Udell, and John S. Selker. Low-cost and precise inline pressure sensor housing and DAQ for use in laboratory experiments. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300213.

Gallup:2020:PNS

[GPO^{+}20] Nicole Gallup, Adam M. Pringle, Shane Oberloier, Nagendra G. Tanikella, and Joshua M. Pearce. Parametric nasopharyngeal swab for sampling COVID-19 and other respiratory viruses: Open source design, SLA 3-D printing and UV curing system. HardwareX, 8:

??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300444.

Guglielmotti:2022:PAA

[GST^{+}22] Victoria Guglielmotti, Nicolás Andrés Saffioti, Ana Laura Tohmé, Martín Gambarotta, Gastón Corthey, and Diego Pallarola. A portable and affordable aligner for the assembly of microfluidic devices. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000931.

Gregory:2020:MLC

[GV20] Calvin Gregory and Andrew Vardy. microUSV: a low-cost platform for indoor marine swarm robotics research. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300134.

Hermann:2020:IIF

[HCS20] Andreas Hermann, Jérôme Chladek, and Daniel Stepputtis. iFO (infrared fish observation) — an open source low-cost infrared underwater video system. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300584.

Hohlbein:2023:OHM

[HF23] Johannes Hohlbein and Sanli Faez. Open hardware in microscopy. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000809.

Herrmann:2021:LCA

[HG21] Achim J. Herrmann and Michelle M. Gehringer. A low-cost automized anaerobic chamber for long-term growth experiments and sample handling. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000675.

Honeycutt:2020:LAS

[HHKB20] Wesley T. Honeycutt, Alyse V. Heaston, Jeffrey F. Kelly, and Eli S. Bridge. LunAero: Automated “smart” hardware for recording video of nocturnal migration. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300146.

Hietanen:2018:AOS

[HHSP18] Iiro Hietanen, Ismo T. S. Heikkinen, Hele Savin, and Joshua M. Pearce. Approaches to open source 3-D printable probe positioners and micromanipulators for probe stations. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300415.

Hermann:2022:TTH

[HJvdG^{+}22] Markus Hermann, Richard Jansen, Johan van de Glind, Edwin T. H. M. Peeters, and Paul J. Van den Brink. A transportable temperature and heatwave control device (TENTA-CLE) for laboratory and field simulations of different climate change scenarios in aquatic microand mesocosms. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000529.

Horst:2022:DVL

[HKMdB22] Rens J. Horst, Antonis Katzourakis, Bastian T. Mei, and Sissi de Beer. Design and validation of a low-cost open-source impedance based quartz crystal microbalance for electrochemical research. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001195.

Humbert:2023:OSC

[HOW23] Jefferson W. Humbert, Kirt L. Onthank, and Kresimir Williams. The open-source camera trap for organism presence and underwater surveillance (OCTOPUS). HardwareX, 13:??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000019.

Hubbard:2020:CSC

[HP20] Benjamin R. Hubbard and Joshua M. Pearce. Conversion of self-contained breathing apparatus mask to open source powered air-purifying particulate respirator for fire fighter COVID-19 response. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300389.

Holder-Pearson:2021:PRF

[HPC21] Lui Holder-Pearson and J. Geoffrey Chase. Physiologicrange flow and pressure sensor for respiratory systems. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000560.

Holder-Pearson:2021:PRT

[HPLC21] Lui Holder-Pearson, Theodore Lerios, and J. Geoffrey Chase. Physiologic-range three/two-way valve for respiratory circuits. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000638.

Hill:2019:ALC

[HPS^{+}19] Andrew P. Hill, Peter Prince, Jake L. Snaddon, C. Patrick Doncaster, and Alex Rogers. AudioMoth: a low-cost acoustic device for monitoring biodiversity and the environment. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300306.

Hsu:2021:TRM

[HPU21] Chun-Yao Hsu, Gurpur Rakesh D. Prabhu, and Pawel L. Urban. Telechemistry 2.0: Remote monitoring of fluorescent chemical reactions. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000742.

Han:2021:LCO

[HS21] Won Han and Joong Ho Shin. Low-cost, opensource 3D printed antibody dispenser for development and small-scale production of lateral flow assay strips. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000171.

Haque:2022:CEP

[HSH^{+}22] Syed Razwanul Haque, Shovon Sudan Saha, A. K. M. Maruf Hossain, Md Hasanur Rahman Sohag, Fozle Rabbi Shafi, Satya Ranjan Sarkar, and Tanzilur Rahman. Cost-effective and power-efficient portable turbine-based emergency ventilator. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000955.

Han:2022:LCO

[HSS22] Won Han, Jaeho Shin, and Joong Ho Shin. Low-cost, open-source contact angle analyzer using a mobile phone, commercial tripods and 3D printed parts. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000724.

Tsan:2023:LCA

[hTFC^{+}23] Vi hung Tsan, Daniel Fan, Sabina Caneva, Carlas S. Smith, and Gerard J. Verbiest. Low-cost acoustic force trap in a microfluidic channel. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000354.

Hasson:2022:CPB

[HW22] Or Hasson and Asher Wishkerman. CultureLED: a 3D printer-based LED illumination cultivation system for multi-well culture plates. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000682.

Heuschele:2019:SOS

[HWR^{+}19] D. Jo Heuschele, Jochum Wiersma, Leonard Reynolds, Amy Mangin, Yvonne Lawley, and Peter Marchetto. The Stalker: an open source force meter for rapid stalk strength phenotyping. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218301184.

Harvie:2021:SCO

[HYdM21] Andrew J. Harvie, Surendra K. Yadav, and John C. de Mello. A sensitive and compact optical detector based on digital lock-in amplification. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000572.

Irgens:2017:ECE

[IBL^{+}17] Peter Irgens, Curtis Bader, Theresa Lé, Devansh Saxena, and Cristinel Ababei. An efficient and cost effective FPGA based implementation of the Viola– Jones face detection algorithm. HardwareX, 1: 68–75, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067216300116.

Iannone:2022:LCP

[ICBL22] Marco Iannone, Diego Caccavo, Anna Angela Barba, and Gaetano Lamberti. A low-cost push–pull syringe pump for continuous flow applications. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000402.

Irving:2021:MCP

[ICY21] P. Irving, R. Cecil, and M. Z. Yates. MYS-TAT: a compact potentiostat/galvanostat for general electrochemistry measurements. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300729.

Idelson:2023:OSB

[ILG23] Antonio Iaccarino Idelson, Miguel Sánchez López, and Roger Groves. An open-source biaxial tensile tester with automated pretensioning for mechanical studies of canvas paintings. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000196.

Ibarra:2018:DCO

[ILL18] David Ibarra, Rodrigo Ledesma, and Edgar Lopez. Design and construction of an omnidirectional sound source with inverse filtering approach for optimization. HardwareX, 4: ??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300901.

Ibrahim:2023:WVW

[INS^{+}23] Nur Fatin Adini Ibrahim, Anas Mohd Noor, Norhayati Sabani, Zulkarnay Zakaria, Asnida Abdul Wahab, Asrulnizam Abd Manaf, and Shazlina Johari. We-VoltamoStat: a wearable potentiostat for voltammetry analysis with a smartphone interface. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000482.

Ismail:2022:ODL

[IPR22] Ismail, Erlanda Augupta Pane, and Reza Abdu Rahman. An open design for a low-cost openloop subsonic wind tunnel for aerodynamic measurement and characterization. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000979.

Jayakantha:2022:DCL

[JBG^{+}22] D. N. P. Ruwan Jayakantha, H. M. N. Bandara, Nadeesha M. Gunawardana, R. P. V. Jayantha Rajapakse, Dulari S. Thilakarathne, Elisabetta Comini, Nanda Gunawardhana, and S. M. M. L. Karunarathne. Design and construction of a low cost air purifier for killing harmful airborne microorganisms using a combination of a strong multidirectional electric-field and an ultra violet light. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000244.

Jiang:2017:HPL

[JC17] Jiming Jiang and Christian Claudel. A high performance, low power computational platform for complex sensing operations in smart cities. HardwareX, 1:22–37, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067216300177.

Jaramillo-Canas:2023:OSP

[JCBBFVC23] Winston Jaramillo-Cañas, Frank Britto-Bisso, Cesar Fernandez-Valiente, and Fanny L. Casado. Opensource perfusion system for medium-scale fabrication of demineralized bone matrix chip grafts. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001237.

Jamieson:2023:LCD

[JCFK^{+}23] Bradley B. Jamieson, Xavier Cano-Ferrer, George Konstantinou, Elisa de Launoit, Nicolas Renier, Albane Imbert, and Johannes Kohl. A low-cost device for cryoanesthesia of neonatal rodents. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722300024X.

Jo:2019:LCS

[JHA^{+}19] Wonse Jo, Yuta Hoashi, Lizbeth Leonor Paredes Aguilar, Mauricio Postigo-Malaga, José M. Garcia-Bravo, and Byung-Cheol Min. A low-cost and small USV platform for water quality monitoring. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300367.

Jeong:2022:VLC

[JKKK22] Hyunmin Jeong, Dohyun Kim, Gyoungdeuk Kim, and Sangkil Kim. Vit-Rad: a low-cost continuous wave Doppler radar system with 3D-printed horn antennas for human vital sign detection. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001067.

Jain:2022:LLC

[JSG^{+}22] Mayank Jain, Vishal Singh Sengar, Isabella Gollini, Michela Bertolotto, Gavin McArdle, and Soumyabrata Dev. LAMSkyCam: a low-cost and miniature ground-based sky camera. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000918.

Jha:2018:RBF

[JSS^{+}18] Rakesh Kumar Jha, Yamini Srivastav, Vedika Sumbli, Trisha, Vidula Gandhi, and Shubha Jain. RFID based food rationing system. HardwareX, 4: ??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300329.

Jonsson:2020:FPL

[JTD20] Alexander Jönsson, Arianna Toppi, and Martin Dufva. The FAST Pump, a low-cost, easy to fabricate, SLA-3D-printed peristaltic pump for multichannel systems in any lab. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300249.

Jadali:2023:FLC

[JZS^{+}23] Nima Jadali, Margaret J.

Zhang, Andrew K. Schulz, Josh Meyerchick, and David L. Hu. Forage-Feeder: a low-cost open source feeder for randomly distributing food. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000123.

Kim:2020:EEH

[KA20] Hubert Kim and Alan T. Asbeck. An elbow exoskeleton for haptic feedback made with a direct drive hobby motor. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300626.

Koe:2020:NEA

[KAB^{+}20] Billy Koe, Colin Abraham, Chris Bailey, Bob Greening, Martin Small, Thomas Connolley, and Jiawei Mi. A novel electromagnetic apparatus for in-situ synchrotron X-ray imaging study of the separation of phases in metal solidification. HardwareX, 7:??, April 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300122.

Kumbol:2018:AAO

[KAT18] Victor Wumbor-Apin Kumbol, Elikplim Kwesi Ampofo, and Mary Ayeko Twumasi. Actifield, an automated open source actimeter for rodents. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300695.

Kehl:2021:OSLa

[KCW21a] Florian Kehl, Vlad F.

Cretu, and Peter A. Willis. Open-source lab hardware: a versatile microfluidic control and sensor platform. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000584.

Kehl:2021:OSLb

[KCW21b] Florian Kehl, Vlad F.

Cretu, and Peter A. Willis. Open-source lab hardware: Driver and temperature controller for high compliance voltage, fiber-coupled butterfly lasers. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000705.

Kitchener:2019:LCB

[KDH^{+}19] Ben G. B. Kitchener, Simon D. Dixon, Kieren O. Howarth, Anthony J. Parsons, John Wainwright, Mark D. Bateman, James R. Cooper, Graham K. Hargrave, Edward J. Long, and Caspar J. M. Hewett. A low-cost bench-top research device for turbidity measurement by radially distributed illumination intensity sensing at multiple wavelengths. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300762.

Kaheman:2023:EMA

[KFB^{+}23] Kadierdan Kaheman, Urban Fasel, Jason J. Bramburger, Benjamin Strom, J. Nathan Kutz, and Steven L. Brunton. The experimental multi-arm pendulum on a cart: a benchmark system for chaos, learning, and control. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722300072X.

Keeratirawee:2022:LCE

[KFH22] Kanchalar Keeratirawee, Jasmine S. Furter, and Peter C. Hauser. Lowcost electronic circuitry for photoacoustic gas sensing. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000256.

Kocak:2022:PLC

[KG22] Mertcan Koçak and Erkin Gezgin. PARS, low-cost portable rehabilitation system for upper arm. HardwareX, 11:??, April 2022.

CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200044X.

Krige:2021:DIL

[KHRC21] Adolf Krige, Jakub Haluška, Ulrika Rova, and Paul Christakopoulos. Design and implementation of a low cost bio-printer modification, allowing for switching between plastic and gel extrusion. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000158.

Kapanaiah:2023:OML

[KK23] Sampath K. T. Kapanaiah and Dennis Kätzel. Open-MAC: a low-cost opensource motorized commutator for electro- and optophysiological recordings in freely moving rodents. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000366.

Koene:2020:ICD

[KKV20] Ivar Koene, Ville Klar, and Raine Viitala. IoT connected device for vibration analysis and measurement. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300171.

Kiarie:2022:DPM

[KKwM22] Gabriel Kiarie, Jason Kabi, and Ciira wa Maina. DSAIL power management board: Powering the Raspberry Pi autonomously off the grid. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000827.

Kazanskaya:2020:LCC

[KLF^{+}20] Rogneda B. Kazanskaya, Alexander V. Lopachev, Tatiana N. Fedorova, Raul R. Gainetdinov, and Anna B. Volnova. A low-cost and customizable alternative for commercial implantable cannula for intracerebral administration in mice. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300298.

Keesey:2022:SLC

[KLS22] Rodolfo Keesey, Robert LeSuer, and Joshua Schrier. Sidekick: a low-cost opensource 3D-printed liquid dispensing robot. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000645.

Kazlovich:2022:OVE

[KMB^{+}22] Kate Kazlovich, Soumya Ranjan Mishra, Kamran Behdinan, Aviv Gladman, Jesse May, and Azad Mashari. Open ventilator evaluation framework: a synthesized database of regulatory requirements and technical standards for emergency use ventilators from Australia, Canada, UK, and US. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000050.

Kujawa:2021:LCP

[KMG^{+}21] Maciej Kujawa, Szymon Motała, Michał Gonet, Rafał Pietrzyk, Tomasz Czechowski, and Mikołaj Baranowski. Low-cost, programmable infusion pump with bolus mode for in-vivo imaging. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000237.

Kulkarni:2022:RNF

[KNAR22] Kanchan Kulkarni, John H.

Nichols, Antonis A. Armoundas, and Jesse D. Roberts, Jr. RespiCo: a novel, flexible, and standalone electronic respiratory coaching device. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000803.

Koch:2023:UGK

[KOBT23] Jack C. Koch, Allyssa M. Oune, Sarah Bodenstein, and Terrence R. Tiersch. Untangling the Gordian knot of Aplysia sea hare egg masses: an integrated open-hardware system for standardized egg strand sizing and packaging for cryopreservation research and application. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000834.

Kodera:2018:AAS

[Kod18] Toshiro Kodera. Adaptive antenna system by ESP32-PICO-D4 and its application to web radio system. HardwareX, 3:91–99, April 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300871.

Klar:2019:YOS

[KPKK19] Ville Klar, Joshua M. Pearce, Pyry Kärki, and Petri Kuosmanen. Ystruder: Open source multifunction extruder with sensing and monitoring capabilities. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300471.

Kassis:2018:PBL

[KPY^{+}18] Timothy Kassis, Paola M.

Perez, Chloe J. W. Yang, Luis R. Soenksen, David L. Trumper, and Linda G. Griffith. PiFlow: a biocompatible low-cost programmable dynamic flow pumping system utilizing a Raspberry Pi Zero and commercial piezoelectric pumps. HardwareX, 4: ??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300130.

Kolbe:2021:TLC

[KS21] Sven Kolbe and Dirk Schindler. TreeMMoSys: a low cost sensor network to measure wind-induced tree response. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000092.

Knierim:2022:SDC

[KSB22] Michael T. Knierim, Max Schemmer, and Niklas Bauer. A simplified design of a cEEGrid ear-electrode adapter for the Open-BCI biosensing platform. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722200102X.

Kurata:2019:OSC

[KST19] Kosaku Kurata, Keita Sumida, and Hiroshi Takamatsu. Open-source cell extension system assembled from laser-cut plates. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300956.

Kemper:2022:TBR

[KTK22] Philip Kemper, Jorg Thöming, and Ekkehard Küstermann. Tailored birdcage resonator for magnetic resonance imaging at 7 T using 3D printing. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000712.

Koch:2021:OSH

[KTT^{+}21] Fritz Koch, Ole Thaden, Kevin Tröndle, Roland Zengerle, Stefan Zimmermann, and Peter Koltay. Open-source hybrid 3Dbioprinter for simultaneous printing of thermoplastics and hydrogels. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000596.

Kurata:2021:OSC

[Kur21] Kosaku Kurata. Opensource colorimeter assembled from laser-cut plates and plug-in circuits. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300705.

Kvalsund:2023:DAB

[KW23] Anniken Semb Kvalsund and Dietmar Winkler. Development of an Arduinobased, open-control interface for hardware in the loop applications. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000950.

Kabi:2023:LCL

[KwMMM23] Jason N. Kabi, Ciira wa Maina, Edwell T. Mharakurwa, and Stephen W. Mathenge. Low cost, LoRa based river water level data acquisition system. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000214.

Kah:2021:LCU

[KWP^{+}21] Delf Kah, Alexander Winterl, Magdalena Přechová, Ulrike Schöler, Werner Schneider, Oliver Friedrich, Martin Gregor, and Ben Fabry. A low-cost uniaxial cell stretcher for six parallel wells. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300717.

Kasparoglu:2022:OHD

[KWP22] Sabin Kasparoglu, Timothy P. Wright, and Markus D. Petters. Openhardware design and characterization of an electrostatic aerosol precipitator. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000116.

Koirala:2021:OSA

[KYA21] Dipak Koirala, Nicholas Yensen, and Peter B. Allen. Open source alliron battery 2.0. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300808.

Liao:2022:OSC

[LAW^{+}22] Hsien-Shun Liao, Imtisal Akhtar, Christian Werner, Roman Slipets, Jorge Pereda, Jen-Hung Wang, Ellen Raun, Laura Olga Nørgaard, Frederikke Elisabet Dons, and Edwin En Te Hwu. Open-source controller for low-cost and high-speed atomic force microscopy imaging of skin corneocyte nanotextures. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000864.

Liardon:2017:AIP

[LB17] Jean-Luc Liardon and D. A. Barry. Adaptable imaging package for remote vehicles. HardwareX, 2:1–12, October 2017. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721730007X.

Luken:2021:ATF

[LBK^{+}21] Arne Lüken, Maike Bruckhaus, Udo Kosfeld, Meike Emondts, and Matthias Wessling. Automated tangential-flow diafiltration device. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000298.

Louw:2022:MDM

[LBK22] Hendrik Louw, André Broekman, and Elsabé Kearsley. MADV-DAQ: Multi-channel Arduinobased differential voltage data acquisition system for remote strain measurement applications. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001055.

Long:2022:PCR

[LDN^{+}22] Matthew Michael Long, Tai The Diep, Sarah Helen Needs, Marta Joan Ross, and Alexander Daniel Edwards. PiRamid: a compact Raspberry Pi imaging box to automate smallscale time-lapse digital analysis, suitable for laboratory and field use. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001225.

Lohss:2023:ATP

[LHN^{+}23] Maxwell Lohss, Jonathan Ho, Nathan Naylor, Stacy Cashman, Roxana Fu, S. Tonya Stefko, and Leah C. Byrne. Adaptable three-pin skull clamp for large animal research. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000792.

Liardon:2018:LIM

[LHZ^{+}18] Jean-Luc Liardon, Lukas Hostettler, Ludovic Zulliger, Karl Kangur, Nawaaz S. Gujja Shaik, and D. A. Barry. Lake imaging and monitoring aerial drone. HardwareX, 3:146– 159, April 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300718.

Liu:2023:DMO

[LKJ^{+}23] Dawei Liu, Apoorv Kulkarni, Victoria F. Jaqua, Christina A. Cole, and Joshua M. Pearce. Distributed manufacturing of an open-source tourniquet testing system. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000494.

Levintal:2021:ERP

[LKL^{+}21] Elad Levintal, Kenneth Lee Kang, Lars Larson, Eli Winkelman, Lloyd Nackley, Noam Weisbrod, John S. Selker, and Chester J. Udell. eGreenhouse: Robotically positioned, low-cost, opensource CO_{2} analyzer and sensor device for greenhouse applications. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000225.

Leal:2021:MDA

[LKN21] Daniel Pinheiro Leal, Jörg Krämer, and Wilfried Nörtershäuser. Motordriven autonomous system for controlling beamline iris diaphragm apertures. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300754.

Lavayssiere:2022:LBS

[LLL22] Camille Lavayssière, Benoît Larroque, and Franck Luthon. Laborem Box: a scalable and open source platform to design remote lab experiments in electronics. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000463.

Lay:2022:HAB

[LLO22] Kenjiro S. Lay, Lingqi Li, and Masataka Okutsu. High altitude balloon testing of Arduino and environmental sensors for CubeSat prototype. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000748.

Lantin:2023:SSP

[LMB^{+}23] Stephen Lantin, Kelli Mc-Court, Nicholas Butcher, Varun Puri, Martha Esposito, Sasha Sanchez, Francisco Ramirez-Loza, Eric McLamore, Melanie Correll, and Aditya Singh. SPOT: Scanning plant IoT facility for highthroughput plant phenotyping. HardwareX, 15: ??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000755.

Lei:2018:IBT

[LMC18] Tian Lei, Abduallah A.

Mohamed, and Christian Claudel. An IMU-based traffic and road condition monitoring system. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300683.

Le:2023:SSB

[LNP^{+}23] Anh Phuc Hoang Le, Quang Lam Nguyen, Bao Hoai Pham, Thien Hoang Minh Cao, Toi Van Vo, Khon Huynh, and Huong Thi Thanh Ha. SALAD: Syringebased Arduino-operated Low-cost Antibody Dispenser. HardwareX, 15: ??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000627.

LeSuer:2018:OOM

[LOSM18] R. J. LeSuer, K. L. Osgood, K. E. Stelnicki, and J. L. Mendez. OMIS: the open millifluidic inquiry system for small scale chemical synthesis and analysis. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806721830052X.

Lund:2018:MBW

[LPB18] John Lund, Anthony Paris, and Jennifer Brock. Mouthguard-based wireless high-bandwidth helmetmounted inertial measurement system. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300087.

Loverde:2023:DBV

[LPK^{+}23] Joseph R. Loverde, Maria Piroli, George J. Klarmann, Joel Gaston, J. Kenneth Wickiser, Jason Barnhill, Kristin H. Gilchrist, and Vincent B. Ho. Development of a bioreactor for in-vitro compression cycling of tissue engineered meniscal implants. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000408.

Lezcano:2023:DVM

[LRP^{+}23] Hernan Lezcano, Jorge Rodas, Julio Pacher, Magno Ayala, and Carlos Romero. Design and validation of a modular control platform for a voltage source inverter. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001353.

Lau:2019:AOS

[LRSC19] Soon Kiat Lau, Felipe Azevedo Ribeiro, Jeyamkondan Subbiah, and Chris R. Calkins. Agenator: an open source computercontrolled dry aging system for beef. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300355.

Lau:2020:HSC

[LS20a] Soon Kiat Lau and Jeyamkondan Subbiah. HumidOSH: a self-contained environmental chamber with controls for relative humidity and fan speed. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722030050X.

Lau:2020:TSO

[LS20b] Soon Kiat Lau and Jeyamkondan Subbiah. TDT Sandwich: an open source dry heat system for characterizing the thermal resistance of microorganisms. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300237.

Laumann:2022:DMP

[LSD22] Daniel Laumann, Dieter Spiehl, and Edgar Dörsam. Device for measuring part adhesion in FFF process. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000037.

Lagos-Susaeta:2020:OOS

[LSSA20] Diego Lagos-Susaeta, Oriana Salazar, and Juan A. Asenjo. openPFGE: an open source and low cost pulsed-field gel electrophoresis equipment. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300377.

Lupetti:2017:SOS

[Lup17] Maria Luce Lupetti. Shybo. An open-source low-anthropomorphic robot for children. HardwareX, 2:50–60, October 2017. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721730038X.

Liao:2022:LCO

[LWS^{+}22] Hsien-Shun Liao, Christian Werner, Roman Slipets, Peter Emil Larsen, Ing-Shouh Hwang, Tien-Jen Chang, Hans Ulrich Danzebrink, Kuang-Yuh Huang, and En-Te Hwu. Lowcost, open-source XYZ nanopositioner for highprecision analytical applications. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000621.

Laganovska:2020:PLC

[LZV^{+}20] Katrina Laganovska, Aleksejs Zolotarjovs, Mercedes Vázquez, Kirsty Mc Donnell, Janis Liepins, Hadar Ben-Yoav, Varis Karitans, and Krisjanis Smits. Portable low-cost opensource wireless spectrophotometer for fast and reliable measurements. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722030016X.

Mouy:2020:FLC

[MBC^{+}20] Xavier Mouy, Morgan Black, Kieran Cox, Jessica Qualley, Callum Mireault, Stan Dosso, and Francis Juanes. FishCam: a low-cost open source autonomous camera for aquatic research. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300195.

Murray:2021:CDA

[MBH^{+}21] Greggory Murray, Samuel Bednarski, Michael Hall, Samuel W. Foster, Si-Jun Jin, Joshua J. Davis, Wei Xue, Eric Constans, and James P. Grinias. Comparison of design approaches for low-cost sampling mechanisms in opensource chemical instrumentation. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000493.

Mottaghi:2023:OSS

[MBKP23] Maryam Mottaghi, Yuntian Bai, Apoorv Kulkarni, and Joshua M. Pearce. Open source scientific bottle roller. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000524.

Massara:2021:IOS

[MBML21] Nazareno Massara, Enrico Boccaleri, Marco Milanesio, and Mattia Lopresti. IETeasy: an open source and low-cost instrument for impulse excitation technique, applied to materials classification by acoustical and mechanical properties assessment. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000602.

Mariola:2019:NAK

[MBP19] M. Mariola, C. Bemont, and F. Petruccione. A novel analogue keyboard for embedded applications, based on integer division truncation. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300099.

Morales:2022:PRO

[MBP^{+}22] Félix Morales, Luis Bernal, Gustavo Pereira, Sandra Pérez-Buitrago, Michael Kammer, and D. H. Stalder. PytuTester: RaspberryPi open-source ventilator tester. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000797.

Mnati:2021:OSN

[MCAR^{+}21] Mohannad Jabbar Mnati, Raad Farhood Chisab, Azhar M. Al-Rawi, Adnan Hussein Ali, and Alex Van den Bossche. An open-source non-contact thermometer using lowcost electronic components. HardwareX, 9: ??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000122.

Medina:2019:STP

[MCLSN19] Deyber Arley Vargas Medina, Luís Felipe Rodriguez Cabal, Fernando Mauro Lanças, and Álvaro José Santos-Neto. Sample treatment platform for automated integration of microextraction techniques and liquid chromatography analysis. HardwareX, 5:??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300907.

McMunn:2017:TSP

[McM17] Marshall S. McMunn. A time-sorting pitfall trap and temperature datalogger for the sampling of surface-active arthropods.

HardwareX, 1:38–45, April 2017. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067216300220.

Mucciarone:2020:AMF

[MD20] David A. Mucciarone and Robert B. Dunbar. Automated multiport flowthrough water pumping and sampling system. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300560.

Mulla:2022:SPF

[MD22] Rafiq Mulla and Charles W.

Dunnill. Sensors-on-paper: Fabrication of graphite thermal sensor arrays on cellulose paper for large area temperature mapping. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000821.

Montoya:2019:MSC

[MDC^{+}19] Rafael Álvarez Montoya, Sara Delgado, José Castilla, José Navarrete, Nuria Díaz Contreras, Juan Ramón Marijuan, Víctor Barrena, Isabel Guillamón, and Hermann Suderow. Methods to simplify cooling of liquid helium cryostats. HardwareX, 5:??, April 2019.

CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300889.

Mucciarone:2020:AUP

[MDD20] David A. Mucciarone, Hans B. DeJong, and Robert B. Dunbar. Autonomous underwater pumping system. HardwareX, 8:??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300638.

Mucciarone:2021:ASM

[MDD^{+}21] David A. Mucciarone, Hans B. DeJong, Robert B. Dunbar, Yui Takeshita, Rebecca Albright, and Keaton Mertz. Autonomous submersible multiport water sampler. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000262.

Mora:2020:COS

[MDR20] Simone Mora, Fábio Duarte, and Carlo Ratti. Can open source hardware mechanical ventilator (OSH-MVs) initiatives help cope with the COVID-19 health crisis? Taxonomy and state of the art. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300596.

Mieszczanek:2021:AME

[MECH21] Pawel Mieszczanek, Sebastian Eggert, Peter Corke, and Dietmar W. Hutmacher. Automated melt electrowritting platform with real-time process monitoring. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000766.

Martinez:2022:LCM

[MGJ^{+}22] Andrés Martínez, Christian González, Adrián Jaramillo, Dorindo Cárdenas, and Alejandro Von Chong. Low-cost, microcontrollerbased phase shift measurement system for a wireless power transfer prototype. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000566.

Mejia-Herrera:2023:LCS

[MHBVBVMA23] M. Mejia-Herrera, J. S.

Botero-Valencia, D. Betancur-Vásquez, and E. A. Moncada-Acevedo. Lowcost system for analysis pedestrian flow from an aerial view using near-infrared, microwave, and temperature sensors. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722300010X.

Mitropoulos:2020:UWS

[MKH^{+}20] Alexander N. Mitropoulos, Kylor T. Kiesewetter, Eric Horne, Jeff Butler, Joseph R. Loverde, and J. Kenneth Wickiser. Uniform wet-Spinning Mechanically Automated (USMA) fiber device. HardwareX, 8:??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722030033X.

Merces:2021:IPB

[MKL^{+}21] George O. T. Merces, Conor Kennedy, Blanca Lenoci, Emmanuel G. Reynaud, Niamh Burke, and Mark Pickering. The incubot: a 3D printer-based microscope for long-term live cell imaging within a tissue culture incubator. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000183.

Marchus:2022:LCO

[MKM^{+}22] Colin R. N. Marchus, Jacob A. Knudson, Alexandra E. Morrison, Isabell K. Strawn, Andrew J. Hartman, Dev Shrestha, Nicholas M. Pancheri, Ian Glasgow, and Nathan R. Schiele. Low-cost, opensource cell culture chamber for regulating physiologic oxygen levels. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000833.

Moustakas:2023:HPG

[MKM^{+}23] Nikolaos G. Moustakas, Marcus Klahn, Bastian T. Mei, Anna Pougin, Martin Dilla, Tim Peppel, Simon Ristig, and Jennifer Strunk. A high-purity gas–solid photoreactor for reliable and reproducible photocatalytic CO_{2} reduction measurements. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722300055X.

Moskalensky:2021:SOL

[MKVS21] Alexander E. Moskalensky, Tatyana Yu. Karogodina, Alexey Yu. Vorobev, and Sergei G. Sokolovski. Singlet oxygen luminescence detector based on low-cost InGaAs avalanche photodiode. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000535.

Medina:2023:OSS

[MLBR^{+}23] Deyber Arley Vargas Medina, Asdrubal Lozada-Blanco, Julie Paulin García Rodríguez, Fernando Mauro Lanças, and Álvaro José Santos-Neto. An opensource smart fraction collector for isocratic preparative liquid chromatography. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722300069X.

Mapley:2020:DVL

[MLG^{+}20] Martin Mapley, Yidi Lu, Shaun D. Gregory, Jo P. Pauls, Geoff Tansley, and Andrew Busch. Development and validation of a low-cost polymer selective laser sintering machine. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300286.

Matheny:2019:LLE

[MMP^{+}19] Ashley M. Matheny, Peter Marchetto, Je’aime Powell, Austin Rechner, Joon yee Chuah, Erica Mc-Cormick, and Suzanne A. Pierce. LEAF: Logger for ecological and atmospheric factors. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721930015X.

Munoz:2022:LCP

[MMR22] Juan D. Muñoz, Víctor H.

Mosquera, and Carlos F. Rengifo. A low-cost, portable, two-dimensional bioimpedance distribution estimation system based on the AD5933 impedance converter. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000190.

Munoz:2023:QCM

[MMR^{+}23] G. G. Muñoz, M. J.

Millicovsky, J. M. Reta, J. I. Cerrudo, A. Peñalva, M. Machtey, R. M. Torres, and M. A. Zalazar. Quartz crystal microbalance with dissipation monitoring for biomedical applications: Open source and low cost prototype with active temperature control. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000238.

McCauely:2022:DML

[MN22] Dalyn M. McCauely and Lloyd L. Nackley. Development of mini-lysimeter system for use in irrigation automation of containergrown crops. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000438.

Mathanlal:2020:AVA

[MNMN^{+}20] Thasshwin Mathanlal, Miracle Israel Nazarious, Roberto Mantas-Nakhai, Maria-Paz Zorzano, and Javier Martin-Torres. ATMOvent: an adapted breathing atmosphere for COVID-19 patients. HardwareX, 8:??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300547.

Mouli:2020:DHS

[MP20] Surej Mouli and Ramaswamy Palaniappan. DIY hybrid SSVEP-P300 LED stimuli for BCI platform using EMOTIV EEG headset. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300225.

Martinez-Prat:2023:MFM

[MPASIM23] Berta Martínez-Prat, Oriol Arteaga, Francesc Sagués, and Jordi Ignés-Mullol. Multimodal fluorescence microscope with fast adaptive polarimetry. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000871.

McLean:2021:LCA

[MPB21] Keegan M. McLean, Alexis L.

Pasulka, and Emily E. Bockmon. A low-cost, accessible, and high-performing Arduino-based seawater pH control system for biological applications. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000778.

Moustafa:2023:FVL

[MPB^{+}23] Tarek E. Moustafa, Edward R. Polanco, Rachel L. Belote, Robert L. Judson-Torres, and Thomas A. Zangle. Fabrication and validation of an LED array microscope for multimodal, quantitative imaging. HardwareX, 13: ??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000068.

McKenzie:2021:LCS

[MPFC21] Lachlan R. McKenzie, Christopher G. Pretty, Benjamin C. Fortune, and Logan T. Chatfield. Lowcost stimulation resistant electromyography. HardwareX, 9:??, April 2021.

CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000079.

Mottaghi:2023:AGP

[MRKP23] Maryam Mottaghi, Motakabbir Rahman, Apoorv Kulkarni, and Joshua M. Pearce. AC/off-grid photovoltaic powered opensource ball mill. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000305.

Mathanlal:2021:PPI

[MRZMT21] Thasshwin Mathanlal, Abhilash Vakkada Ramachandran, Maria-Paz Zorzano, and Javier Martin-Torres. PACKMAN — a portable instrument to investigate space weather. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722030078X.

Moreno:2023:OSM

[MSR^{+}23] Xavier Casas Moreno, Mariline Mendes Silva, Johannes Roos, Francesca Pennacchietti, Nils Norlin, and Ilaria Testa. An opensource microscopy framework for simultaneous control of image acquisition, reconstruction, and analysis. HardwareX, 13:

??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722300007X.

McCarthy:2021:BFP

[MSWC21] David T. McCarthy, Baiqian Shi, Miao Wang, and Stephen Catsamas. BoSL FAL pump: a small, lowcost, easily constructed, 3D-printed peristaltic pump for sampling of waters. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000432.

Myers:2020:LCD

[MTPK20] Audun D. Myers, Joshua R.

Tempelman, David Petrushenko, and Firas A. Khasawneh. Low-cost double pendulum for high-quality data collection with open-source video tracking and analysis. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722030047X.

Martin:2022:LCP

[MWG^{+}22] Vincent Martin, Jean-François Witz, Frédéric Gillon, Denis Najjar, Philippe Quaegebeur, Abdelkader Benabou, Michel Hecquet, Emmanuel Berté, François Lesaffre, Matthieu Meersdam, and Delphine Auzene. Low cost 3D printing of metals using filled polymer pellets. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000372.

Mahant:2023:OHC

[MYG^{+}23] Sunandan Mahant, Shweta Yadav, Cameron Gilbert, Eva R. Kjærgaard, Mads M. Jensen, Tommy Kessler, Merete Bilde, and Markus D. Petters. An open-hardware community ice nucleation cold stage for research and teaching. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000986.

Mirshahidi:2020:HTT

[MZLG20] Kiana Mirshahidi, Kamran Alasvand Zarasvand, Wenting Luo, and Kevin Golovin. A high throughput tensile ice adhesion measurement system. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300559.

McDowell:2021:CTC

[MZW^{+}21] Kyle McDowell, Yang Zhong, Kira Webster, Hector Jaime Gonzalez, A. Zachary Trimble, and Camilo Mora. Comprehensive temperature controller with internet connectivity for plant growth experiments. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000687.

Ng:2022:CDA

[NAC^{+}22] Qing Arn Ng, Christopher Yew Shuen Ang, Yeong Shiong Chiew, Xin Wang, Chee Pin Tan, Mohd Basri Mat Nor, Nor Salwa Damanhuri, and J. Geoffrey Chase. CAREDAQ: Data acquisition device for mechanical ventilation waveform monitoring. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001031.

Netto:2019:OSA

[NAN19] Guilherme Tomaschewski Netto and Jorge Arigony-Neto. Open-source Automatic Weather Station and Electronic Ablation Station for measuring the impacts of climate change on glaciers. HardwareX, 5:??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300749.

Nwankwo:2023:RRB

[NFBR23] Linus Nwankwo, Clemens Fritze, Konrad Bartsch, and Elmar Rueckert. ROMR: a ROS-based open-source mobile robot. HardwareX, 14:??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000330.

Nedbal:2020:BIO

[NGS20] Jakub Nedbal, Lu Gao, and Klaus Suhling. Bottomilluminated orbital shaker for microalgae cultivation. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300523.

Nguyen:2021:HBL

[NGSU21] Bao Nguyen, Bryson Goto, John S. Selker, and Chet Udell. Hypnos board: a low-cost all-in-one solution for environment sensor power management, data storage, and task scheduling. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000420.

Nicholson:2021:MFF

[NHAH^{+}21] Kyle Nicholson, Ashley Henke-Adams, Daniel M. Henke, Alexxai V. Kravitz, and Hiram A. Gay. Modified full-face snorkel mask as COVID-19 personal protective equipment: Quantitative results. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000146.

Niketa:2021:ACV

[NIKK21] A. K. Niketa, Md Aasif Ikbal, Susmitha Kothapalli, and Shishir Kumar. An automated chemical vapor deposition setup for 2D materials. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300742.

Naglak:2021:CDS

[NKM^{+}21] John E. Naglak, Caleb Kase, Max McGinty, Casey D. Majhor, Carl S. Greene, Jeremy P. Bos, and Wayne W. Weaver. Cable deployment system for unmanned ground vehicle (UGV) mobile microgrids. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000341.

Nickl:2022:ETH

[NKS22] Julius Nickl, Sven Kolbe, and Dirk Schindler. Enhancing TreeMMoSys with a high-precision strain gauge to measure the windinduced response of trees down to the ground. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001249.

Neumann:2023:PAW

[NLYB23] Kyle C. Neumann, Daniel La, Hyemin Yoo, and Deron E. Burkepile. Programmable Autonomous Water Samplers (PAWS): an inexpensive, adaptable and robust submersible system for time-integrated water sampling in freshwater and marine ecosystems. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001377.

Nazarious:2020:POP

[NMZMT20] Miracle Israel Nazarious, Thasshwin Mathanlal, Maria-Paz Zorzano, and Javier Martin-Torres. Pressure Optimized Pow-Ered Respirator (PROPER): a miniaturized wearable cleanroom and biosafety system for aerially transmitted viral infections such as COVID-19. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300535.

Nguyen:2022:LCS

[NND^{+}22] Dung Kim Nguyen, Huy Quang Nguyen, Huyen Thuy T. Dang, Viet Quoc Nguyen, and Linh Nguyen. A lowcost system for monitoring pH, dissolved oxygen and algal density in continuous culture of microalgae. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000980.

Nagalingam:2023:LCF

[NRK^{+}23] Nagaraj Nagalingam, Aswin Raghunathan, Vikram Korede, Edwin F. J. Overmars, Shih-Te Hung, Remco Hartkamp, Johan T. Padding, Carlas S. Smith, and Huseyin Burak Eral. Low-cost fluorescence microscope with microfluidic device fabrication for optofluidic applications. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000226.

Nasir:2022:DIV

[NSHA22] Muhammad Nasir, Muhammad Shoaib, Muhammad Umar Hassan, and Muhammad Sabieh Anwar. Design and implementation of a versatile magnetic field mapper for 3D volumes. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001018.

Niederauer:2023:KOS

[NSZ^{+}23] Christian Niederauer, Marco Seynen, Jan Zomerdijk, Marko Kamp, and Kristina A. Ganzinger. The K2: Opensource simultaneous triplecolor TIRF microscope for live-cell and singlemolecule imaging. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000111.

Oliveira:2021:DFM

[OCS21] Franklin Oliveira, Daniel G. Costa, and Ivanovitch Silva. On the development of flexible mobile multisensor units based on opensource hardware platforms and a reference framework. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000730.

Onthank:2023:OAT

[OFP^{+}23] Kirt L. Onthank, James Foster, E. Preston Carman Jr, John E. Foster, Monica Culler-Juarez, Eliam Calvo, Wesley Duerksen, Trevor Natiuk, and Lucas Saca. The Open acidification Tank Controller: an open-source device for the control of pH and temperature in ocean acidification experiments. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000421.

Oberloier:2022:OSD

[OGP22] S. Oberloier, N. Gallup, and J. M. Pearce. Overcoming supply disruptions during pandemics by utilizing found hardware for open source gentle ventilation. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000857.

Oh:2017:OSA

[OHF17] Jinook Oh, Riccardo Hofer, and W. Tecumseh Fitch. An open source automatic feeder for animal experiments. HardwareX, 1: 13–21, April 2017. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067216300050.

Owida:2022:DIL

[OhMA22] Hamza Abu Owida, Bashar Al haj Moh’d, and Mohammad Al Takrouri. Designing an integrated lowcost electrospinning device for nanofibrous scaffold fabrication. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000808.

Owida:2023:FON

[OhMAN23] Hamza Abu Owida, Bashar Al haj Moh’d, and Feras Al-Naimat. Fabricating orientated nanofibrous meshes with a bespoke ultra-cost-effective electrospinning machine. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000901.

Ohara:2022:FHP

[OKH22] Yuki Ohara, Junichi Kanie, and Katsutoshi Hori. Fabrication of a highly protective 3D-printed mask and evaluation of its viral filtration efficiency using a human head mannequin. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000591.

Oberloier:2018:OSL

[OP18] Shane Oberloier and Joshua M. Pearce. Open source low-cost power monitoring system. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300610.

Oberloier:2023:OSF

[OWHP23] Shane Oberloier, Nicholas G.

Whisman, Finn Hafting, and Joshua M. Pearce. Open source framework for a broadly expandable and reconfigurable data acquisition and automation device (BREAD). HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000743.

Perez-Alvarez:2023:MKB

[PÁBS23] Diego T. Pérez-Álvarez, Jacob Brown, and Jason Stafford. Modification of kitchen blenders into controllable laboratory mixers for mechanochemical synthesis of atomically thin materials. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000780.

Pocero:2017:OSI

[PAMC17] Lidia Pocero, Dimitrios Amaxilatis, Georgios Mylonas, and Ioannis Chatzigiannakis. Open source IoT meter devices for smart and energy-efficient school buildings. HardwareX, 1:54–67, April 2017. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067216300293.

Potticary:2018:DWS

[PAMH18] Jason Potticary, Michael P. Avery, Doug Mills, and Simon R. Hall. DON-ALD: a 2.5T wide sample space permanent magnet. HardwareX, 3:39–48, April 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067217300846.

Pontarolli:2022:DSD

[PBD^{+}22] Ricardo P. Pontarolli, Jeferson A. Bigheti, Felipe O. Domingues, Lucas B. R. de Sá, and Eduardo P. Godoy. Distributed I/O as a service: a data acquisition solution to Industry 4.0. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001006.

Pasta:2023:CBP

[PCCS23] Salvatore Pasta, Chiara Catalano, Fabrizio Crascì, and Roberta Scuoppo. A custom-built planar biaxial system for soft tissue material testing. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000822.

Poquita-Du:2023:ELC

[PDDT23] Rosa Celia Poquita-Du, Ian Peter Morgia Du, and Peter A. Todd. EmerSense: a low-cost multiparameter logger to monitor occurrence and duration of emersion events within intertidal zones. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000172.

Pantos:2023:RMG

[PDMV23] Christos Pantos, Jurrian Doornbos, Gonzalo Mier, and João Valente. The ReFiBot makers guide: Fostering academic open science and circularity with a robotic educational kit. HardwareX, 16:??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000913.

Ponzetti:2022:HPB

[PDR^{+}22] Marco Ponzetti, Ganga Chinna Rao Devarapu, Nadia Rucci, Armando Carlone, and Vittorio Saggiomo. HistoEnder: a 3D printer-based histological slide autostainer that retains 3D printer functions.

HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001158.

Pearce:2020:ESS

[Pea20] Joshua M. Pearce. Economic savings for scientific free and open source technology: a review. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300481.

Pearce:2021:AAW

[Pea21] Joshua M. Pearce. Authors from all over the world share their tech in HardwareX to battle COVID-

19. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000195.

Poulsen:2022:LDD

[PEJ^{+}22] Ebbe Poulsen, Mathias Eggertsen, Erik H. Jepsen, Claus Melvad, and Søren Rysgaard. Lightweight drone-deployed autonomous ocean profiler for repeated measurements in hazardous areas — example from glacier fronts in NE Greenland. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200058X.

Pusch:2018:LVS

[PHF18] Kira Pusch, Thomas J.

Hinton, and Adam W. Feinberg. Large volume syringe pump extruder for desktop 3D printers. HardwareX, 3:49–61, April 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300822.

Purser:2020:PMC

[PHL^{+}20] Autun Purser, Ulrich Hoge, Johannes Lemburg, Yasemin Bodur, Elena Schiller, Janine Ludszuweit, Jens Greinert, Simon Dreutter, Boris Dorschel, and Frank Wenzhöfer. PlasPI marine cameras: Open-source, affordable camera systems for time series marine studies. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300919.

Patel:2022:OST

[PHLM22] Trishna Patel, Jamie Hendren, Nathan Lee, and Aaron D. Mickle. Open source timed pressure control hardware and software for delivery of air mediated distensions in animal models. HardwareX, 11:??, April 2022.

CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000165.

Peck:2022:PDT

[PJS22] Charlie Peck, Faith Jackobs, and Emmett Smith. The PortaLyzer, a DIY tool that allows environmental DNA extraction in the field. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001183.

Park:2020:OSA

[PK20] Hyeonjun Park and Donghan Kim. An open-source anthropomorphic robot hand system: HRI hand. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300092.

Pazdzior:2021:PSC

[PK21] Robert Pazdzior and Stefan Kubicek. PlateFlo

— a software-controllable plate-scale perfusion system for culture of adherent cells. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000511.

Pazaitis:2023:TMO

[PK23] Nikolaos Pazaitis and Alexandra Kaiser. TMA-Mate: an open-source modular toolkit for constructing tissue microarrays of arbitrary layouts. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000263.

Potorti:2023:EPS

[PLP23] Francesco Potortì, Davide La Rosa, and Filippo Palumbo. Enerduino-pro: Smart meter led probe using Arduino. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000688.

Perez:2023:TDT

[PMA^{+}23] Manuel Pérez, Diego Mendez, Diego Avellaneda, Arturo Fajardo, and Carlos I. Páez-Rueda. Timedomain transmission sensor system for on-site dielectric permittivity measurements in soil: a compact, low-cost and standalone solution. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000056.

Pinares-Mamani:2020:LCD

[PMCL20] Omar Gustavo Celso Pinares-Mamani and Juan C. Cutipa-Luque. A low-cost didactic module for testing advanced control algorithms. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300572.

Pedrolli:2023:KAT

[PNM^{+}23] Lorenzo Pedrolli, Sadegh Nadimi, Sadaf Maramizonouz, Beatriz Achiaga Menor, and Alejandro López. Kinetic adhesion test to determine particle surface energy. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000445.

Poor:2022:LCO

[Poó22] Viktor Soma Poór. A low-cost and open-source mini benchtop centrifuge for molecular biology labs. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000736.

Poor:2023:PCA

[Poó23] Viktor Soma Poór. Phone cam array — an opensource, modular photogrammetry system made of Android phones. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000457.

Pringle:2020:OSA

[POP^{+}20] Adam M. Pringle, Shane Oberloier, Aliaksei L. Petsiuk, Paul G. Sanders, and Joshua M. Pearce. Open source arc analyzer: Multi-sensor monitoring of wire arc additive manufacturing. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300468.

Phooplub:2022:PDD

[POT^{+}22] Kittirat Phooplub, Sirirat Ouiganon, Panote Thavarungkul, Proespichaya Kanatharana, and Chittanon Buranachai. Portable device for dual detection of fluorescence and absorbance for biosensing or chemical sensing applications. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200013X.

Pandey:2021:SLC

[PPAP21] Santosh Pandey, Yunsoo Park, Ankita Ankita, and Gregory J. Phillips.

Scan4CFU: Low-cost, opensource bacterial colony tracking over large areas and extended incubation times. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000791.

Payne:2022:DOS

[PPF^{+}22] Matthew Payne, Francis Pooke, Harrison Fulton, Hamish Shaw, Tom Coulson, Jennifer Knopp, Lui Holder-Pearson, Jake Campbell, and J. Geoffrey Chase. Design of an open source ultra low cost insulin pump. HardwareX, 12:??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001201.

Pineda:2022:MOS

[PPG^{+}22] Daniel Pineda, Juan Pérez, Daniel Gaviria, Karen Ospino-Villalba, and Omar Camargo. MEDUSA: an open-source and webcam based multispectral imaging system. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722200027X.

Pooke:2023:LCL

[PPHP^{+}23] Francis Pooke, Matthew Payne, Lui Holder-Pearson, Doug Heaton, Jake Campbell, and J. Geoffrey Chase. Low-cost, lowpower, clockwork syringe pump. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000767.

Price:2019:APA

[Pri19] Aaron Price. An apparatus for personalized atmospheric and flight data collection aboard high altitude weather balloons. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300579.

Pacher:2023:DVM

[PRR^{+}23] Julio Pacher, Jorge Rodas, Alfredo Renault, Magno Ayala, Leonardo Comparatore, and Raul Gregor. Design and validation of a multilevel voltage source inverter based on modular H-bridge cells. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000597.

Porr:2020:BIL

[PSL^{+}20] Marc Porr, Sebastian Schwarz, Ferdinand Lange, Laura Niemeyer, Thorleif Hentrop, Daniel Marquard, Patrick Lindner, Thomas Scheper, and Sascha Beutel. Bringing IoT to the lab: SiLA2 and open-source-powered gateway module for integrating legacy devices into the digital laboratory. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300274.

Petsiuk:2020:PRA

[PTD^{+}20] Aliaksei Petsiuk, Nagendra G. Tanikella, Samantha Dertinger, Adam Pringle, Shane Oberloier, and Joshua M. Pearce. Partially RepRapable automated open source bag valve mask-based ventilator. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300407.

Palfy:2021:PRD

[PTKG21] Tamás Gábor Pálfy, Luca Török, Péter Kalicz, and Zoltán Gribovszki. 3D printed receptacle with diffuser membrane for manipulating pressurized air and water. HardwareX, 9:??, April 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000213.

Podgorski:2023:PBP

[PWC23] Rafał Podgórski, Michał Wojasiński, and Tomasz Ciach. Pushing boundaries in 3D printing: Economic pressure filament extruder for producing polymeric and polymer-ceramic filaments for 3D printers. HardwareX, 16:??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000937.

Peters:2023:DFL

[PZF23] Anna Peters, Zhu Zhang, and Sanli Faez. Dark-field light scattering microscope with focus stabilization. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000317.

Reyer:2020:HPL

[RAMM20] Sebastian Reyer, Sebastian Awiszus, Klaus Meissner, and Joachim Müller. High precision laboratory dryer for thin layer and bulk drying with adjustable temperature, relative humidity and velocity of the drying air. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300420.

Rysgaard:2022:MOP

[RBB^{+}22] Søren Rysgaard, Kim Bjerge, Wieter Boone, Egon Frandsen, Michael Graversen, Toke Thomas Høye, Bjarne Jensen, Geoffrey Johnen, Marcin Antoni Jackowicz-Korczynski, Jeffrey Taylor Kerby, Simon Kortegaard, Mikhail Mastepanov, Claus Melvad, Peter Schmidt Mikkelsen, Keld Mortensen, Carsten Nørgaard, Ebbe Poulsen, Tenna Riis, Lotte Sørensen, and Torben Røjle Christensen. A mobile observatory powered by sun and wind for near real time measurements of atmospheric, glacial, terrestrial, limnic and coastal oceanic conditions in remote off-grid areas. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000761.

Regnacq:2023:BVP

[RBRK23] Louis Regnacq, Yannick Bornat, Olivier Romain, and Florian Kolbl. BIMMS: a versatile and portable system for biological tissue and electrodetissue interface electrical characterization. HardwareX, 13:??, March 2023.

CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001328.

Reinecke:2018:IFO

[RC18] Tobias Reinecke and Brian H.

Clowers. Implementation of a flexible, open-source platform for ion mobility spectrometry. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300294.

Read:2021:VOS

[RCM21] Robert L. Read, Lauria Clarke, and Geoff Mulligan. VentMon: an open source inline ventilator tester and monitor. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000249.

Rotermund:2019:OHN

[REP19] David Rotermund, Udo A.

Ernst, and Klaus R. Pawelzik. Open hardware for neuro-prosthesis research: a study about a closed-loop multi-channel system for electrical surface stimulations and measurements. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218301202.

Rhudy:2020:WLD

[RGM20] Matthew B. Rhudy, Nathan Greenauer, and Catherine Mello. Wearable light data logger for studying physiological and psychological effects of light data. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300663.

Rodriguez:2023:LCI

[RGSGD^{+}23] Cristian F. Rodríguez, Paula Guzmán-Sastoque, Mónica Gantiva-Diaz, Saúl C. Gómez, Valentina Quezada, Carolina Muñoz-Camargo, Johann F. Osma, Luis H. Reyes, and Juan C. Cruz. Low-cost inertial microfluidic device for microparticle separation: a laserablated PMMA lab-on-a-chip approach without a cleanroom. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223001001.

Rosero-Herrera:2022:LLR

[RHAB22] J. D. Rosero-Herrera and W. Acuña-Bravo. A lower limb rehabilitation platform with mirror therapy, electrical stimulation and virtual reality for people with limited dorsiflexion movement. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200030X.

Robke:2019:SLD

[RHR19] Rhiannon Robke, Parastoo Hashemi, and Eric Ramsson. A simplified LED-driven switch for fastscan controlled-adsorption voltammetry instrumentation. HardwareX, 5: ??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721830035X.

Russell:2022:SDL

[RHWP22] Felix Russell, Alastair Hales, Gavin White, and Yatish Patel. A system for determining Li-ion cell cooling coefficients. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000025.

Rosa:2023:POM

[RJL^{+}23] Camarillo-Escobedo Rosa, Flores-Nuñez Jorge, García-Muñoz Luis, Camarillo-Escobedo Juana, and Peña-Dominguez Edgar. 3D printed opto-microfluidic autonomous analyzer for photometric applications.

HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000135.

Romuli:2023:LCD

[RJN^{+}23] Sebastian Romuli, Achim Jesser, Charles Ikenna Nwankwo, Ludger Herrmann, and Joachim Müller. Low-cost drum granulator for mechanized seedball production. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000044.

Regan:2022:DFL

[RKD^{+}22] Brian Regan, David Kinahan, Philip Daly, Richard O’Kennedy, and David Collins. Design and fabrication of a low-cost wireless camera imaging system for centrifugal microfluidics. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000049.

Raymond:2018:OSL

[RMB18] Martin A. Raymond, Thomas G. Mast, and Joseph M. Breza. An open-source lickometer and microstructure analysis program. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300105.

Romero-Morales:2019:SUM

[RMOB^{+}19] Alejandra I. Romero-Morales, Brian J. O’Grady, Kylie M. Balotin, Leon M. Bellan, Ethan S. Lippmann, and Vivian Gama. Spin∞: an updated miniaturized spinning bioreactor design for the generation of human cerebral organoids from pluripotent stem cells. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300422.

Rattan:2023:HLS

[RNRP23] Ravneet S. Rattan, Nathan Nauta, Alessia Romani, and Joshua M. Pearce. Hangprinter for large scale additive manufacturing using fused particle fabrication with recycled plastic and continuous feeding. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000081.

Romanov:2021:EIL

[Rom21] Alexey M. Romanov. An easy to implement logic analyzer for long-term precise measurements. HardwareX, 9:??, April 2021.

CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300730.

Rathod:2023:MDL

[RPHS23] Tejas Rathod, Vinay Patil, R. Harikrishnan, and Priti Shahane. Multipurpose deep learningpowered UAV for forest fire prevention and emergency response. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722300086X.

Rubio:2023:CPO

[RRS^{+}23] Joaquina Rubio, Christiam Rojas, Midori Sanchez, Daniela Gómez-Alzate, Mauricio Córdova, Verónica Montoya, Benjamin Castaneda, Javier Chang, and Sandra Pérez-Buitrago. COVOX: Providing oxygen during the COVID-19 health emergency. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001286.

Raudmae:2023:ROS

[RSV^{+}23] Renno Raudmäe, Sandra Schumann, Veiko Vunder, Maarika Oidekivi, Madis Kaspar Nigol, Robert Valner, Houman Masnavi, Arun Kumar Singh, Alvo Aabloo, and Karl Kruusamäe. ROBOTONT

— open-source and ROSsupported omnidirectional mobile robot for education and research. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000433.

Rubio:2022:MPS

[RWK22] Dani Carbonell Rubio, Willi Weber, and Enrico Klotzsch. Maasi: a 3D printed spin coater with touchscreen. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200061X.

Ryu:2022:PLC

[Ryu22a] Jae Hyeon Ryu. Prototyping a low-cost opensource autonomous unmanned surface vehicle for real-time water quality monitoring and visualization. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001146.

Ryu:2022:UBR

[Ryu22b] Jae Hyeon Ryu. UASbased real-time water quality monitoring, sampling, and visualization platform (UASWQP). HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000220.

Slocum:2018:RSD

[SAB^{+}18] Richard K. Slocum, Rachel K. Adams, Kamilah Buker, David S. Hurwitz, H. Benjamin Mason, Christopher E. Parrish, and Michael H. Scott. Response spectrum devices for active learning in earthquake engineering education. HardwareX, 4: ??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806721830021X.

Sahar:2020:DFN

[SBT20] Muhammad Sana Ullah Sahar, Matthew Barton, and Geoffrey Tansley. Design and fabrication of a nerve-stretching device for in vivo mechanotransduction of peripheral nerve fibers. HardwareX, 7: ??, April 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300018.

Suciu:2022:AOD

[SBTPV22] Ioana Suciu, Guillem Boquet, Pere Tuset-Peiró, and Xavier Vilajosana.

ADO: an open digital endto-end tank based aquaculture platform. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000281.

Sanchez:2020:OOS

[SDDL20] Carlos Sánchez, Paolo Dessì, Maeve Duffy, and Piet N. L. Lens. OpenTCC: an open source low-cost temperature-control chamber. HardwareX, 7: ??, April 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300080.

Seifer:2023:SSP

[SE23] Shahar Seifer and Michael Elbaum. Synchronization of scanning probe and pixelated sensor for image-guided diffraction microscopy. HardwareX, 14:??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722300038X.

Sun:2022:CFE

[SFWW22] Wenhuan Sun, Adam Feinberg, and Victoria Webster-Wood. Continuous fiber extruder for desktop 3D printers toward long fiber embedded hydrogel 3D printing. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000426.

Saun:2021:DVI

[SG21] Tomas J. Saun and Teodor P. Grantcharov. Design and validation of an inertial measurement unit (IMU)- based sensor for capturing camera movement in the operating room. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000080.

Spinelli:2019:LCA

[SGD19] Gerardo M. Spinelli, Zach L. Gottesman, and Jonathan Deenik. A low-cost Arduino-based datalogger with cellular modem and FTP communication for irrigation water use monitoring to enable access to CropManage. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067219300070.

Samieian:2022:LSI

[SGD^{+}22] Mohammad Amin Samieian, Carlos E. Garcia, Laura Bravo Diaz, Alastair Hales, Yatish Patel, and Gregory J. Offer. Large scale immersion bath for isothermal testing of lithiumion cells. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001043.

Staat:2023:PTM

[SGG23] Christoph Staat, Norbert Gattinger, and Bernhard Gleich. PLUSPULS: a transcranial magnetic stimulator with extended pulse protocols. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001250.

Susko:2018:AFC

[SGH^{+}18] Alexander Q. Susko, Fletcher Gilbertson, D. Jo Heuschele, Kevin Smith, and Peter Marchetto. An automatable, field camera track system for phenotyping crop lodging and crop movement. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300348.

Schnabel:2023:LCT

[SHS^{+}23] Tobias Schnabel, Robert Honke, Andreas Schmid, Simon Mehling, Rene’ Göhring, Oldrich Simek, Axel Wolfram, Andre Wetterauer, and Christian Springer. Low-cost test rig for characterization of photocatalytic planar materials using photonically sized UV-A LED light sources. HardwareX, 16: ??, December 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000949.

Schlatter:2018:PPV

[SIR18] Samuel Schlatter, Patrin Illenberger, and Samuel Rosset. Peta-pico-Voltron: an open-source high voltage power supply. HardwareX, 4:??, October 2018. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300531.

Serioli:2023:AOS

[SIY^{+}23] Laura Serioli, Atsushi Ishimoto, Akinobu Yamaguchi, Kinga Zór, Anja Boisen, and En-Te Hwu. APELLA: Open-source, miniaturized all-in-one powered lab-on-a-disc platform. HardwareX, 15: ??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000561.

SreeSaiGanesh:2022:IBP

[SJB22] Kakumanu Vamsi Sree Sai Ganesh, S. P. Shibu Jeyanth, and A. Ruhan Bevi. IOT based portable heart rate and SpO2 pulse oximeter. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000542.

Soleymani:2021:GAU

[SK21] Hamid Soleymani and Steven Kidder. A Griggs apparatus upgrade for stress-controlled testing of geological material at high temperature and pressure. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000018.

Sands:2023:LCO

[SK23] Matthew Sands and Jinki Kim. A low-cost and open-source measurement system to determine the Young’s and shear moduli and Poisson’s ratio of soft materials using a Raspberry Pi camera module and 3D printed parts. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001316.

Salem:2020:DVR

[SKC^{+}20] Ghadi Salem, Jonathan Krynitsky, Noah Cubert, Alex Pu, Simeon Anfinrud, Jonathan Pedersen, Joshua Lehman, Ajith Kanuri, and Thomas Pohida. Digital video recorder for Raspberry PI cameras with multi-camera synchronous acquisition. HardwareX, 8:??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300699.

Siguerdidjane:2022:LCO

[SKG22] Wassime Siguerdidjane, Farbod Khameneifar, and Frédérick P. Gosselin. A low-cost open-source automated shot peen forming system. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000086.

Smolka:2023:SPW

[SKK^{+}23] Jan Smolka, Kamila Kempná, Petr Kučera, Kamil Kempný, Eleni Asimakopoulou, and Pavel Danihelka. Setup of a 3D printed wind tunnel: Application for calibrating bi-directional velocity probes used in fire engineering applications. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000470.

Singh:2023:SLC

[SKS23] Saurabh Kumar Singh, Avinash Kumar, and Pranav R. Shirhatti. A simple and low-cost setup for part per billion level frequency stabilization and characterization of red He– Ne laser. HardwareX, 14:??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000287.

Sahu:2022:HLS

[SL22] Govind N. Sahu and Mohit Law. Hardware-in-the-loop simulator for emulation and active control of chatter. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000189.

Sandby:2021:CEL

[SMB^{+}21] Lucas M. Sandby, Jens E. B. Mejdahl, Simon H. Bjerregaard, Claus Melvad, and Søren Rysgaard. A cost-efficient low-weight autonomous profiler for measurements in polar coastal waters and other regions with strong density gradients. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000365.

Scutelnic:2022:TEB

[SMS^{+}22] Dumitru Scutelnic, Giacomo Marchioro, Salvatore Siracusano, Paolo Fiorini, Riccardo Muradore, and Claudia Daffara. Thermal endoscope based on cost-effective LWIR camera cores. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000451.

Samples:2023:OAP

[SMS^{+}23] Robert Samples, Riko Mukoyama, Jacob Shaffer, Jill Mikucki, and Lesley-Ann Giddings. OpenASAP: an affordable 3D printed atmospheric solids analysis probe (ASAP) mass spectrometry system for direct analysis of solid and liquid samples. HardwareX, 16:??, December 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000974.

Stucker:2021:CCN

[SMSR21] Andrew M. Stucker, Ethan Morris, Christopher J. Stubbs, and Daniel J. Robertson. The Crop Clamp — a non-destructive electromechanical pinch test to evaluate stalk lodging resistance. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000559.

Semenzin:2020:OSA

[SMW^{+}20] Clayton S. Semenzin, Martin Mapley, Eric Wu, Jo P. Pauls, Benjamin Simpson, Shaun D. Gregory, and Geoff Tansley. Opensource automated centrifugal pump test rig. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300493.

Sugimoto:2020:MSS

[SNAO20] Yasuhiro Sugimoto, Keisuke Naniwa, Hitoshi Aonuma, and Koichi Osuka. Microinjection support system for small biological subjects. HardwareX, 7: ??, April 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300110.

Shnier:2023:RAC

[SOB^{+}23] Adam Shnier, Francis Otieno, Caren Billing, Daniel Wamwangi, and David G. Billing. Robust Arduino controlled spin coater using a novel and simple gravity chuck design. HardwareX, 14: ??, June 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000299.

Sabattini:2022:AAS

[SRB^{+}22] J. A. Sabattini, J. M. Reta, L. A. Bugnon, J. I. Cerrudo, R. A. Sabattini, A. Peñalva, M. Bollazzi, M. O. Paz, and F. Sturniolo. AntVideo-Record: Autonomous system to capture the locomotor activity of leafcutter ants. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000153.

Smith:2022:LCC

[SRC^{+}22] Rachel Smith, Amelia Rolfe, Chris Cameron, Geoffrey M. Shaw, J. Geoffrey Chase, and Christopher G. Pretty. Low cost circulatory pressure acquisition and fluid infusion rate measurement system for clinical research. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000633.

Schaufler:2023:LCM

[SS23] Ruby L. Schaufler and Niall C. Slowey. Lowcost modular chromatography column rack and vial holders. HardwareX, 13:??, March 2023. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200133X.

Scholze:2020:STT

[SSL^{+}20] Mario Scholze, Sarah Safavi, Kai Chun Li, Benjamin Ondruschka, Michael Werner, Johann Zwirner, and Niels Hammer. Standardized tensile testing of soft tissue using a 3D printed clamping system. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300687.

Smith:2022:URD

[SSM^{+}22] Corinne Smith, Joud Satme, Jacob Martin, Austin R. J. Downey, Nikolaos Vitzilaios, and Jasim Imran. UAV rapidlydeployable stage sensor with electro-permanent magnet docking mechanism for flood monitoring in undersampled watersheds. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000700.

Shiwarski:2020:PBS

[STE^{+}20] Daniel J. Shiwarski, Joshua W.

Tashman, Amity F. Eaton, Gerard Apodaca, and Adam W. Feinberg. 3D printed biaxial stretcher compatible with live fluorescence microscopy. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300031.

Skrzypczak:2020:OSH

[STP20] Noah G. Skrzypczak, Nagendra G. Tanikella, and Joshua M. Pearce. Open source high-temperature RepRap for 3-D printing heat-sterilizable PPE and other applications. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300390.

Suriano:2021:PAQ

[Sur21]  Domenico Suriano.  A

portable air quality monitoring unit and a modular, flexible tool for on-field evaluation and calibration of low-cost gas sensors. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000274.

Salazar-Vazquez:2020:PPH

[SVMV20] Jairo Salazar-Vazquez and Andres Mendez-Vazquez. A plug-and-play hyperspectral imaging sensor using low-cost equipment.

HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300513.

Shaid:2019:LCB

[SWPG19] Abu Shaid, Lijing Wang, Rajiv Padhye, and Martin Gregory. Low cost bench scale apparatus for measuring the thermal resistance of multilayered textile fabric against radiative and contact heat transfer. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218301226.

Sebe:2022:NMV

[SZZ22] István Sebe, László Zsidai, and Romána Zelkó. Novel modified vertical diffusion cell for testing of in vitro drug release (IVRT) of topical patches. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000384.

Toschke:2021:MOH

[TBBI21] Yannic Toschke, Bjoern Bourdon, Dirk Berben, and Mirco Imlau. A modular optical honeycomb breadboard realized with 3D-printable building bricks and industrial aluminum extrusions. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000110.

Townsley:2020:MAU

[TBM20] Cody R. Townsley, Joseph M. Breza, and Thomas G. Mast. Movement assay for the undergraduate neuroscience laboratory. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722030002X.

Thompson:2018:PMF

[TCC^{+}18] Alison L. Thompson, Alex Conrad, Matthew M. Conley, Henry Shrock, Bryce Taft, Charles Miksch, Talon Mills, and John M. Dyer. Professor: a motorized field-based phenotyping cart. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300038.

Tai:2021:OMR

[TCG^{+}21] Albert Tai, Michael Chun, Yuqiu Gan, Mert Selamet, and Hod Lipson. PARA: a one-meter reach, two-kg payload, three-DoF open source robotic arm with customizable end effector. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000389.

Thomsen:2020:AMI

[THL^{+}20] Steffen Thomsen, Mads Holm Hansen, Jeppe Pinholt Lillethorup, Frederik Sebastian Tirsgaard, Adam Flytkjær, Claus Melvad, Søren Rysgaard, and Daniel F. Carlson. An affordable and miniature ice coring drill for rapid acquisition of small iceberg samples. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300109.

Toloza:2022:MHS

[THMZ22] Juan Manuel Toloza, Matías Hirsch, Cristian Mateos, and Alejandro Zunino. Motrol: a hardwaresoftware device for batch benchmarking and profiling of in-lab mobile device clusters. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000852.

Tsow:2021:LCW

[TKHB21] Francis Tsow, Anupam Kumar, SM Hadi Hosseini, and Audrey Bowden. A low-cost, wearable, do-it-yourself functional near-infrared spectroscopy (DIY–fNIRS) headband. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722100033X.

Toschke:2022:DLB

[TLO^{+}22] Yannic Toschke, Janet Lusmoeller, Lars Otte, Johann Schmidt, Svenja Meyer, Alexander Tessmer, Christian Brockmann, Milena Ahuis, Emma Hüer, Christian Kirberger, and Dirk Berben. Distributed LoRa based CO_{2} monitoring network

— a standalone open source system for contagion prevention by controlled ventilation. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000062.

Triwiyanto:2023:BME

[TLP^{+}23] Triwiyanto Triwiyanto, Sari Luthfiyah, I. Putu Alit Pawana, Abdussalam Ali Ahmed, and Alcham Andrian. Bilateral mode exoskeleton for hand rehabilitation with wireless control using 3D printing technology based on IMU sensor. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000391.

Tuljak:2022:LCP

[TLŠ^{+}22] Marko Tuljak, David Lajevec, Rok Štanc, Špela Zemljič Jokhadar, and Jure Derganc. Low-cost programable stroboscopic illumination with sub-microsecond pulses for high-throughput microfluidic applications. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001122.

Tiersch:2022:DPV

[TPLT22] Nolan J. Tiersch, Jacqueline Paulsen, Yue Liu, and Terrence R. Tiersch. A 3-D printed vitrification device integrated with French straws. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222001110.

Trevathan:2022:OSI

[TS22] Jarrod Trevathan and Simon Schmidtke. Opensource Internet of Things remote aquatic environmental sensing. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000815.

Tashman:2021:HPO

[TSF21] Joshua W. Tashman, Daniel J. Shiwarski, and Adam W. Feinberg. A high performance opensource syringe extruder optimized for extrusion and retraction during FRESH 3D bioprinting. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300791.

Takara:2021:IRG

[TTA^{+}21] Grant Takara, A. Zachary Trimble, Reika Arata, Shane Brown, Hector Jaime Gonzalez, and Camilo Mora. An inexpensive robotic gantry to screen and control soil moisture for plant experiments. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000031.

Utter:2020:OSF

[UB20] Brent Utter and Alexander Brown. Open-source five degree of freedom motion platform for investigating fish-robot interaction. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300158.

Ulrich:2019:OSW

[Ulr19] Burkhard Ulrich. Opensource wideband (DC to MHz range) isolated current sensor. HardwareX, 5:??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300592.

Utter:2019:OSE

[UMEA19] Brent Utter, Rachel Marbaker, Kevin Eschen, and Julianna Abel. Opensource experimental setup for investigating the actuation behavior of active textiles. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300264.

Villanueva:2021:OSP

[VA21] Emanuel Martinez Villanueva and Rafiq Ahmad. An open-source powered and ergonomic personal protective respirator for frontline COVID-19 response. HardwareX, 10: ??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000523.

Venkatesh:2022:DOS

[VAG^{+}22] Kavyashree Venkatesh, Sourabh Suresh Alagundagi, Vikas Garg, Krupakar Pasala, Deval Karia, and Manish Arora. DripOMeter: an open-source optoelectronic system for intravenous (IV) infusion monitoring. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000906.

vanAmelrooij:2022:BBB

[vAvdGP^{+}22] Edward van Amelrooij, Nick van de Giesen, Jeroen Plomp, Michel Thijs, and Tomáš Fico. BLOSM: Boron-based large-scale observation of soil moisture: First laboratory results of a cost-efficient neutron detector. HardwareX, 12:??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000876.

Vadivel:2022:MPC

[VBK^{+}22] Dhanalakshmi Vadivel, Diego Savio Branciforti, Othman Kerroumi, Marco Dondi, and Daniele Dondi. Mostly 3D printed chemical synthesis robot. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000554.

Vaishampayan:2023:MPB

[VCS^{+}23] Vijay Vaishampayan, Oinam Robita Chanu, Balasubramanian Sivasamy, Muthamilselvi Ponnuchamy, Varshini Karthik, Ambar Pendharkar, Lohith Srinivas Thotakura, Aryan Prabhu, Venkatesan Dhananjeyan, and Ashish Kapoor. Microfluidic paper-based device coupled with 3D printed imaging box for colorimetric detection in resource-limited settings. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000639.

Veldscholte:2022:OHC

[VdB22] Lars B. Veldscholte and Sissi de Beer. Open-Humidistat: Humiditycontrolled experiments for everyone. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000335.

vandenBerg:2022:OBD

[vdBSRW22] Sander C. van den Berg, Rob B. N. Scharff, Zoltán Rusák, and Jun Wu. Open-Fish: Biomimetic design of a soft robotic fish for high speed locomotion. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000657.

Vlasov:2023:LCE

[VDV23] Evgenii Vlasov, Nikita Denisov, and Johan Verbeeck. Low-cost electron detector for scanning electron microscope. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000202.

Valtr:2023:SPM

[VKM^{+}23] Miroslav Valtr, Petr Klapetek, Jan Martinek, Ondřej Novotný, Zdeněk Jelínek, Václav Hortvík, and David Nečas. Scanning probe microscopy controller with advanced sampling support. HardwareX, 15:??, September 2023. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000585.

Valencia-Ortiz:2022:DSA

[VOS22] Milton Valencia-Ortiz and Sindhuja Sankaran. Development of a semiautomated volatile organic compounds (VOCs) sampling system for field asymmetric ion mobility spectrometry (FAIMS) analysis. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S246806722200089X.

Vestad:2022:LCV

[VS22] Håvard Vestad and Martin Steinert. A low-cost vibration isolation chamber — making high precision experiments accessible. HardwareX, 11: ??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000098.

Vaut:2019:FRA

[VSTB19] Lukas Vaut, Ermes Scarano, Guido Tosello, and Anja Boisen. Fully replicable and automated retention measurement setup for characterization of bioadhesion. HardwareX, 6: ??, October 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300173.

Winters:2021:POS

[WBD^{+}21] Brandon J. Winters, Nick Banfield, Cassandra Dixon, Anna Swensen, Dakota Holman, and Braxton Fillbrown. 3D-Printable and open-source modular smartphone visible spectrophotometer. HardwareX, 10:??, October 2021. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000614.

Wright:2022:FOS

[WCR22] Harry C. Wright, Duncan D. Cameron, and Anthony J. Ryan. FoamPi: an open-source Raspberry Pi based apparatus for monitoring polyurethane foam reactions. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001109.

Wu:2021:PPL

[WDW21] Jin Wu, David Dellal, and Steven Wasserman. Prokaryote playhouse: a low-cost, laser-cut acrylic incubator for optogenetic bacterial culture. HardwareX, 9:??, April 2021. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067221000134.

Wang:2023:HPP

[WLML23] Sheng Wang, Baichen Li, David McLeod, and Zhenyu Li. A handheld plug-and-play microfluidic liquid handling automation platform for immunoassays. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000275.

Williams:2019:TFB

[WM19] James Williams and Ilya Mikhelson. Triple frame buffer FPGA implementation. HardwareX, 5:

??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300798.

Woern:2018:RRO

[WMPP18] Aubrey L. Woern, Joseph R.

McCaslin, Adam M. Pringle, and Joshua M. Pearce. RepRapable recyclebot: Open source 3-D printable extruder for converting plastic to 3-D printing filament. HardwareX, 4:??, October 2018. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218300208.

Wankhede:2022:PDS

[WNKP22] Pankaj Wankhede, Nara Guru Narayanaswamy, Suresh Kurra, and Amrita Priyadarshini. A portable device for single point strain analysis in sheet metal forming processes. HardwareX, 12: ??, October 2022. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806722200116X.

Wiranata:2022:ETT

[WOM^{+}22] Ardi Wiranata, Yunosuke Ohsugi, Ayato Minaminosono, Yu Kuwajima, and Shingo Maeda. Electromechanical tensile test equipment for stretchable conductive materials.

HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000323.

Winterl:2020:MCT

[WRH^{+}20] Alexander Winterl, Sebastian Richter, Aymeric Houstin, Anna P. Nesterova, Francesco Bonadonna, Werner Schneider, Ben Fabry, Céline Le Bohec, and Daniel P. Zitterbart. micrObs — a customizable time-lapse camera for ecological studies. HardwareX, 8:??, October 2020. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067220300432.

Woo:2023:WOI

[WRSU23] Winnie Woo, William Richards, John Selker, and Chet Udell. Weather-Chimes: an open IoT weather station and data sonification system. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067223000093.

White:2018:CML

[WS18a] Jonathan A. White and Aaron M. Streets. Controller for microfluidic large-scale integration. HardwareX, 3:135–145, April 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S246806721730041X.

Winters:2018:POC

[WS18b] Brandon J. Winters and David Shepler. 3D printable optomechanical cage system with enclosure. HardwareX, 3:62–81, April 2018. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067217300421.

Watson:2019:AOA

[WS19] Craig Watson and Samuel Senyo. All-in-one automated microfluidics control system. HardwareX, 5:??, April 2019. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067218301147.

Williams:2020:LCO

[WSB^{+}20] Mitchell Williams, Stuart Sater, Colin Burkhalter, Stephen Schoonen, Jacob Miller, Dev Shrestha, Michele R. Brumley, and Nathan R. Schiele. Lowcost, open-source, variable speed and incline treadmill for studying impacts of neonatal locomotion. HardwareX, 7:??, April 2020. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300055.

Wijayanti:2022:ULC

[WSI^{+}22] Ika Dewi Wijayanti, Ari Kurniawan Saputra, Faris Ibrahim, Amaliya Rasyida, Putu Suwarta, and Indra Sidharta. An ultra-low-cost and adjustable inhouse electrospinning machine to produce PVA nanofiber. HardwareX, 11:??, April 2022. CO-DEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222000608.

Wang:2019:OEF

[WSL^{+}19] Benjamin Wang, Rayan Sud, Michael Leung, Mimi Yang, Jesse A. Rodriguez, Ricky Lee, and Mark Cappelli. OpenEM — electromagnetic field mapping robot for microwave and RF measurements. HardwareX, 5:??, April 2019. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067218300622.

Weaver:2023:OSH

[WYT23] Isaac A. Weaver, S. Aryana Yousefzadeh, and Michael R. Tadross. An open-source head-fixation and implantprotection system for mice. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2468067222001365.

Yensen:2019:OSA

[YA19] Nicholas Yensen and Peter B. Allen. Open source all-iron battery for renewable energy storage. HardwareX, 6:??, October 2019. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067219300318.

Yang:2023:PCM

[YLW^{+}23] Anxiong Yang, Falk Nicolas Lein, Joana Weiler, Julian Drechsel, Vanessa Schumann, Felix Erichson, André Streek, and Richard Börner. Pressurecontrolled microfluidics for automated single-molecule sample preparation. HardwareX, 14:??, June 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000329.

Yingst:2023:PTL

[YM23] Andrew Yingst and Vuk Marojevic. Power tether for long duration multicopter flight. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000731.

Yang:2021:DLC

[YXES21] Yusheng Yang, Jun Xu, Willemijn S. Elkhuizen, and Yu Song. The development of a low-cost photogrammetry-based 3D hand scanner. HardwareX, 10:??, October 2021. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067221000419.

Zitterbart:2022:TLC

[ZBOB22] Daniel P. Zitterbart, Alessandro Bocconcelli, Miles Ochs, and Julien Bonnel. TOSSIT: a low-cost, hand deployable, rope-less and acoustically silent mooring for underwater passive acoustic monitoring. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000499.

Zimmermann:2020:LZO

[ZCB^{+}20] M. J. Y. Zimmermann, A. Maia Chagas, P. Bartel, S. Pop, L. L. Prieto-Godino, and T. Baden. LED Zappelin’: an open source LED controller for arbitrary spectrum visual stimulation and optogenetics during 2-photon imaging. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300365.

Zhang:2020:RLB

[ZR20] Haipeng Zhang and Sangjin Ryu. Rotating-liquid-based hydrogel bead generator. HardwareX, 8: ??, October 2020. CO-DEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067220300304.

Zinzindohoue:2023:PTA

[ZSLF23] Coffi Gérard Franck Zinzindohoué, Timm Schoening, Estanislau Baptista Lima, and Björn Fiedler. PlasPi TDM: Augmentation of a low-cost camera platform for advanced underwater physical-ecological observations. HardwareX, 15:??, September 2023. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067223000779.

Zhang:2022:COS

[ZTS22] Yushen Zhang, Tsun-Ming Tseng, and Ulf Schlichtmann. ColoriSens: an open-source and low-cost portable color sensor board for microfluidic integration with wireless communication and fluorescence detection. HardwareX, 11:??, April 2022. CODEN ???? ISSN 2468-0672. URL http://

www.sciencedirect.com/
science/article/pii/S2468067222000578.